A Complete Bibliography of Publications in the Journal of Statistical Computation and

Simulation: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 April 2024 Version 2.28

Title word cross-reference

$ [Sha23d]. 1 [SYA23]. $109.00 [Sha20d]. $110.00 [Sha23e]. $118.10 [Sha22b]. $129.95 [Sha21c]. $130.00 [Sha23f]. $131.04 [Ano22]. $131.49 [Sha21b]. $139.95 [Sta21]. $144.49 [Sha22b]. $159.99 [Sha23a]. $164.00 [Sha20b]. $17.99 [Sha20i]. $18.99 [Sha20i]. 2 [SYA23]. $208.94 [Sha20e]. $244.76 [Sha23e]. $30.99 [Sha21e]. $37.37 [Sha20c]. $50.99 [Sha21d]. $67.15 [Che20a]. $74.21 [Sha23c]. $87.70 [Sha20a]. $93.22 [Sha22b]. $94.75 [Sha23b]. $94.99 [Ano23a]. $99.95 [Sha22a, Sha23d]. α [KLdS21]. X̄ [MLA^{+}20]. X̄ − S [ACHS23]. C [MP20, AZL21]. C(α) [AA22b]. C_{pmk} [DZS21]. C_{pyk} [Bal21]. χ^{2} [YZK22]. d [SYVV20, SSV21]. F [Che22b, CLUX21]. K [CHEH21, Nik22, WXWS21, ZADA22, ACS21b, DK22, JS22b, PY24, RLACE20]. l_{0} [KLL^{+}23]. l_{1} [BYH23]. L_{2} [Zhe21]. M [DWWW23, LWG20, WKS23, DS22]. n [DK22]. p [AKA22, GMCL20, GCF^{+}20, HD22b, LJW23, TTWG20, WX21]. P (X > Y) [JRM20]. ϕ [DWWW23]. q [AK23]. r [dSdNB22]. S_{pmk} [DS20]. T [KKLL24, PB23, HVAVD20, Ros22, TAWL21]. T^{2} [KK21, SAR20]. U [BVC24, CS23, Das22, FSB20, WW22a]. W [PB23]. X^{−} [TMMCS21]. z [Che22b].

-ANA [DS22]. -digests [Ros22]. -divergence [MP20]. -estimates [LWG20]. -estimators [PB23, WKS23]. -finite [Nik22]. -Hotelling [SAR20]. -inflated [ACS21b]. -larger [dSdNB22]. -means [AZL21]. -mixed [ZADA22]. -mixing [DWWW23]. -order [GMCL20]. -out-of- [DK22]. -posterior [SSV21]. -records [JS22b]. -risk [SYVV20]. -square [KKLL24]. -stable [KLdS21]. -statistic [CLUX21, WW22a]. -statistics [BVC24, CS23, Das22, FSB20]. -test [Che22b, Che22b]. -type [AKA22]. -value [TTWG20]. -values [HD22b, WX21].

0 [Sha20i, Sha23e].

1 [Che20a, Sha21e]. 1-119-28201-2 [Che20a]. 19 [SOBC23]. 1d [LC20].

2 [Che20a]. 2-stage [BPT23]. 2PL [TC21].

3 [Ano23a, Sha22b]. 3-component [CA20]. 33487 [Sha23c].

4 [Sha20c, Sha23f]. 4th [Sha21e, Sta21].

5 [Sha22a].

6 [Sha20d, Sha23b].

7 [Sha21d, Sha22b].

8 [Ano22].

9 [Sha22b, Sta21]. 978 [Ano22, Ano23a, Sha20c, Sha20d, Sha20i, Sha21d, Sha21e, Sha22b, Sha22a, Sha23b, Sha23f, Sha23e, Sta21]. 978-0-367-15206-2 [Sha21b]. 978-0-367-27798-7 [Sha23c, Sha23e]. 978-0-367-36547-9 [Sha22b]. 978-0-367-55054-7 [Sha22b]. 978-0-367-85940-4 [Sha23f]. 978-1-032-15373-5 [Sha23d]. 978-1-138-59152-3 [Sha22b]. 978-1-1381-9745-9 [Sta21]. 978-1-285-05088-1 [Sha21e]. 978-1-4665-1585-7 [Sha21d]. 978-1-4822-3742-9 [Sha21c]. 978-1-498-75202-2 [Sha20e]. 978-1-4987-2798-3 [Sha20a]. 978-1-4987-3716-6 [Sha23b]. 978-1-4987-6460-5 [Sha22a]. 978-1-5225-2512-7 [Sha20b]. 978-1-5416-1851-0 [Sha20i]. 978-1-5416-9947-2 [Sha20i]. 978-3-030-72436-8 [Ano22].

978-3-030-72439-9 [Sha23a]. 978-3-11-069780-3 [Ano23a]. 978-3-319-388865-6 [Sha20d]. 978-3-718-65314-0 [Sha23e]. 978-981-4616-32-4 [Sha20c].

ABC [Tsi20]. abrupt [QH21]. ACC [JNNTMS24]. accelerated [AA22a, CLY22, CC21, FH22, HLC21, MTW21]. acceptance [DSM22]. accuracy [WS20, ZHF21]. accurate [SC21]. Active [DXF^{+}23]. Active-set [DXF^{+}23]. Adaptation [Avc24]. Adaptive [GJQ22, HYWW21, JZY24, WB22, YZ21, ZWW21, AFS21, CF21, EES24, EN23, FFMZ22, GJP24, HR20, HKLA21, HSK22, HA23, HCZ^{+}21, KSL22, KV23, LTL20, LW21b, LSL21, MC21, MYM23, uAR21, uAS23, SCA20, Sab23, SG23, WWW22, YXC21, ZSWN21]. Adaptive-weighted [YZ21]. Additive [ZZY22, ASP22, Che20c, DXF^{+}23, EGA21, FSLW21, HTX23, HW20, Kim22, LLZ20, MHB23, ZFZ22, ZY22, ZLY23, ZLL24]. ade4 [Sha20n]. Adjusted [ZLZ24, CM21, KAN21, SAMA21]. advances [Sha21g]. AEPD [KL22b]. affects [Sha20c]. AFT [SRU23]. AG [Sha23a]. age [BKM21]. ageing [FPMC20]. aggregation [HFM21]. agreement [AH22]. aid [CTPY22]. Alan [Sha20e]. algebra [RH21]. algorithm [BHAK21, BPT23, CH21, CHW21, CC21, DXF^{+}23, DJA22, LZZ22, MYM23, MH21, NSK21, Nis23, PP24, SG23, Sto22, TZ22, VS20, WZ21, WW21, YY20, ZG21]. algorithm-based [DXF^{+}23]. Algorithmic [Sha23b]. algorithms [BC23, CHG21, DGMN22, HCN20, MMH22, NI23, NF23, RNSR20, RN21]. allocated [DN22]. allocation [Das22, ES21, HAV20, KAQ21, MNAP22, NAJ21, VPMA23, ZADA22]. allowing [WTZ20]. almost [PFMT20]. alter [SSS20]. altered [MAY23]. Alternative [San23, Chr22, HMB22, Li24]. Alzheimer [CHJS20]. among [PZWL23, QJFY20, RDSG21]. ANA [DWXW22, DS22]. analysing [GMFB21, MN22, MAY23]. Analysis [AA23, EN23, GSH20, MAKRK22, MMH20, Sha22a, XLF22, ADA22, AMO20, BSH22, BP24b, CIPCR23, CA20, CHJS20, Che22a, CG21, DC23, DSH24, FASuA21, FSB20, GMCL20, GB22b, Han23, HLC21, HD22b, HH21, HWF22, IS23b, KKS20, Kim22, KCC22, Lee22, LLM23, LSL21, LTW^{+}23b, LTY^{+}24, MN22, PG22, RMF^{+}21, SK21b, SB20, SBH21, Sha20n, Sha21k, Sha22b, SYG24, SRU23, TC21, WXWS21, WWW22, XYZ20, ZCD22, ZRB^{+}20, ZLLZ22, tB22, Sha20d, Sha20a, Sha21b, Sta21]. analytical [PS21]. analytics [Sha20o, TLGW20, Ano23a]. analyzing [Fer23, HMB22]. Anderson [KB22]. answer [ZSS23]. any [MGA21]. Application [BLZ21, MHKS22, ADA22, Akg23, AFAK24, Ano23c, AÖ21, BHA22, Cha23, CHJS20, DM22, DWXW22, DSSNL20, DDD23, EES24, GPH23, GM22, GCF^{+}20, GYL24, IS23b, JS22b, KSJ24, KJL23, Lee20, LK22, LWWD20, LBHX23, MWMCR20, MN22, NN22, NN23a, PB21, QMK21, QH21, SAB21, TLGW20, TÖSY21, WKS23, WL21]. applications [ASAP21, AKAA22, DSW20, DS22, Han23, HZD23, KAQ23, KM21, KN22, LTY^{+}24, MAKRK22, MSR23, NAV23, NN23b, PMW22, Sha20m, SZT^{+}23, WW23, Sha20g]. applied [NJ20, dSdNB22]. approach [Aça23, ALAS21, Ali20, BP21a, BP24a, BHK23, CM21, DHS24, EGA21, GSRY23, HMB22, HPV23, KD20, KLdS21, KAQ21, KKZ23, KÖ21, LGM20, MN23, DRG23, MH21, QH21, RMM24, SRG22, Sha21k, SSV21, SCK23, SMZ20, SK21c, Ten20, Tsi20, WLLH22, WTZ20, WWW22, YK20, YH20, ZXDD22, dSdNB22, Sha21c]. approaches [DDL^{+}20, MHB23, TWS23, dSCdC22]. Approximate [BGV20, KH23, SCKW21, DGMN22, TTWG20]. Approximately [WLZ21]. Approximation [Pho23, DSW20, GFPG22, NS22, YSJL20]. approximations [MJG21, SC21, ZGO24]. Archimax [GH23]. Archimedean [GH23]. area [LLP^{+}23, NET^{+}24, RMN21, Sha20p, SKO20]. Arieh [Sha20c]. ARIMA [XQ21]. arising [NS22]. arm [LWS20]. ARMA [AJN21, CCM20, CF21, JNFTM^{+}21, JNNTMS24, SK20]. arrival [MRNLB20]. ARSV [DT23]. Art [Sha20i]. artificial [FPM22, TWN22, Sha23f]. aspect [Sha20c]. Assessing [AP20b, NJR22, NN22]. assessment [Kel21, Mod22, TWZ23]. assignable [SAR20]. assignment [KR21]. assisted [KV23]. associated [AMO22, LLP^{+}23, ZYW22]. association [Sha21f]. assuming [CA20]. assurance [KJB22]. ASTD [KL22b]. Asymmetric [MWMCR20, Ros22, Tsi20, YY23, dSAF22]. asymmetry [BHK23]. Asymptotic [DWWW23, EH22, LLBW23, ATH20]. Asymptotically [GYL24]. attributes [QJFY20, QHCB20, TK24]. AUC [LLX24]. augmentation [LCH20, PG22, SM21, WWW22]. auto [MWMCR20]. auto-regressive [MWMCR20]. autocorrelated [CF21, JNNTMS24, SCA20, ZRB^{+}20]. autocorrelation [Sab23, WS22]. Automotive [FPM22]. autoregressive [AZS^{+}24, BBC21, BL20b, GYL24, HPV23, KL22a, KSAR21, KCO20, KNM23, KB24, KL24, Lee20, LLM23, LJ23, LG20, LYP20, LWS^{+}23, MN22, dSMA22, MG23, YWW23a, YLYD24]. auxiliary [ANA^{+}23, ARA^{+}21, CLL23, CTPY22, GAN^{+}20, KRK^{+}21, Ozg21, SSS^{+}22, THL^{+}22, XCXZ22]. auxiliary-information [ANA^{+}23]. availability [GWL21]. Avenue [Sha20b]. average [AAA^{+}22, ARA^{+}21, ASAA23, AK23, CLL23, HPV23, HZXS24, ISZ^{+}23, KCC22, dSMA22, Mir21, MSR23, uAT20, uAR21, SMZ20, TMMCS21, WQZS23]. average-CUSUM [ASAA23]. averaging [DXY21]. avoid [Sha20c]. AZIAD [MAY23].

Backward [CJK21]. Bahadur [WYBW22]. Bahadur-type [WYBW22]. balance [DJA22]. balanced [Kel21, OSG20, SÇK24]. bands [BM22]. Bandwidth [HPP23, Ten20, FSGMM23, JIK22, KK20]. bar [QHCB20]. barycenters [PDD^{+}21]. based [AE23, Akg23, AMD21, ANRS20, AZL21, AKA22, Asa22, Bal21, BS21, BLA20, BA21, BHAK21, BAAR21, BGGH21, CXH21, CMMP20, CH21, CHW21, CLY23, CLL23, CMZ24, CS23, Das22, DK22, DXF^{+}23, DSSNL20, DA23, ES21, ESA23, FWC24, GTMP22, GvRS23, GMJ23, HFM21, HRB21, HLC21, JNFTM^{+}21, JIK22, Jia21, JRM20, KB20, KAN21, KB22, KL24, KM22b, KGK24, Lee20, LKS21, LS21, LK23, LJ23, Li20, Li21, LRN22, LTS^{+}23, MXYY22, MTW21, MP20, MB22, MN22, MK20, Mod22, MC21, MP21, MB23, MH21, NET^{+}24, NSK21, uA20, NFP21, NN23a, Ozg21, PRG22, Pan22, PY24, PTG21, PPCC21, PYF20, QH21, QHCB20, RLACE20, RANF20, SKSS22, Sha20m, SYA23, SZ24, SYG24, SAMA21, SK21c, SJB23, SN20, SM21, SZT^{+}23, SU20, SH23, SO20, Tak21, TL21, TA20, Ten20, TZ22]. based [WW22a, WS22, WWW22, XP20, YLHZ20, YWW23b, YY23, ZG21, ZYW22, ZWW21]. baseline [FSLW21]. Basic [Sha20i]. bathtub [BLA20]. bathtub-shaped [BLA20]. Bayes [JHW24, SRU23, SZS21, ZYW23]. Bayesian [Sha21c, Sha23e, ADA22, ALAS21, Ali20, AHB24, AVMS23, AMO20, BLA20, BGV20, CdSL22, CLL21, CA20, CLY20, CLY22, CLY23, DN22, DZS21, DSH24, DK21, FASuA21, FL23, GMCL20, GWNW22, Gho20, Han20, Han23, HLC21, HPW23, IS23b, IS23a, Jan23, JS22a, JZ22a, KEC20, KRK^{+}21, KKLK22, KLdS21, KH23, Kel21, Kim22, KJL23, KB24, KK20, KGK24, LGM20, LGK20, LW21b, LSL21, LLMY20, LWW20, MG23, MB23, MJK20, PG22, Par21, PHW21, PEMCPRC22, PYF20, PMKP23, RMF^{+}21, RT20, RH21, SK21a, Sha23e, SYG24, SCKW21, TK24, TR22, Üns22, WLLH22, WTZ20, YY23, ZW21, ZXDD22, dSPG20, dSdNB22, dSCdC22, dSAF22, Sha23c]. behaviour [vdNPCvB22]. behind [Sha20l]. being [JS22b]. Bell [MAA22]. Ben [Sha20c]. Ben-Naim [Sha20c]. Berlin [Ano23a]. Bernoulli [SAMA21]. Berry [WYBW22]. Best [ZLY23, HCY21]. Beta [DSW20, QMK21, YXC21]. Between [MHKS22, Ali20, AH22, CAA20, EGM20, HCZ^{+}21, KCC22, Mah23, RK22, XSHC20, ZLFW24]. Bézier [SU20]. Bias [BP24b, Bal21, De 20, GCF^{+}20, WFY22, ZSS23]. bias-corrected [Bal21, WFY22]. biased [AKAA22, AÖ21, BAM22, Che20b, KK20, SGY23]. Big [Sha20b]. bilinear [RM21]. billions [Sha20r]. bimetal [ASAP21]. Bin [Sha22b]. BINAR [KCO20, SZ21]. binary [Ano23c, DM22, DDD23, LLX24, MC22, MHB23, SO20, TC21]. binomial [Dab24, DHS24, EN23, FWC24, GMFB21, KSJ24, LWW20, SZ21, VR20, YWW23a, ZXDD22]. Bioconductor [Sha21f]. biological [AP20a, BP21b, TLGW20]. biomarkers [FLPV21]. biomedicine [SZT^{+}23]. Birnbaum [DDL^{+}20, KJB22, LBHX23, LCY22, PTG21, PNN22]. birth [KRK^{+}21]. Bivariate [KSJ24, WS22, AHB24, HA23, JG22, KCO20, KJAK23, LJ23, LBHK22, MD22, Nik22, SZ21, Sul22]. blanket [Nik22]. Blind [ZFLL23]. block [FFNM20, SGY23, XP20]. blockwise [LZCF23]. Boca [Sha20a, Sha20e, Sha21b, Sha21d, Sha21c, Sha22b, Sha22a, Sha23b, Sha23c, Sha23d, Sha23f, Sha23e, Sta21]. Bolin [Sha21d]. bone [LBHX23]. Book [Ano22, Ano23a, Che20a, Sha20a, Sha20c, Sha20b, Sha20d, Sha20f, Sha20h, Sha20g, Sha20e, Sha20i, Sha21b, Sha21d, Sha21c, Sha21e, Sha22b, Sha22a, Sha23b, Sha23c, Sha23a, Sha23d, Sha23f, Sha23e, Sta21, Sha23a, Sha23d]. Books [Sha20i, Sha21h, Sha21i, Sha21j, Sha22b, Sha23e]. boosting [YLS22]. Bootstrap [Cha20, DS20, FSGMM23, YDL23, Bal21, BSH22, De 20, LW21b, Mar21, SC21, SM21]. Bootstrapped [AL20]. Bootstrapping [CS23, Sha21a]. bootstraps [RMN21]. both [RK22]. boundaries [ASP22]. bounded [GAB21, MRNLB20, OKF^{+}21, PTdS^{+}23]. bounds [WYBW22]. Box [FPM22]. break [Cha20]. breaks [CS23]. Broken [Sha21b, Sha21d, Sha23c]. Brownian [Mah23]. BSC [OSG20]. buffered [BHK23]. builder [BBBS23]. Burr [Akg23, MTW21, STG22].

C [Ano21]. calcium [AFAK24]. calculation [HGH22, TTWG20]. Calibrated [ASS20, LWW20]. Calibration [WX21, ASS20, JAZ22, Ozg21, RH21, XCXZ22, ZYFW20, ZC21, ZZY22]. call [JC20]. capability [Bal21, DS20, GL22, PRG22, SJB23]. capture [CB20]. capture-recapture [CB20]. care [Sha20r]. Carlo [BHAK21, ENK^{+}23, GCF^{+}20, OSG20, TÖSY21, YSJL20]. CARR [KJT21]. Case [FPM22, MWS23, NAV23, SCK23, WXWS21]. case-cohort [MWS23, SCK23]. cases [Üns22]. categorical [CHEH21]. Causal [BS22a, Sha23f, KR21, MWS23]. cause [AMO20, MMH20]. causes [DLK24, SAR20, TR22]. CAViaR [Tsi20]. CDF [AAR24]. CDN [Che20a]. cells [GAB21]. cellwise [STM24]. CEM [NF23]. Cengage [Sha21e]. censored [AR21, AMD21, AMO20, Çet21, CLY20, CLY22, CLY23, Dav21, ED22, FASuA21, GSH20, HRB21, HCN20, HMS21, Jia21, JZY24, JRM20, KGK24, Lon22, MWS23, MTW21, MC21, OV21, PMKP23, SA24, STG22, SRU23, SLER21, TG21, TÖSY21, WXWS21, YZZW23, ZHF21]. censoring [AA22a, BA21, BM22, ÇSK22, Che20c, DSSNL20, DLK24, EES24, EN23, FSLW21, KKK23, LTW21, MTSR22, MKK20, NET^{+}24, PB23, SCG21, SÇK24, WXWS21, ZFZ23]. censorship [ADA22, DCMCCB20]. central [DWXW22, Pho23, Won22]. centroid [DHN22]. CERES [Avc24]. Cham [Ano22]. Chambers [Sha20a]. Chan [Sha20d]. Chang [Sha20d]. Change [LLM23, Ano23c, ASB^{+}23, CLY23, EH22, LC20, LK22, LYZ20, LGK20, PK22a, SA24, SK20, SK21c, YWW23b]. change-point [LYZ20, LGK20, PK22a]. changepoint [DM22]. changes [QH21]. Chapman [Sha23b, Sha23c, Sha23d]. characteristic [CMZ24, DZS21]. characteristics [EES24, PS24, TAWL21]. characterization [ANRS20]. chart [ANA^{+}22, ACHS23, ARA^{+}21, AFR23, ASAA23, CLL23, CTPY22, CF21, HSK22, HCZ^{+}21, ISZ^{+}23, JNFTM^{+}21, JZY24, KAN21, MGT21, MLA^{+}20, Mir21, MSR23, uAT20, uAR21, uAS23, PPCC21, PPC^{+}23, QH21, QHCB20, Sab23, SAR20, SAMA21, TL21, THL^{+}22, YLHZ20, YLK^{+}22, YZZW23]. charts [AF20, AAN^{+}23, ANA^{+}23, ASAP21, AAA^{+}22, Ali20, Ano21, CTB20, EH22, HR20, Haq20, HKLA21, HA23, HNK24, HZXS24, JNNTMS24, JP21, KK21, KCC22, LG21, LT23, MB22, TH21, WS22, ZWW21]. Chaudhuri [PS24]. checking [WX21]. Chen [Ano22, Sha23a]. Cheng [Sha20d]. Cheng-Few [Sha20d]. chi [MXYY22]. chi-square [MXYY22]. Chocolate [Sha20b]. choice [GGA22]. Cholesky [KD20]. Chun [Sha20a]. CircSpaceTime [LSM20]. circular [Aça23, APAAC21, LSM20, MVFFC22]. CL1 [Nik20]. claims [MAKRK22]. class [AZS^{+}24, AA22b, BP24a, QL22, SSS^{+}22, vdNPCvB22]. classes [GK20, RK22]. Classical [DZS21, DWWW23, KEC20]. classification [CXH21, GGA22, HD22a, LLX24, MC22, MN21, RANF20, YDL23]. classified [NJR22]. classifier [DHN22, MP21]. classifiers [RANF20]. clinical [CXH21, Sha20j]. closed [KJAK23]. closed-form [KJAK23]. closeness [Dav21]. cluster [BC23, FMH^{+}22, GTMP22, HWF22, VPMA23, ZG21]. cluster-weighted [GTMP22]. clustered [GMFB21, JS24, KHK20]. clustering [AZL21, BVC24, CH21, DK21, GTMP22, KB22, LYZ20, LXC23, Mod22, MT20, MH21, NN22, PDD^{+}21, SG23, WS20, YY20]. clusters [PDD^{+}21, ZCD22]. clusterwise [HFM21, DRG23]. coded [NSK21]. Coefficient [Ano21, AML22, AAA^{+}22, FWC24, FFMZ22, HBKB22, HTX23, HZXS24, HJL22, HZ23, LYP20, San23, SCW21, SBAC21, SL23b, TH21, YDS^{+}20, YLK^{+}22, YZ21, ZLLB21, ZYW23, ZZLL20, ZFZ23]. coefficients [AP20b, RNSR20, SBAC21, YWW23a, YLYD24]. cognitive [LTS^{+}23]. cohort [MWS23, SCK23]. cold [DJA22, LTW^{+}23b]. cold-standby [DJA22, LTW^{+}23b]. coloured [LGM20]. combination [PMW22]. combined [KK21, Mar21]. Combining [HD22b, QAAM22]. common [BYH23, DT23, JG22, JT24, WW22b]. communication [KLL^{+}23]. communication-efficient [KLL^{+}23]. comparative [AAN^{+}23, DDL^{+}20, HD22b, HMS21]. compare [RNSR20, RN21]. Comparing [MHB23, WR22]. Comparison [EGM20, GA20, PA23, PS21, SKH21, SDP24, SZ21, TWN22, AJD21, ASB^{+}23, DLO22, FMH^{+}22, GCF^{+}20, GMJ23, JC20, Jia21, LLP^{+}23, MBM20, SM21]. comparisons [KLC23, YZK22]. competing [AHB24, AMO20, DLK24, Lee22, LTW21, MMH20, MMH22, PY24, PMKP23, SK21a, SK24, SCK23, TG21, TR22, WTZ20]. Complete [DS22, QZ22]. complex [MT23, Pho23]. complexity [Ano23a]. complier [MWS23]. Component [TA20, AALA22, Çet21, CA20, DC23, DT23, GB22b, LLM23, NAJ21, XYZ20, ZRB^{+}20, ZLLZ22]. components [BM23, JHW24, Sir20]. Composite [GH23, ZFZ23, Li20, Won22]. Compositional [LTY^{+}24, ISZ^{+}23]. Compound [Dab24, MAKRK22]. comprehensive [DLO22]. Compromise [HAV20, KAQ21]. Computation [LBHK22, SCKW21]. Computational [BBBS23, RN21, DGMN22, KJL23]. Computationally [MJG21]. Computing [PMW22]. concerning [Li24]. concomitants [MB22]. condition [ZXG20]. Conditional [YLS22, AZL21, Chr22, KCC22, PHW21, RLACE20, WS22, ZY21]. conditionally [KL22b]. conditions [WWW23]. Confidence [BM22, BSH22, LW21a, NAV23, PNN22, SN20, TR24, Bal21, Cha20, DS20, DZS21, KLC23, Li21, LRN22, LLMY20, MKK20, San23, WGF22, ZXZ^{+}22]. confirmatory [LWS20]. conformance [LW21a]. confounder [BS22a]. confounding [Sha22b]. conjoint [Sul22]. conjugate [DN22, PHW21, SZS21]. consecutive [DK22]. considering [JC20, LTS^{+}23, Sab23]. consistency [WWW23, ZYW22]. consistent [GFPG22, JF20, ZYW23]. constant [SLER21, VR20]. constant-stress [SLER21]. Constrained [JS22a]. constraint [BA21]. constraints [ASS20, JAZ22, SYVV20, ZW21]. Construction [ACHS23, KC22]. consumer [JB22]. containing [SN20]. contaminated [ASOH23, MG23]. contamination [STM24]. Content [IAO^{+}23]. context [LYZ20, OSG20]. contingency [Sul21]. continuous [LZZ22, PTdS^{+}23, RK22, TH22]. Control [LT23, AF20, ANA^{+}22, ACHS23, AAA^{+}22, Ano21, ASAA23, Asa22, CTPY22, CF21, HZXS24, ISZ^{+}23, JNFTM^{+}21, JF20, JZY24, JP21, KK21, KAN21, LG21, MGT21, MLA^{+}20, MSR23, uAT20, uAR21, QHCB20, Sab23, TL21, TH21, THL^{+}22, YLHZ20, ZLFW24]. controlled [LL21]. controlling [PP24]. conventional [BAAR21]. convergence [DS22, Li24]. convergent [SYA23]. convolutional [CXH21]. Conway [Abo23, GMFB21, dSMA22, SAA^{+}22, SBA23]. Coordinate [WZ21]. Copula [BHAK21, ALAS21, ES22, KHK20, Mah23, Nik22, SGY23, WS22]. copula-based [WS22]. Copula-frailty [BHAK21]. copulas [GH23]. corrected [Bal21, WFY22]. Correction [Ano21, Ano23b, Ano23c, De 20, HAC20, Li24, MLA^{+}20]. correlated [ASW24, GT22, KKS20, Mah23, MSH20, Sta21, ZLL24]. Correlation [CXH21, FSGMM23, LXH20, XSHC20, XLF22, ZYW23]. Correlation-driven [CXH21]. Correlations [MD22, LT23]. Cosine [JIK22]. Cosine-based [JIK22]. cost [BA21, VPMA23]. costs [FNY21]. count [Ano23c, ACS21b, DM22, GMFB21, HMB22, Jan23, MSBS22, MSR23, SR23, SW20, WWWZ21, dSLP^{+}23]. counterpart [RMM24]. counting [PA23]. counts [AJ21, LKS21, LK23]. Covariance [SLM^{+}22, AA23, BM23, WLS^{+}21]. covariate [Che20b]. covariates [Che20c, DXY21, HJL22, KR21, LE21, MN21, RNSR20, TWS23, XLF22, YXYL22]. coverage [DHS24]. COVID [SOBC23]. COVID-19 [SOBC23]. Cox [AR21, AP20b, DXF^{+}23, KAQ23, SZ24, YXYL22, ZSWN21]. Cox-optimized [SZ24]. Cramér [AFM22, TTWG20]. CRC [Sha20a, Sha20e, Sha21b, Sha21d, Sha21c, Sha22b, Sha22a, Sha23b, Sha23c, Sha23d, Sha23f, Sha23e, Sta21]. credible [LWW20]. crime [Sha20l]. criteria [BP21b, CHW21, Nik20, RANF20, SM21, vdNPCvB22]. criteria-based [CHW21]. criterion [LC20, LTW23a, PDD^{+}21]. critical [Ber22]. cross [BGV20, NAV23]. cross-validation [BGV20]. crossover [NSK21]. Cryptocurrency [CHG21]. crystals [AFAK24]. CT [Sha21e]. Cucconi [NM20]. cum [IMAI20, KKZ23]. Cummings [Sha21b]. cumulant [CCL^{+}20]. Cumulative [DCMCCB20, NS22, PYF20, TR22]. Cumulative/dynamic [DCMCCB20]. curable [WTZ20]. cure [WTZ20]. cured [WXWS21]. current [AML22, YDS^{+}20]. curve [DCMCCB20, NET^{+}24, SU20, vdNPCvB22]. CUSCORE [HNK24]. CUSUM [Ano21, Ali20, ASAA23, De 20, EH22, HR20, HKLA21, HA23, KAN21, Lee20, QH21, TH21, WS22]. cvmgof [AFM22]. CVT [IAO^{+}23]. cyclostationary [PFMT20].

D [Ano23a, Sha23a, SYA23]. DAG [ZTP21]. dam [Akg23]. Daniel [Sha23b]. Danish [WL21]. Data [BB22, Aça23, AR21, ACHS23, AMD21, AML22, AJN21, AFAK24, Ano23c, AVMS23, ACS21b, AMO20, BHA22, BS21, BHAK21, BAM22, BPT23, BP24b, CIPCR23, Cha23, CB20, CLY20, Che20b, CLY22, CLY23, CTY23, CC21, Chr22, CF21, DHN22, Dav21, DM22, DSSNL20, DDD23, DF20, DK21, DLK24, FFMZ22, FFS23, GW22, GSH20, GMFB21, GSRY23, GA20, GMJ23, HGH22, HMB22, HRB21, HD22a, HCN20, HH21, HMS21, HWF22, ISZ^{+}23, Jan23, Jia21, JZY24, JS24, KRK^{+}21, KM22a, KKS20, Kim22, KK20, KCC22, KHK20, LLWZ20, LCZ^{+}20, LSM20, Lee22, Li21, LRN22, LXC23, LZCF23, LWWD20, LCH20, LBHX23, LTY^{+}24, Lon22, LXH20, MXYY22, MWS23, MTW21, MWMCR20, MAKRK22, MMH20, MMH22, MT20, MAY23, MSR23, NS22, NN22, OKF^{+}21, OV21, PG22, Pan22, PY24, PDD^{+}21, PHW21, PTdS^{+}23, PYF20, PB23, PMKP23, QZ22]. data [QH21, RLACE20, RN21, SA24, SK21a, SRG22, SR23, SG23, Sha21a, Sha20n, Sha20o, Sha20q, Sha21c, Sha21k, Sha22a, Sha23c, Sha23e, SOBC23, SRU23, SLER21, SBB21, SCK23, Sta21, SN20, SW20, SDML20, SL23b, SO20, TAWL21, TA20, TWN22, TR24, TG21, TÖSY21, TR22, TH22, VR20, WWWZ21, WXWS21, WKS23, WR22, WL21, XSHC20, YKM23, YDS^{+}20, YZZW23, YH20, ZCD22, ZRB^{+}20, ZHF21, Zha21a, ZXDD22, ZLL24, ZLLZ22, dSdNB22, dSLP^{+}23, Sha20i, Sha20a, Sha20b, Ano23a, Sha20b]. data-driven [KM22a]. datasets [CLL21, Fer23, KH20, LLX24, MLG23]. date [Cha20]. David [Sha20i]. Dawoud [DA21]. deal [VR20]. decision [BLZ21]. decomposition [KD20]. decrease [WW22b]. decreasing [BT24, GM22]. Deep [TZ22, CHG21, KV23]. deferred [KJB22]. Defining [STÖ21]. deflation [XYZ20]. degradation [HLC21]. delayed [JNFTM^{+}21]. delta [AH22]. demand [FPM22]. Density [LA20, BAM22, CH21, GJQ22, GFPG22, HPP23, JIK22, KM22a, LJ23, MC22, Nis23, NC23, NJ20, SGY23, SK21c, Ten20]. Dependence [BAM22, CB20, JIK22]. dependencies [dSAF22]. dependency [ES22]. dependent [AZS^{+}24, AHB24, CM21, DLK24, GSH20, GWL21, MMH20, SK21a, SW20, TK24, WWWZ21]. depth [BS21, HM23, Pan22]. depth-based [Pan22]. derivation [Li24]. derivatives [Ber22]. descent [WZ21]. descriptive [Sha21a]. Design [uAS23, BLZ21, BS22b, CF21, Das22, DSM22, EGM20, JNNTMS24, KK21, TWS23, VPMA23, WLZ21, Won22, ZWW21]. designed [BP21b]. Designing [CLY20, CLY22, CLY23, JB22, ANA^{+}23, ISZ^{+}23, MJSB23, SMZ20]. designs [DGMN22, Kel21]. Detecting [JP21, EH22, HNK24, LWS^{+}23, QH21]. Detection [GB22b, AFAK24, Ano23c, DM22, Fer23, FMH^{+}22, HLK24, HW20, LYZ20, PK22a, SK20, TK22, WW22b, YZZW23]. Determination [KJB22, SJB23]. Developing [PRG22, AR21]. deviations [LLBW23]. device [ZSS23]. diagnosing [EH22]. Diagnostic [GB22a, CIPCR23, RNSR20, RN21, TLGW20, vdNPCvB22]. diagnostics [KAQ23, OCNN20]. Dice [Sha20i]. difference [BSH22, GT22, HGH22, JG22, PNN22, SB20]. differences [MKK20, YKM23]. different [AFAK24, AAR24, BM23, DS20, ES21, Han20, Han23, IS23b, JAZ22, Jia21, PY24, TdL22]. differential [KN22, LCH20, SYA23, ZLFW24, ZQAH23]. digests [Ros22]. dimension [CLL21, Chr22, DC23, KB22, LZ20, ZZLL20]. dimensional [BHA22, CHJS20, Che22b, CTY23, CC24, CC21, CG21, DK21, GJP24, HD22a, HTF21, Kaz24, KKS20, LLWZ20, LCZ^{+}20, Li20, LZZ22, LZCF23, LLZ20, LTS^{+}23, MXYY22, NS22, PZWL23, PEMCPRC22, QZ22, RT20, WLS^{+}21, YH20, YLS22, ZLZ24]. Ding [Ano22]. Ding-Geng [Ano22]. Dinov [Ano23a]. direct [NAV23, Ten20]. direct-inverse [NAV23]. direction [CHG21]. directional [PG22, Pan22]. Dirichlet [VS20]. Discover [Sha20c]. discovery [PP24, ZLFW24]. Discrete [TdL22, FLPV21, GPH23, Lee22, RNSR20, ZTP21]. discriminant [CHJS20, Che22a, GvRS23, WWW22]. Discriminating [RK22]. disease [CXH21, CHJS20]. dispersed [AJ21, KSJ24, KH20, Li21]. dispersion [LT23, MGT21, PPC^{+}23, VR20]. distance [DJA22, KJT21, LS21, LW23, LA20, Mod22]. distance-based [Mod22]. distances [HFM21, NN22]. distortion [ZYFW20, ZXG20, ZC21, ZY22, ZZY22]. distortions [Zha22]. distributed [DZS21]. distribution [AZIE21, ADA22, Akg23, AMO22, AMD21, ASP22, AJD21, AHB24, AAR24, Bal21, BLA20, BM22, BAAR21, BCM21, CA20, CLUX21, Dav21, DPS21, DLO22, DA23, DSH24, DLK24, ESA23, GvRS23, Han20, HCY21, IS23a, JG22, Jia21, JZY24, JRM20, KJB22, KB20, KLdS21, KZ21, KJAK23, KCC22, KKK23, KGK24, LG21, Lon22, LCY22, MTW21, MTSR22, MAKRK22, MJGAF21, MC21, NSK21, NS22, NC23, NJ20, OKF^{+}21, PTG21, Pho23, PYF20, PMW22, PB21, RLACE20, RMF^{+}21, SSN20, SCG21, STG22, SYVV20, SK21b, SBH21, San23, Sha21a, SYG24, SLER21, SMZ20, SZS21, SU20, SH23, TdL22, TL21, TG21, THL^{+}22, YGWH21, YY23, ZXDD22, dSdNB22]. distribution-based [TL21]. distribution-free [SMZ20]. distributions [BLZ21, BT24, CdSL22, CGGE20, FASuA21, GM22, HGH22, HPP23, KC22, KH20, Kur24, LTW21, MKK20, MLA^{+}20, MD22, MJK20, PHW21, PK22b, PNN22, WFY22, XG22, ZY21, dSPG20]. divergence [KN22, KL24, LJ23, MP20, RANF20, SK21c]. divergence-based [RANF20, SK21c]. diverging [XLF22]. diverse [TWZ23]. do [MGA21, Sha20i]. domain [Zha21a]. domains [ZTP21]. Double [Sha21a, ARA^{+}21, CP22, EGM20, KK21, Lon22, MK20, uAT20, SAR20, SJB23, TR24]. doubly [HLC21, JS22a]. drifts [KSL22]. driven [BB22, CXH21, KM22a]. DTRreg [SMWP20]. Dual [Tac24, CM21, HNK24]. Dual-rank [Tac24]. dual-record [CM21]. due [SSS20]. DWSurv [SMWP20]. Dynamic [dSdNB22, DCMCCB20, SMWP20, SMZ20]. dynamics [KSL22].

E-Bayesian [Han20, Han23, IS23b, IS23a, PYF20, SYG24]. E-MSE [Han20].

E-MSEs [IS23a]. E-Posterior [IS23a, Han23]. early [BLZ21, TK22]. ecological [Sha20e, Sha20n]. Economic [CF21, EGM20, KK21]. Economic-statistical [CF21]. ed [Sta21]. Edgeworth [GvRS23]. Edgeworth-type [GvRS23]. edited [Ano22, Sha20a, Sha20e, Sha23a]. effect [MGA21, RCPA21, RMM24, SAB21, SSS20, TMMCS21, WQZS23, WR22, YK21, ZLY23]. effective [BP24a, HGH22]. Effects [JAZ22, AL20, BB22, Cha23, GWNW22, HVAVD20, MWS23, SO20, TAWL21, TH22, YDL23, YY23, Zha21b, ZXDD22]. effects-based [SO20]. efficiency [AP20b, ARA^{+}21, RLACE20]. Efficient [AML22, DGMN22, HZ23, Li20, SSS^{+}22, WWWZ21, YDS^{+}20, ZH22, ANA^{+}23, ASAA23, CLY23, DPS21, HPV23, ISZ^{+}23, KLL^{+}23, KJAK23, LTW23a, LE21, MJG21, SS22, SC21, SMZ20, WB22, WWW22]. efficiently [ZFLL23]. elastic [BPT23, JS22a, MB23]. electrodes [TG21]. ellipses [LBHK22]. elliptical [Ema20, LTL20, ZY21]. emerging [TK22]. emphasis [BSH22]. Empirical [LRN22, QZ22, BL20a, CTY23, DLO22, HTF21, JZ22a, JZ22b, KKS20, LLP^{+}23, LLX24, MB23, PP24, SN20, SZS21, SL23b, YWW23b]. Empirical-likelihood-based [LRN22, YWW23b]. endpoints [SMWP20]. energy [LTW23a]. Enhanced [HKLA21, CLL23, GAN^{+}20, KKZ23]. Enhancing [uAT20]. ensemble [BHA22, MP21, WW22a]. ensembling [TWZ23]. entropy [CH21, DSH24, HRB21, NJ20]. entropy-based [CH21]. enumeration [vdNPCvB22]. envelope [ZH22]. environmental [Fer23, Sha20e]. epochs [MRNLB20]. equality [BLZ21, CGGE20]. equations [LCH20, Nik20, SYA23, ZQAH23]. equivalence [CMZ24, dSCdC22]. error [CdSL22, Che20b, Che20c, Ema20, GB22a, HAC20, JF20, JZ22a, Kel21, KK22, KKZ23, LE21, LWG20, uAR21, SYVV20, Sir20, YK20, dSPG20]. error-prone [LE21]. Errors [ASW24, ASP22, CIPCR23, CPX22, DWWW23, HVAVD20, HAC20, HJL22, IS23a, KSAR21, Mah23, Sab23, SMGAV23, SBB21, TMMCS21, YK21, YZ21, ZYFW20, ZXG20, ZC21, ZKG21, Zha22, ZY22, ZYW22, ZZY22]. Errors-In-Variables [ASW24, YK21]. Esséen [WYBW22]. essentials [Sha20t, Sha20d]. estimable [Li21]. Estimate [DT23, WLLH22, ZG21]. estimated [ASAP21, AP20b, MSMTK22, TMMCS21]. estimates [JHW24, LWG20, RMN21]. Estimating [ASB^{+}23, LCP23, ASP22, AI23, AAR24, EES24, FSB20, KLdS21, Nik20]. Estimation [AE23, Akg23, Bal21, BS22b, BT24, GWL21, HAC20, JRM20, KL22a, KGSA22, LJW23, Lon22, MWS23, MN23, MRNLB20, MC21, NFP21, PTG21, SSN20, SCG21, SS22, SYVV20, SÇK24, XG22, ZLL24, AC20, AZIE21, AJ21, AML22, AZS^{+}24, AJD21, AB23, AKA22, AÖ21, BBBS23, BLA20, BAM22, BPT23, Ber22, BB22, BP21a, BP24a, BCM21, BSH22, CIPCR23, Çet21, CCL^{+}20, CB20, CM21, Che20b, Che20c, De 20, DK22, DS20, DZS21, DCMCCB20, DHS24, ES21, ESA23, Ema20, FZMZ20, FFMZ22, FH22, GJQ22, GT22, GFPG22, GSRY23, GCF^{+}20, GH23, GAN^{+}20, Han20, Han23, HRB21, HPP23, HPV23, HDP21, HPW23, HZ23, IS23a, JG22, JIK22, KD20, KIS20, KL24, KK20, KK22, KKZ23, KGK24, LJ23, LWS20, LW23, Li24, LLX24, LXH20, LW22, LA20, MGB21, MAA22, MW24, MHB23, Nis23, PS21]. estimation [QAAM22, RLACE20, STG22, SSS^{+}22, SRG22, San23, SKH21, Sha20p, SL23a, SGY23, SOBC23, SAB21, Sto22, SKO20, SU20, SH23, SO20, TdL22, TA20, Ten20, TÖSY21, TR22, VS20, WWWZ21, WQZS23, WYBW22, WW22b, WS20, XP20, XCXZ22, YDS^{+}20, YK20, YZ21, YLS22, YXYL22, ZKG21]. estimations [HYWW21, PYF20, SYG24]. estimator [Abo23, Aça23, AFS21, AAQ20, AALA22, AGY21, BL20a, BSH22, DA21, DPS21, ED22, GB22a, HCY21, KM22a, KKZ23, LAW20, LE21, NN23a, Ozg21, QAAM22, RMM24, SAA^{+}22, Sir20, STÖ21, WFY22, YK21, ZLLZ22]. estimators [ASOH23, AKAA22, ASS20, Bal21, BAAR21, BP24a, CMMP20, DWWW23, Fer23, HWF22, IMAI20, JAZ22, KJT21, KZ21, KJAK23, KK20, LW21b, LLP^{+}23, MBM20, MP20, MGA21, uA20, NJ20, NN23b, PB23, QMK21, SSS^{+}22, SS22, SLM^{+}22, SBA23, SKSS22, SZS21, Üns22, WKS23, WWW23, ZW21, ZYW22]. ethical [Das22]. Euclidean [NN22]. Euler [ZQAH23]. euro [Cha23]. EuroBarometer [Cha23]. European [FNY21, JC20]. EV [WWW23]. Evaluating [CHEH21, Mah23, SAB21]. Evaluation [FPM22, MJSB23, ATH20, DDL^{+}20, IS23a]. evasive [ZSS23]. event [BHAK21, BSH22, CC21, KHK20, LYZ20, ZHF21]. Events [MHKS22, Ali20, HCZ^{+}21, KCC22, LGK20, SBH21]. every [Sha20c]. evidence [STÖ21]. EWMA [Ali20, EH22, HKLA21, HSK22, HCZ^{+}21, JZY24, PPCC21, PPC^{+}23, TL21, YZZW23, ZWW21]. EWMA-type [PPCC21]. Exact [ÇSK22, CLUX21, DM22, LCY22, KCC22, PK22b, PPCC21, PPC^{+}23, PMW22, SBH21, Ano23c]. examples [Sha22a]. Excel [Sha20d]. excellent [Sha21h, Sha21i, Sha22b]. excess [HMB22]. exciting [YLYD24]. expansion [GvRS23]. expected [IS23a, MRNLB20, WYBW22]. expectile [CZ23, LZZ22, PZWL23]. experiment [SK21b]. explanatory [QL22, SZ21]. Exploring [DC23]. exponent [BL20a]. exponential [AMO22, AA22b, BM22, ÇSK22, Dav21, GSRY23, JZY24, KL24, KCC22, KKK23, LKS21, LT23, MKK20, MAKRK22, MC21, uAT20, PYF20, SSN20, SS22, SRG22, SKSS22, San23, SKO20, TG21, ZY20, Zha22, ZZY22]. exponential-polynomial [KL24]. exponential-type [SS22]. exponentiality [ANRS20]. Exponentially [AAA^{+}22, HZXS24, ISZ^{+}23, Mir21, MSR23, uAR21, SMZ20]. exponentiated [AMO22, CA20, EES24, LTW21]. expression [ASAP21]. extended [ZY20]. extending [AH22]. extension [BYH23, OV21]. extensions [GH23]. extra [MSBS22]. Extraction [LG21]. extrapolation [LLMY20]. extreme [CHEH21, DDD23, LWS^{+}23, OCNN20, THL^{+}22, WFY22, dSdNB22]. extreme- [CHEH21]. extreme-value [OCNN20]. extropy [HRB21, JS22b].

factor [Che22b, ZYW23]. factorization [ASB^{+}23]. factors [Won22]. failure [AMO20, BKM21, CC21, DLK24, FH22, GM22, Han23, KKK23, MWS23, SA24, SCG21, STG22, Sha21k, TA20, WXWS21]. failure-time [TA20]. failures [FASuA21]. false [PP24, TR24, ZLFW24]. false-positive [TR24]. families [SKO20]. family [AKA22, BA21, DK22, FFNM20, JF20, KCO20, LKS21, LTW21, MB22, MK20, PK22b, SA24, ZY20]. family-wise [JF20]. Fast [CC24, RH21, XYZ20, CC21, PP24, SMZ20, ZSWN21]. fat [HPP23]. fat-tailed [HPP23]. fault [KKLL24]. FDA [RLACE20]. Feature [LLWZ20, HD22a, Kaz24, LCZ^{+}20, LLZ20, MXYY22, MH21, SDML20, WWW22, XSHC20, ZLZ24]. features [CZ23]. feed [Tak21]. feed-forward [Tak21]. Few [Sha20d]. fiducial [GYL24, YGWH21]. filter [ZY20]. filtering [BYH23, MN21]. finance [PB21]. Financial [Sha23b, BHK23, LWWD20, MWMCR20, Sha20d]. Finch [Sha21d]. Finite [PB23, CdSL22, DK21, Nik22, SCKW21]. Finite-sample [PB23]. fire [WL21]. First [Sha21e, YWW23a, Ber22, KKK23, SCG21, STG22, WS22]. first-failure [STG22]. First-order [YWW23a, WS22]. Firth [AR21]. fit [ACS21a, AFM22, BKM21, DLO22, LS21, MVFFC22, NN23a, OKF^{+}21, PS21, SU20, vdNPCvB22]. fitness [DJA22]. fitness-distance [DJA22]. fixed [BB22, TR24, WR22]. fixed-inverse [TR24]. FL [Sha20e, Sha21b, Sha21d, Sha21c, Sha22b, Sha22a, Sha23b, Sha23c, Sha23d, Sha23f, Sha23e, Sta21]. flash [Cha23]. Flexible [YY23, ACS21b, HAV20, LWS20, SOBC23, dSPG20]. Floor [Sha21e]. folding [LZ20]. follow [SKSS22]. follow-up [SKSS22]. forecasting [CHG21, FPM22, SOBC23, Tak21]. forensic [FSB20]. forest [WW21]. forests [GMJ23]. form [CLUX21, KJAK23]. forward [Tak21]. four [SKSS22, Sha21h]. four-phase [SKSS22]. fourth [De 20]. fractal [DJA22]. fractional [Mah23, ZQAH23]. frailties [JLZ21]. frailty [Asa22, BHAK21, Dab24, JLZ21, KAN21, LBHX23]. frailty-based [KAN21]. framework [AR21, CXH21, MLG23, WW22a, ZG21]. fraud [Sha20r]. free [HSK22, HNK24, LCZ^{+}20, MXYY22, SMZ20]. frequency [BGGH21]. frequentist [AP20a]. Fu [Sha22a]. Fuentes [Sha20e]. full [PHW21]. full-conditional [PHW21]. fully [PHW21]. function [AZS^{+}24, AJD21, BKM21, BAAR21, BGGH21, CMZ24, ESA23, FSLW21, FSGMM23, GFPG22, GvRS23, KM22a, MN23, MW24, Pho23, SA24, SZS21, TR22, VPMA23]. function-on-scalar [MW24]. Functional [LXC23, ZLLZ22, AC20, ASW24, DHN22, FZMZ20, HH21, PRG22, SRG22, Sha21a, SBAC21, ZCD22, ZFZ22, ZW23]. functional/longitudinal [DHN22]. functions [AFAK24, BM22, CH21, Han20, Han23, HYWW21, IS23b, LLWZ20, LG21, Li21, MC22, NS22, NJ20, NFP21, Par21, Ros22, Tak21]. fundamental [MC22]. fused [LC20]. future [BGV20]. Fuzzy [AZL21, ACHS23, HAV20, KAQ21, Tak21].

G [Sha23a]. gambler [Hus22]. game [Sha20k]. Gamma [SP21, AKAA22, Asa22, GT22, HGH22, IS23b, KGSA22, Kur24, LSL21, MGB21, MBM20, NC23, PTdS^{+}23, QAAM22, RMF^{+}21, SRU23, SZS21, FL23, PG22, VLRZ21, VPMA23]. gamma-frailty [Asa22]. Gamma-related [SP21]. Gan [Sha22a]. GARCH [SK20]. Gaussian [AAQ20, AALA22, ALAS21, BP24b, CvNR23, HLC21, KB20, LGM20, LTY^{+}24, MN23, NS22, RT20, ZTP21, ZLFW24, ZLLZ22]. GBM [FNY21]. GEE [XLF22]. Gegenbauer [HPW23]. Gelfand [Sha20e]. gene [ASAP21]. general [AZIE21, FSLW21, Gra20, KL24, LTW21, MLG23, PHW21, ZXG20]. Generalized [AJ21, HPV23, HTX23, MLA^{+}20, SKSS22, YGWH21, AJD21, Bal21, BBSM23, CP22, ÇSK22, CTY23, DPS21, DXF^{+}23, DDD23, EGA21, GT22, GYL24, HA23, HW20, HZD23, KNM23, KH20, KKK23, LTL20, Li21, LWWD20, LTW21, MLG23, MHB23, NM20, OCNN20, PMW22, RM21, RMF^{+}21, RCPA21, SSN20, SCG21, SYG24, SRU23, SLER21, TK24, VLRZ21, YY23, Zha21b, ZFZ22, dSLP^{+}23]. generally [CLL23]. generated [GA20]. generating [Sul22]. genetic [NSK21, NI23, Nis23]. Geng [Ano22]. geodesic [PDD^{+}21]. geometric [AB23, KGSA22, KH20, MJGAF21, RK22, Sto22]. geostatistics [Sha20m]. Germany [Ano23a]. Gibbs [KH23, ZTP21]. given [MRNLB20]. GLD [Pan22]. GLD-plot [Pan22]. Global [CG21, Sha20b, WLS^{+}21, LTW23a, Sha20m, ZFLL23]. GmbH [Ano23a]. goals [HAV20]. God [Sha20i]. Goodness [MVFFC22, ACS21a, AFM22, BKM21, DLO22, LS21, NN23a, PS21, SU20]. Goodness-of-fit [MVFFC22, ACS21a, AFM22, DLO22, NN23a, SU20]. GQL [SO20]. graded [FL23]. gradient [BBSM23, KSL22]. graph [CXH21]. Graphical [AZ20, AE23, AP20a, LGM20, PZWL23, Pan22, ZLFW24]. group [AZ20, CHJS20, JB22, KJB22, YZK22]. Grouped [HD22a]. grouping [ZLY23]. groups [BVC24, ES22]. GrowNet [SZ24]. Gruyter [Ano23a]. Guangquan [Sha21c]. Gumbel [MTSR22].

H [Ano21]. Hachette [Sha20i]. Haenszel [BSH22]. Haining [Sha21c]. half [SLER21]. half-normal [SLER21]. Hall [Sha23b, Sha23c, Sha23d]. Hamiltonian [YSJL20]. Handbook [Sha20f, Sha20h, Sha20g, Sha20e]. handling [DA21]. handwriting [FSB20]. hardback [Che20a, Sha23b, Sha23c, Sha23a, Sha23d, Sta21]. hardcover [Ano22, Ano23a, Sha20a, Sha20b, Sha20e, Sha20i, Sha21b, Sha21c, Sha22b, Sha22a, Sha23f, Sha23e]. Hardy [Sha23b]. hazard [AR21, DK22, FSLW21, HCN20, NET^{+}24, PYF20]. hazards [BA21, Che20c, FSLW21, KAQ23, WTZ20, YDS^{+}20, ZLY23]. HDBRR [PEMCPRC22]. health [Ano22, DDD23, Sha20m, Sha20r, Sha21j, Sha23a]. health/medical [Sha21j]. healthcare [EGA21, MSR23, SAMA21]. heavily [AR21]. Heavy [XP20, BPT23, MWMCR20]. heavy-tailed [BPT23, MWMCR20]. Hellinger [LW23]. Hermite [Ten20]. Hershey [Sha20b]. Heston [JC20]. heterogeneity [AL20, LZZ22]. heterogeneous [Asa22, KB24]. heterogeneously [HYWW21]. heteroscedastic [KL22b, LK22, YDL23, YZ21, ZFZ23]. Heteroscedasticity [WKL^{+}24, SCW21]. heteroskedastic [BBSM23, PHW21]. heteroskedasticity [CLUX21]. Heuchenne [Ano21]. heuristic [KV23].

Hidden [SW20, BBBS23, Gho20, Tak21, VS20]. hierarchical [GH23, IS23b, NN22, SYG24]. High [PZWL23, BHA22, CHJS20, Che22b, CTY23, CC24, CC21, CG21, DSSNL20, DK21, GJP24, HD22a, HTF21, KKS20, Li20, LZZ22, LZCF23, NS22, PEMCPRC22, QZ22, RT20, WLS^{+}21, YH20, YLS22, ZLZ24]. High-dimensional [PZWL23, BHA22, CHJS20, Che22b, CTY23, CC24, CC21, CG21, DK21, HTF21, KKS20, Li20, LZZ22, LZCF23, NS22, PEMCPRC22, QZ22, RT20, WLS^{+}21, YH20, YLS22]. higher [CCL^{+}20, KL22a]. higher-order [CCL^{+}20, KL22a]. highly [Fer23, YZZW23]. Hoboken [Che20a]. Hoeting [Sha20e]. Holmes [Sha21d]. homogeneity [Ana20, PK22b, SBH21]. homogeneous [KSJ24, MRNLB20]. homogeneously [ASAA23, TMMCS21]. honest [GYL24]. Hotelling [KKLL24, SAR20]. Huber [HLK24]. Huberized [KH23]. hybrid [AMO20, ÇSK22, Dav21, DLK24, EN23, GT22, LTW21, Lon22, MTSR22, WB22, XQ21]. hypergeometric [YXC21]. Hypervolume [FLPV21]. hypotheses [CMZ24, FFNM20]. Hypothesis [KB20, TAVP20, BS22b, GM22, QL20, ZW23].

Ian [Sha20i]. identical [AKA22, Çet21]. identifiability [ZXG20]. Identification [KSAR21, GB22b, WKL^{+}24, YH20]. identify [KKLL24]. IGI [Sha20b]. ignorable [AMO20, GSRY23, SRG22]. II [AFR23, BA21, BM22, Çet21, EES24, EN23, GSH20, HRB21, JZY24, JRM20, Kel21, KCC22, MTW21, MTSR22, MSMTK22, MKK20, MC21, NET^{+}24, TG21, YLHZ20]. imbalanced [LLX24]. Imperfect [MHKS22, JNNTMS24]. implementation [Sha23e]. Implications [LLX24]. importance [LL21]. imprint [Sha20b]. improve [NI23]. Improved [BBSM23, DHS24, Ema20, Fer23, MGB21, ASAA23, CLUX21, KD20, LZZ22, Mir21, Ozg21]. improvement [ZSS23]. Improving [KÖ21, WW21]. Imputation [Won22, LZCF23, MMH22, RN21, SBB21]. impute [KR21]. INAR [FWC24, HZD23, KSJ24, LLBW23, MSBS22, QL22, Sto22, YWW23b]. Incidence [Sha21b, TR22, Üns22]. included [SYA23]. incomplete [MN21]. incorporating [PZWL23]. Increasing [ARA^{+}21]. independence [ASM20, CAA20, MJG21, QZ22, RDSG21, Sul21, SH23]. independent [GB22b, HD22b, LLM23]. index [Bal21, DS20, DZS21, HJL22, HDP21, Li20, MXYY22, MK20, PB23, SJB23, SBAC21, WL21, XP20, ZZLL20]. indicators [ZFZ23]. indices [KSJ24, PRG22]. Indirect [CCM20, Par21]. individual [HBB21, WR22]. individuals [JP21]. induced [SCK23]. induced-smoothing [SCK23]. industry [FPM22]. inefficiency [Üns22]. inequality [GM22]. Inference [AMD21, AA22a, KKK23, LTW21, SMGAV23, SLER21, TG21, AP20a, AMO22, ATH20, AVMS23, AL20, BS22a, BVC24, ÇSK22, CvNR23, CCM20, CLUX21, DXF^{+}23, FL23, GT22, Gho20, GYL24, HJL22, KEC20, LLWZ20, LC20, LKS21, LK23, LLMY20, LLBW23, LBHX23, MTW21, MTSR22, MK20, MG23, MJK20, NET^{+}24, Par21, PK22a, QL20, RT20, SK21a, SC21, Sha23f, SL23b, YGWH21, YDL23, Sha23f]. inferences [ATH20, DLK24, LCY22]. inferential [Ano23a]. infinite [WW23]. inflated [ACS21b, DDD23, GMCL20, GMFB21, LKS21, LK23, LE21, MSBS22, MAY23, PB21, SR23, SZT^{+}23, ZXZ^{+}22]. Influence [KAQ23, OCNN20, AP20b, Kel21, KK22, TC21, WS20]. Influential [AFAK24, GSRY23, SRG22]. Information [RANF20, Sha20b, ANA^{+}23, AZL21, CHW21, CLL23, CTPY22, HW20, LC20, LXC23, Nik20, NN23a, THL^{+}22]. informative [CA20, WXWS21]. INGARCH [LKS21, LK23]. inhomogeneous [FSGMM23]. initial [SMZ20]. INLA [CvNR23]. INMA [AK23]. innovations [AZS^{+}24, CCM20, GMCL20, KCO20, KSJ24, KL22b, SZ21]. Innovative [Sha20j, MSH20, ZSS23]. inspecting [QHCB20]. instrumental [YLS22]. insurance [MAKRK22]. Integer [RM21, AK23, BL20b, GMCL20, KCO20, KL24, LJ23, SOBC23, YLYD24]. Integer-valued [RM21, AK23, BL20b, GMCL20, KCO20, KL24, LJ23, SOBC23, YLYD24]. Integrated [Sha20a, BBC21, JNFTM^{+}21]. integration [CPX22]. integrative [CHJS20]. integro [ZQAH23]. integro-differential [ZQAH23]. intelligence [Sha23f]. intercept [TLGW20]. interpoint [Mod22]. Interval [Ano21, GT22, JG22, AMD21, BSH22, DCMCCB20, DHS24, HMS21, HZXS24, KK21, KLC23, MWS23, San23, SLER21, TH21, WXWS21, XQ21, YLK^{+}22]. interval-censored [HMS21, MWS23, WXWS21]. interval-valued [XQ21]. intervals [Bal21, Cha20, DS20, DZS21, HSK22, KC22, LW21a, Li21, LRN22, LLP^{+}23, LWW20, MKK20, NAV23, PNN22, SN20, TR24, ZXZ^{+}22]. intractable [Par21]. intrinsic [KB22]. introducing [SMWP20]. Introduction [Sha23d, Sha22b, Sha23d]. Inverse [AAQ20, AAR24, AALA22, AJD21, Bal21, CLL21, HLC21, IS23b, KB20, LZ20, LTY^{+}24, MGB21, NAV23, RMF^{+}21, SAB21, SZS21, TR24]. inverted [KKK23, LTW21, YXC21]. investigate [Pan22]. investigation [HTF21]. Invited [DDL^{+}20]. irregularly [Zha21a]. IRT [TC21]. ISBN [Ano22, Ano23a, Che20a, Sha20a, Sha20c, Sha20b, Sha20d, Sha20e, Sha20i, Sha21b, Sha21d, Sha21c, Sha21e, Sha22b, Sha22a, Sha23b, Sha23c, Sha23a, Sha23d, Sha23f, Sha23e, Sta21]. ISBN-13 [Sha20e, Sha21e]. isolation [WW22b]. isotropic [Zha21a]. issues [BBBS23, SSS20]. Istanbul [Akg23]. Item [Sha20f, ASB^{+}23, Sha20h, Sha20g, dSAF22]. items [QJFY20]. Iteratively [CZ23, Zha21b]. IV [Gho20]. Ivo [Ano23a]. ix [Sha20b, Sha20i, Sha22b, Sha23d].

jackknife [CH21, JZ22a, JZ22b]. Jenkins [FPM22]. Jennifer [Sha20e]. Jessica [Sha21e]. Jin [Sha22b]. Jocelyn [Sha21d]. John [Che20a, Sha20d]. John-Ran [Sha20d]. Joint [CvNR23, SAR20, AHB24, KK22, SMZ20]. jointly [Çet21, ÇSK22, EH22]. jump [MJK20].

K-L [AZL21]. kappa [RNSR20]. Kelley [Sha21d]. Ken [Sha21d]. Kendall [SU20, SH23, XYCZ23, ZYW23]. Kernel [Nis23, YK21, AZS^{+}24, BAM22, CZ23, ES21, ESA23, HPP23, KM22a, KK20, SKH21, Ten20]. kernel-based [ES21, ESA23]. kernels [Tsi20]. KFC [HFM21]. Kibria [DA21]. Kies [EES24]. knots [OV21]. Korosteleva [Sha22b, Sha23d]. Kriging [TWN22, ZFLL23]. KS [OKF^{+}21]. Kullback [NN23a]. Kumaraswamy [MC21, SYG24]. Kumaraswamy-exponential [MC21]. Kwartler [Che20a].

L [Sha20a, Sha20e, AZL21]. lambda [DPS21]. Langevin [KSL22]. language [Sha22b]. Laplace [DLO22, DA23, YY23]. large [HWF22, MP21, SG23, WW23]. large-sample [MP21]. large-scale [SG23]. larger [dSdNB22]. Lasso [JS22a, WW21, CHJS20, KH23, LW21b, WB22, AAY21, LC20, DRG23, ZSWN21]. LASSO-penalized [DRG23]. latent [ASB^{+}23, CvNR23, FL23, SO20]. law [OKF^{+}21, WW23]. layer [Tak21]. Learn [Sha20i]. Learning [Sha21e, BHA22, CHG21, DC23, GWNW22, GC20, HFM21, KV23, MMH22, ZTP21]. least [CZ23, GA20, Jia21, Zha21b]. least-squares [Jia21]. leave [BGV20, NC23]. leave-future-out [BGV20]. leave-one-out [NC23]. Lee [Sha20d]. left [Che20c, SRU23, ZHF21]. left-truncated [ZHF21]. left-truncation [Che20c]. Leibler [NN23a]. length [BAM22, Che20b, KK20, PPCC21, PPC^{+}23]. length-biased [BAM22, Che20b, KK20]. Leone [Lon22]. levels [ES22]. Leveraging [KR21]. Li [Sha20a, Sha21c, Sha22b]. Li-Chun [Sha20a]. life [AA22a, CLY22, FPMC20, KJB22, MTW21, NFP21, PYF20, SK24, Sha20c]. lifetime [BLA20, EES24, HCN20, KAN21, KKK23, Sha21g]. lifetimes [KM21]. lifting [KV23]. Likelihood [LBHX23, SZT^{+}23, AFS21, AZIE21, AJ21, Bal21, BL20a, ÇSK22, CM21, CTY23, GvRS23, GH23, JZ22a, JZ22b, KZ21, KGK24, LG20, Li21, LRN22, Li24, MN23, MB23, NC23, PP24, QAAM22, QZ22, QH21, SN20, SL23b, YLHZ20, YWW23b, Zha21b]. Likelihood-based [SZT^{+}23, GvRS23]. likelihood-type [SN20]. limit [DWXW22]. limiting [GWL21]. limits [Fer23, QHCB20]. Lindley [KCO20, AAR24, Bal21, BCM21, EES24, JRM20, KGK24, MSBS22]. Linear [ZW21, AF20, AFS21, AML22, ASW24, ASS20, Ana20, BBSM23, CMMP20, CP22, Cha20, CHJS20, CTY23, CC24, DA21, DWWW23, Ema20, ED22, FZMZ20, GB22a, GJP24, HBB21, HCY21, HTF21, JS22a, JZ22a, JHW24, KL22a, KKLK22, KLL^{+}23, KSAR21, KÖ21, LG20, LYP20, Li21, LW23, MLG23, DRG23, NN23a, NF23, PRG22, PMW22, RLACE20, SYA23, SL23b, WKS23, WWW23, YK21, YDS^{+}20, ZYFW20, ZXG20, Zha21b, ZH22, ZY22, ZW23, ZLL24, ZFZ23, dSPG20, dSdNB22, dSLP^{+}23]. link [AFAK24]. linked [GW22]. Liouville [ZQAH23]. Literacy [Sha20b]. Liu [Abo23, AAQ20, ED22, GB22a, LAW20, MAA22, SAA^{+}22, TÖSY21]. Liu-type [AAQ20]. Liya [Sha22a]. LMARS [BP21b]. Local [TC21, DFLT21, HYWW21, NN23a, RLACE20]. Locally [LE21, AF20, CCM20, FZMZ20, HTX23]. Location [Lee20, MB22, ANA^{+}22, AAN^{+}23, BS21, KL22b, KM22b, LK22, Mar21, NM20, NC23, PK22b, QL20, SMZ20, ZY21]. location-scale [BS21, KL22b, LK22, PK22b, QL20, ZY21]. log [BCM21, HCY21, MN23, ZXZ^{+}22]. log-Lindley [BCM21]. log-logistic [HCY21]. log-normal [ZXZ^{+}22]. Logarithmic [Sul21, ZYFW20, ZC21, SBB21]. logistic [AFAK24, AL20, HCY21, LAW20, MSMTK22, NSK21, PMW22, SL23a, SO20, TLGW20, VLRZ21, ZXDD22]. logit [Cha23]. lognormal [CGGE20, JG22]. logrank [DF20]. Lomax [DSH24, YGWH21]. Long [AP20a, De 20, JIK22, SDP24, WLLH22]. long-range [JIK22]. Long-tailed [AP20a]. longitudinal [AJ21, AHB24, ASB^{+}23, BP24b, CTY23, DHN22, FFMZ22, FFS23, HWF22, Jan23, JS24, LLWZ20, LRN22, LXH20, Nik20, Sha22a, SL23b, TAWL21, YH20, ZXDD22, dSAF22, vdNPCvB22, Sha22a]. loss [AJD21, Han20, Han23, IS23b, PRG22, SZS21, WL21]. loss-based [PRG22]. lot [MJSB23]. lower [JS22b, NET^{+}24, SSN20]. LS [WWW23].

M [Sha21e, Sta21]. machine [Kaz24, MMH22]. Mahalanobis [NN22]. Maintained [MHKS22]. maintenance [AA22a, JNFTM^{+}21, SAR20]. majorization [WZ21, Zhe21]. majorization-minimization [Zhe21]. management [OSG20]. Mandelbrot [Mah23]. manifold [FLPV21, GJQ22]. Mantel [BSH22]. manufacturing [MSR23]. many [AH22, SAB21, SN20]. mapping [OSG20]. Marcinkiewicz [WW23]. marginal [Sha21k]. Markov [BBBS23, LWWD20, SW20, VS20, YWW23a]. Markov-switching [YWW23a]. marrow [LBHX23]. MARS [BP21b]. Marshall [AHB24]. Maruyama [ZQAH23]. masked [MMH20, MMH22]. masking [MMH20]. matching [ZGO24]. Mathematical [Sha20k, STÖ21]. Mathematics [Sha20i, Sha23b]. matrices [DFLT21, SLM^{+}22, WLS^{+}21]. matrix [ASB^{+}23, GTMP22, KD20, LXH20, RH21, XLF22]. matrix-variate [GTMP22]. max [Sab23, QL22]. max-INAR [QL22]. max-type [Sab23]. Maxence [Sha23b]. maxima [Mir21]. maximization [LXC23]. maximizing [Sha20o]. Maximum [HR20, KZ21, NC23, AFS21, AZIE21, Bal21, BLA20, LG20, QAAM22, Zha21b]. maximum-likelihood [Bal21]. Maxwell [Abo23, dSMA22, SBA23, GMFB21, IS23a, SCG21, SAA^{+}22]. MCMC [EES24, GPH23, MJK20, PEMCPRC22, YLYD24]. MDDRCINAR [LJW23]. Mean [SRG22, ZKG21, AF20, ASOH23, AI23, AFR23, ASAA23, BKM21, BP24a, CPX22, CS23, CTB20, GM22, GCF^{+}20, GB22b, HR20, Haq20, HKLA21, HSK22, HNK24, IMAI20, IS23a, JG22, JT24, KB20, KIS20, uA20, uAS23, PYF20, QHCB20, SCA20, SSS^{+}22, SS22, SYVV20, SKSS22, San23, TL21, ZWW21]. mean-of-order- [GCF^{+}20]. means [AZL21, ASB^{+}23, CGGE20, GT22, KLC23, KÖ21, LLP^{+}23, SB20, TAVP20, ZXZ^{+}22]. measure [AH22, BAM22, KN22, Mod22]. Measurement [Zha22, ASP22, CdSL22, Che20b, Che20c, Ema20, GB22a, HAC20, HJL22, KL22b, KK22, uAR21, Sab23, SMGAV23, SBB21, Sir20, TMMCS21, ZYFW20, ZXG20, ZC21, ZKG21, ZY22, ZZY22]. measurements [WR22]. measures [AL20, BAAR21, GB22a, GAN^{+}20, HRB21, ZHF21, ZY21]. Measuring [Sha20l]. Mechanism [YY20, MMH20]. mechanisms [FSLW21, SM21]. media [KV23]. median [KIS20, LCY22]. medians [JG22]. mediation [Sha22b]. medical [Kel21, Sha21j]. Memory [uA20, AAN^{+}23, ANA^{+}23, SDP24, WLLH22]. memory-type [AAN^{+}23, ANA^{+}23]. merge [DN22]. merge-split [DN22]. Meta [SK21b, Tak21, Avc24, BSH22, KV23, SB20, SBH21]. Meta-analysis [SK21b, BSH22, SB20, SBH21]. meta-heuristic [KV23]. meta-regression [Avc24]. method [AR21, AÖ21, BHA22, EES24, FNY21, GT22, GC20, JS24, KLL^{+}23, KAQ23, LYZ20, LGK20, LG20, LZCF23, LTW^{+}23b, LW22, NC23, OKF^{+}21, OSG20, QMK21, QZ22, SYA23, WW21, WFY22, ZQAH23]. methods [AJD21, ATH20, Ano22, AAR24, ASB^{+}23, BB22, CC24, DS20, EGM20, FPM22, FMH^{+}22, GT22, GA20, Jia21, KJL23, NN22, PS21, QAAM22, RN21, Sha20m, Sha20t, Sha22b, Sha23a, SN20, Sul22, SZT^{+}23, TdL22, WW22a, Won22]. metrics [FSB20, TWZ23]. Metropolis [MYM23]. MEWMA [ASAP21, Haq20, uAS23]. microsimulation [Sha20p]. MIDAS [BGGH21]. Milen [Ano23a]. minification [Sto22]. minimization [Zhe21]. Minimum [LW23, CMMP20, KJT21, LJ23, LTW23a, MP20, Sul21]. mining [Che20a]. Minitab [Sha20d]. mis [vdNPCvB22]. misclassification [TR24]. Mises [AFM22, MJK20]. missing [AMO20, BHA22, DXY21, GSRY23, GMJ23, KR21, LZCF23, RLACE20, RN21, SRG22, SBB21, TR22, WQZS23, Won22, ZFZ23, ZLZ24]. missingness [BS22a]. misspecification [QL20]. misspecified [JLZ21]. misusing [Sha20c]. Mixed [FASuA21, HVAVD20, ASAA23, ACS21a, BGGH21, CJK21, Ema20, GWNW22, HZD23, JS22a, JHW24, KAQ21, KÖ21, MNAP22, MHB23, NJR22, NF23, PTdS^{+}23, RCPA21, SDML20, SO20, TAWL21, WR22, YK21, YY23, ZADA22, Zha21b, ZXDD22, dSLP^{+}23]. mixed-effect [YK21]. mixed-effects [TAWL21, ZXDD22]. mixed-frequency [BGGH21]. mixed-homogeneously [ASAA23]. mixing [DWWW23, NAJ21]. Mixture [DA23, IMAI20, KIS20, CA20, DSSNL20, DK21, LW22, Nik22, SCKW21, WTZ20, XLF22, ZY21, vdNPCvB22]. mixture-based [DSSNL20]. mixtures [CdSL22, DA23, DK21, GWNW22, KK20, MJK20, NFP21, NF23]. MOBK [DLK24]. Modal [ZY22]. Model [AAQ20, DXY21, GTMP22, GSRY23, MXYY22, Sha20m, vdNPCvB22, Abo23, AFS21, AP20a, ADA22, AALA22, AKAA22, AML22, AP20b, AB23, ACS21b, AGY21, BBBS23, BGGH21, BP21b, CMMP20, CIPCR23, Çet21, Cha23, Che20b, Che20c, CMZ24, CC21, CG21, Dab24, DA21, DSSNL20, DDD23, EES24, ES22, FWC24, FSLW21, FNY21, FFNM20, FH22, FL23, GMFB21, GGA22, HMB22, HLC21, HD22a, HCN20, HAC20, HTX23, HW20, HJL22, HZD23, IS23b, JNFTM^{+}21, JC20, JHW24, KRK^{+}21, KL22a, KKLK22, KAQ23, KSAR21, KJT21, KSJ24, KB22, KJL23, KV23, KKK23, LCZ^{+}20, LLM23, LGM20, LG20, LAW20, LYP20, LZZ22, LT23, LWWD20, LLMY20, LLZ20, LE21, LJW23, LCP23, LBHX23, LWS^{+}23, LWG20, MNAP22, MBM20, MAA22, MN22, dSMA22, MG23, MSH20, NJR22, OCNN20, OSG20, PHW21, QL20, RM21, RCPA21, RH21, STG22, SK24, SAA^{+}22, SBA23]. model [SZ24, SLER21, SSS20, SAMA21, Sto22, SW20, SM21, SBAC21, SZT^{+}23, SO20, TAWL21, TLGW20, TG21, TÖSY21, TC21, VS20, WTZ20, WX21, XQ21, YK21, YDS^{+}20, YDL23, YK20, YXYL22, ZADA22, ZY20, ZH22, ZYW22, ZFLL23]. Model-based [GTMP22, Sha20m, KB22]. model-equivalence [CMZ24]. Model-free [MXYY22, LCZ^{+}20]. Modeling [KH20, LKS21, LK23, Sha20p, Sha21d, Sha21k, Sha22b, Sha23c, Sha23e, SOBC23, dSAF22]. Modelling [PY24, SR23, Aça23, AHB24, DA23, JC20, KB24, LSM20, LLMY20, LTS^{+}23, DRG23, MSBS22, MT20, WLZ21, vdNPCvB22, Sha21c]. models [AC20, AR21, AE23, ASW24, AZS^{+}24, Ana20, AK23, Asa22, AL20, AFM22, BBBS23, BBC21, BBSM23, BHAK21, BB22, BGV20, CdSL22, CP22, Che22b, CTY23, CvNR23, DN22, DDL^{+}20, DWXW22, DWWW23, DXF^{+}23, DT23, Ema20, ED22, ENK^{+}23, FFMZ22, FFS23, GTMP22, GPH23, GB22a, Gho20, GJP24, GYL24, GC20, HTF21, HVAVD20, HPV23, HDP21, HZ23, JS22a, JLZ21, JZ22a, KLL^{+}23, KM21, KNM23, KL22b, KJL23, KL24, KK22, KKZ23, KÖ21, KHK20, Lee20, LKS21, LK22, LJ23, LGM20, LW21b, Li21, LRN22, LW23, LTY^{+}24, LW22, MLG23, MN23, MGB21, MN22, MW24, MHB23, NET^{+}24, NF23, PY24, Par21, PTdS^{+}23, QL20, RK22, RMM24, SMGAV23, SKH21, Sha21g, SCW21, SRU23, SCKW21, Sir20, SK20, dSSL23, SZ21, SL23b, TWN22, TWZ23, Tsi20, VR20, WKS23, WR22, WWW23, YY23, YZ21, ZYFW20]. models [ZXG20, Zha21b, ZW21, ZC21, ZKG21, ZHF21, ZLLB21, Zha22, ZFZ22, ZY22, ZW23, ZLFW24, ZZLL20, ZXDD22, ZLL24, ZZY22, ZFZ23, dSPG20, dSdNB22, dSLP^{+}23, Sha20f]. moderate [LLBW23]. moderation [Sha22b]. modern [Sha20j, Sha23a, Ano22, Sha23a]. modification [AI23]. Modified [BAAR21, HCY21, PG22, WKS23, ZFLL23, Abo23, AFS21, ES22, FWC24, KD20, LC20, RCPA21, SAA^{+}22, SJB23, ZQAH23]. Mohamed [Sta21]. moment [AMO22, BAAR21, CCL^{+}20, De 20, DS22, GM22, WW23, ZXG20]. moments [JZ22b]. Momiao [Sha23f]. monitor [MSR23, uAS23, QHCB20]. Monitoring [HBB21, HBKB22, LK22, MHKS22, TH21, ANA^{+}22, AAN^{+}23, ASAP21, AAA^{+}22, Ali20, AFR23, ASAA23, CLL23, CTPY22, CTB20, EGM20, EGA21, HR20, Haq20, HA23, HCZ^{+}21, HZXS24, JNFTM^{+}21, JZY24, KAN21, KCC22, LT23, MSMTK22, Mir21, SCA20, SMZ20, TL21, TMMCS21, WS22, YLHZ20, ZWW21, Ano21]. monotonic [PA23]. Monte [YSJL20, BHAK21, ENK^{+}23, GCF^{+}20, OSG20, TÖSY21]. Monte-Carlo [GCF^{+}20]. Montse [Sha20e]. Morgenstern [MB22]. most [MGA21]. motion [Mah23]. moving [AAA^{+}22, ARA^{+}21, ASAA23, AK23, CLL23, HPV23, HZXS24, ISZ^{+}23, dSMA22, Mir21, MSR23, uAT20, uAR21, SMZ20, TMMCS21]. Mroz [TÖSY21]. MSE [Han20, MGA21]. MSEs [IS23a]. MSPE [LLP^{+}23]. Mukerjee [PS24]. Multi [AH22, GGA22, EH22, KV23, LWS20, LTS^{+}23, PK22b, QH21, SSS^{+}22, XSHC20]. multi-arm [LWS20]. multi-auxiliary [SSS^{+}22]. multi-chart [QH21]. multi-charts [EH22]. Multi-choice [GGA22]. multi-dimensional [LTS^{+}23]. multi-media [KV23]. Multi-rater [AH22]. multi-response [XSHC20]. multi-sample [PK22b]. multichannel [Ano23c, DM22]. multicollinear [AAY21, RMM24]. multicollinearity [DA21, MAA22]. multicomponent [Akg23, KKK23, KGK24, STG22]. multicomponents [AKA22]. multidimensional [FL23, SO20]. Multilevel [Cha23, AL20, GMFB21, Sha21d]. multinomial [Cha23, MK20, YXYL22]. Multiple [LK23, MMH22, ANA^{+}23, AHB24, BVC24, CMMP20, DXY21, FZMZ20, HM23, KJB22, LC20, LW23, LZCF23, MVFFC22, MG23, PK22a, RN21, SAR20, SSV21, SW20, TAWL21, TK24, WKL^{+}24, WR22, WW22b, XCXZ22, YZK22]. multiple-dependent [TK24]. multiple-output [FZMZ20]. Multiplicative [ZXG20, LWG20, ZYFW20, ZC21, Zha22]. Multiply [WQZS23]. Multiresolution [GFPG22]. multisample [NM20]. multistage [SAMA21]. Multivariate [Sha20n, TAWL21, ZFZ22, dSLP^{+}23, ALAS21, ACS21a, BS21, BT24, CC24, CTPY22, CJK21, CTB20, DSM22, GL22, GFPG22, GB22b, HAV20, HKLA21, HBB21, ISZ^{+}23, JS24, KK21, LLZ20, LXH20, LW22, LA20, NN22, PB21, RDSG21, SCA20, Sab23, SMGAV23, Sha21k, VPMA23, XG22, ZLLB21, ZH22, ZWW21, dSAF22]. mutual [RDSG21].

Naim [Sha20c]. natural [SKO20]. Nature [Ano22, Sha23a, Sha23a]. nearest [MP21, RDSG21]. nearest-neighbor-based [MP21]. nearly [LLBW23]. necessary [WWW23]. negative [ASB^{+}23, Dab24, DHS24, FWC24, GMFB21, KSJ24, SZ21]. negatively [ZYW22]. Nehren [Sha23b]. neighbor [MP21]. neighbourhood [SG23, SAB21]. neighbours [RDSG21]. Ness [Mah23]. nested [BS22b]. net [BPT23, JS22a, MB23]. Network [FFMZ22, CXH21, Che22a, FPM22, GW22, MT23, QH21, TWN22, ZLFW24]. Network-adaptive [FFMZ22]. network-linked [GW22]. networks [AP20a, Ano23c, BPT23, BP21b, DM22, JC20, Tak21]. neural [FPM22, JC20, MT23, Tak21, TWN22]. neutrosophic [AA23]. newly [TK22]. Neyman [AA22b]. NJ [Che20a]. NN [RLACE20]. noise [SYA23, dSSL23]. nominal [AH22]. nomination [Mir21]. Non [ANA^{+}22, HRB21, AJ21, ASS20, ASM20, AKA22, ASB^{+}23, AMO20, Ber22, BAAR21, BP21a, BM23, Çet21, CA20, GSRY23, JP21, KSJ24, KK22, KKZ23, LGK20, MN23, MRNLB20, MJG21, Pho23, SRG22, SKSS22, SSS20, TAVP20, WTZ20, ZKG21, ZLLZ22]. non-central [Pho23]. non-conventional [BAAR21]. non-curable [WTZ20]. non-Gaussian [MN23, ZLLZ22]. non-homogeneous [KSJ24, MRNLB20]. non-identical [AKA22, Çet21]. non-ignorable [AMO20, GSRY23, SRG22]. non-informative [CA20]. non-linear [ASS20]. non-negative [ASB^{+}23]. non-normal [BM23]. Non-parametric [ANA^{+}22, HRB21, ASM20, Ber22, LGK20, MJG21, TAVP20]. non-random [JP21]. non-response [ASB^{+}23, BP21a, KK22, KKZ23, SKSS22, SSS20, ZKG21]. non-stationary [AJ21]. nonconvex [WZ21, YY20]. nonhomogeneous [CAA20]. noninformative [DSH24]. nonlinear [AZS^{+}24, GL22, OCNN20, WS22, YY23]. Nonmyopic [TWS23]. nonnormality [BS22b]. Nonparametric [APAAC21, Che22a, HDP21, NJ20, THL^{+}22, WYBW22, AAN^{+}23, DN22, DWXW22, DFLT21, GK20, JIK22, KK20, LW22, Mod22, QH21, dSSL23, TK22, ZC21]. nonparanormal [AE23]. nonstationary [RT20]. norm [TZ22]. normal [AZIE21, BT24, BM23, CdSL22, CIPCR23, DA23, HA23, IS23b, JT24, KC22, KLC23, LWS20, LBHK22, Nik22, PB21, SCA20, SYVV20, SLER21, Sul22, TL21, YDL23, YXC21, ZXZ^{+}22, ZGO24]. normal-hypergeometric-inverted-beta [YXC21]. normality [ALAS21, Kel21]. normalizing [Par21]. normally [DZS21]. note [JLZ21]. notions [FPMC20]. Novel [BP21b, KV23, AB23, Hus22, MH21, NI23, WLLH22, XQ21]. nuclear [TZ22]. nuisance [SSS20]. null [SBH21]. null-distribution [SBH21]. number [ACHS23, PDD^{+}21, XLF22, ZCD22]. numbers [Sul22, WW23]. NW [Sha21b, Sha21d, Sha23c]. NY [Sha20i].

Objective [ADA22, KKLK22, HLC21, RMF^{+}21]. observation [AFAK24, MRNLB20, WWWZ21, ZY20]. observational [KKZ23, YZK22]. observations [Asa22, HBB21, Kim22, LTY^{+}24]. observed [DLK24, LTW21]. occupancy [Akg23]. odds [ATH20]. offline [Ano23c, DM22]. Olga [Sha22b, Sha23d]. Olkin [AHB24]. Omic [Sha21f]. Omnibus [LS21]. One [Haq20, KM22b, MSBS22, NC23, SAA^{+}22, Sha20f, SZT^{+}23, WLS^{+}21, YDL23]. One-sample [KM22b, WLS^{+}21]. One-sided [Haq20]. one-way [YDL23]. ones [MSBS22, SN20]. Online [GWNW22, MLG23, YZZW23]. Open [Sha20b, HGH22]. operator [FWC24, NSK21, NI23, RM21]. Optimal [BA21, BP21a, DSM22, KK21, PMKP23, SMWP20, TK24, CG21, Das22, DGMN22, ES21, KSGS20, MSH20, MJSB23, PK22a, TWS23, WLZ21, ZG21]. optimality [MP21]. optimisation [SAR20]. Optimization [NAJ21, CC24, DJA22, GC20, LTW23a, WB22, ZWW21]. optimized [EH22, SZ24]. Optimum [AGY21, VPMA23]. option [FNY21]. optional [ZKG21]. order [AMO22, CCL^{+}20, CPX22, De 20, GMCL20, GCF^{+}20, KL22a, WS22, XP20, YWW23a, ZY20, dSdNB22]. ordered [HCN20, NFP21]. ordering [ACS21a]. ordinal [Nik20, TH22, YXYL22]. organization [LG21]. Ornstein [Gra20, SP21]. ORRT [KK22, KKZ23, PS24]. orthogonal [FFNM20]. other [ASB^{+}23, GA20]. outcome [ZG21]. outcomes [AHB24, LWS20, MHB23, WQZS23]. outlier [HLK24, RMM24]. outliers [DA21, SK21c, TL21]. output [FZMZ20]. over-dispersed [AJ21, KSJ24, KH20, Li21]. over-dispersion [VR20]. overlap [PS24]. overview [Sha20t]. oxalate [AFAK24].

P [Ano21]. PA [Sha20b]. package [AFM22, DFLT21, LSM20, MT20, MAY23, PEMCPRC22, SMWP20]. pair [FSGMM23, GH23]. pair-tractable [GH23]. paired [uAT20, TH22, YKM23]. Palfrey [Sha23e]. panel [BB22, WWWZ21].

Paperback [Sha21e, Sha20c, Sha20d, Sha21d]. Parallel [LTL20, LTW23a]. parameter [Abo23, AMD21, AA22b, AGY21, BBBS23, BLA20, ENK^{+}23, Han20, HSK22, HCY21, JZY24, KÖ21, LC20, LK22, MKK20, MGA21, NC23, OV21, PK22a, PTG21, SAA^{+}22, SBA23, San23, SL23a, STÖ21, SK21c, SÇK24, SZS21, VS20, WLLH22, ZWW21]. parameter-free [HSK22]. parameters [ASAP21, AP20a, AL20, DPS21, DHS24, EES24, GSRY23, GK20, GC20, KGSA22, KLdS21, LJW23, MSMTK22, MLA^{+}20, MC21, NM20, SCA20, Sto22, TMMCS21]. Parametric [DLK24, ANA^{+}22, ASM20, Ber22, HRB21, LGK20, MJG21, SKH21, TAVP20, YZZW23, Zha22]. parent [BM23]. Pareto [Gho20, Han20, MD22, PB23]. Parkway [Sha21b, Sha21d, Sha23c]. parsimonious [KM21]. Part [Sha23a]. partial [GA20, NET^{+}24, RNSR20, ZYFW20, ZH22]. partially [AML22, AA22a, DLK24, GCF^{+}20, KL22a, KSAR21, LG20, LYP20, LTW21, MT20, SL23b, TK22, YK21, YDS^{+}20, ZLL24, ZFZ23]. Particle [CHW21, GC20, ZY20]. partition [LSL21]. partly [AMD21]. Parts [FPM22, Won22]. PaSS [CHW21]. patient [KAN21]. patterned [HNK24]. patterns [CHEH21, JP21, LG21]. Paul [Sha22a]. PDF [AAR24]. Pearson [MXYY22]. Penalized [CTY23, KHK20, LG20, OKF^{+}21, AR21, DC23, FFMZ22, DRG23, OV21, WZ21, YXYL22, ZFLL23]. penalty [MBM20]. percentiles [FPMC20]. Performance [HWF22, uAR21, AKAA22, EGA21, FPM22, FSB20, HCZ^{+}21, KCC22, NI23, uAT20, PB23, TMMCS21, WGF22]. period [MRNLB20]. Periodic [BBC21, AK23, Mah23, MN22]. periodically [Mah23]. Permutation [ACS21a, HTF21, DF20, XYCZ23, tB22]. Peter [Sha21b]. Phase [KCC22, AFR23, KAN21, Ozg21, SKSS22, MSMTK22, SAMA21, YLHZ20]. phenomena [TK22]. Pitman [Dav21]. PL [KCO20]. Place [Sha21e]. plan [JB22, KJB22, PMKP23, SJB23, TK24]. plans [CLY20, CLY22, CLY23, DSM22]. plate [CIPCR23]. Play [Sha20i]. player [Hus22]. PLMIX [MT20]. plot [Avc24, MJGAF21, Pan22]. plug [Ten20]. plug-in [Ten20]. point [Ano23c, CAA20, FSGMM23, LLM23, LYZ20, LGK20, MGB21, PK22a, SK20]. points [LC20]. Poisson [Abo23, AVMS23, CJK21, GMCL20, GAB21, HMB22, HZD23, KCO20, LE21, MRNLB20, dSMA22, MAKRK22, MSBS22, RCPA21, SAA^{+}22, SBA23, SZT^{+}23, VR20, WW22b]. Poisson-Voronoi [GAB21]. policy [AA22a, SAR20]. Pólya [FL23, PG22, VLRZ21]. Pólya-Gamma [VLRZ21]. polynomial [HYWW21, KL24, YLHZ20]. population [ASOH23, AI23, BP21a, BP24a, CB20, CM21, GAN^{+}20, IMAI20, KIS20, uA20, TAVP20, WS20]. populations [BS21, BM23, ÇSK22, JT24, LBHK22, MK20, ZXZ^{+}22]. portfolios [Li20]. positive [dSSL23, TR24]. possible [LTY^{+}24]. Post [KKLL24, CB20]. post-stratified [CB20]. Posterior [IS23a, CvNR23, Han23, SSV21, WX21]. posteriori [KZ21]. potential [ZG21]. power [AA22b, BAAR21, DLO22, JRM20, KGK24, LJ23, OKF^{+}21, RM21, SK21c]. powerful [DF20]. powers [JT24]. pp [Ano22, Ano23a, Che20a, Sha20a, Sha20c, Sha20b, Sha20d, Sha20e, Sha20i, Sha21b, Sha21d, Sha21c, Sha21e, Sha22b, Sha22a, Sha23b, Sha23c, Sha23a, Sha23d, Sha23f, Sha23e, Sta21]. Practical [Sha20o, HGH22, RLACE20]. practice [Che20a, Sha20s]. precedence [BLZ21]. precision [KD20]. Prediction [PFMT20, ZHF21, CHJS20, GA20, JLZ21, KÖ21, LLP^{+}23, Lon22, NJR22, TWN22]. predictive [AR21, Ali20, RN21, Sha22b, WX21]. Predictor [tB22, Che22a, YK21]. predictors [PZWL23]. preliminary [Kel21, MBM20]. presence [BP21a, MAA22, uAR21, RNSR20, SKSS22, SAB21, SBB21, SK21c, TWS23, TL21, WXWS21, ZKG21]. Press [Ano22, Ano23a, Sha20a, Sha20c, Sha20d, Sha20e, Sha20i, Sha21b, Sha21d, Sha21c, Sha21e, Sha22b, Sha22a, Sha23c, Sha23a, Sha23d, Sha23f, Sha23e, Sta21]. Pretest [FH22]. PRHR [AKA22]. price [JC20]. Principal [AALA22, Sir20, ZRB^{+}20, ZZLL20, BM23, DC23, XYZ20, ZLLZ22]. prior [SZS21]. priors [CA20, DSH24, LTL20, RMF^{+}21, YXC21]. probabilistic [DC23]. probability [AZL21, CH21, Dab24, DHS24, Han23, JZ22b, KZ21, KKS20, MLA^{+}20, NJ20, Sha20c]. problem [HAV20, Hus22, MNAP22, Mar21, MC22, NAJ21, SSV21, ZADA22]. problems [CHEH21, MT23]. procedure [AJ21, BAM22, CP22, HFM21, JF20, KKLL24, PP24, PK22b, SKSS22]. procedures [SDML20, WGF22, YKM23]. Process [GL22, ANA^{+}22, AAN^{+}23, AJN21, AFR23, ASAA23, Asa22, Bal21, BL20b, CLL23, CTPY22, DS20, GMCL20, HR20, Haq20, HKLA21, HBKB22, HSK22, HA23, HLC21, KGSA22, KCO20, LT23, MGT21, MLA^{+}20, MSBS22, PRG22, QHCB20, SCA20, SOBC23, SJB23, TL21, WWWZ21]. processes [BC23, CAA20, CJK21, CCM20, CTB20, FSGMM23, Gra20, GB22b, HPW23, JNNTMS24, LSL21, LLBW23, Mah23, MWMCR20, MRNLB20, PA23, PFMT20, QL22, RT20, SP21, Sha22b, Sha23d, SAMA21, WS22, WW22b, YWW23a, YLYD24, YWW23b]. Prodromou [Sha20b]. product [KKZ23, NAV23, uA20]. production [JNFTM^{+}21]. products [CLY20]. profile [CM21, EGM20, LG20, LT23, LW23, PRG22]. profile-likelihood [CM21]. profiles [EGA21, GL22, HBB21, MSMTK22, YLHZ20]. programming [KAQ21, Sha20k, STÖ21]. progressive [ANA^{+}22, AA22a, BM22, ÇSK22, EES24, HRB21, KKK23, LTW21, MTW21, MKK20, MC21, NET^{+}24, SCG21, SLER21, SÇK24]. progressive-stress [MTW21]. progressively [BA21, EN23, FASuA21, GSH20, JRM20, KGK24, MTW21, STG22, TG21]. projected [AZIE21]. Projection [XSHC20]. projections [XYZ20]. prone [LE21]. propensity [SAB21]. Properties [TTWG20, DWWW23, GAB21, PPCC21, PPC^{+}23, Zha21a]. proportion [LW21a, TR24]. proportional [AR21, BA21, DK22, HCN20, KAQ23, NET^{+}24, WTZ20, YDS^{+}20]. proportions [LWW20, Mir21]. proposed [MGA21]. prospective [ATH20]. protection [JB22]. proximity [LCP23]. pseudo [AVMS23, ES22, GH23, LTL20, MN23, Sul22, TWS23]. pseudo-copula [ES22]. pseudo-likelihood [GH23]. pseudo-log-likelihood [MN23]. pseudo-nonmyopic [TWS23]. pseudo-Poisson [AVMS23]. pseudo-priors [LTL20]. pseudo-random [Sul22]. pseudodistance [CMMP20]. public [Sha20m]. Pull [RMN21].

Qingzhao [Sha22b]. quadratic [LLWZ20, LS21]. qualitative [KRK^{+}21]. quality [CTPY22, DZS21, JNFTM^{+}21]. Quantifying [MJGAF21, WW22a]. Quantile [SCK23, CC24, DXY21, ES21, FFS23, Jan23, JS22a, Li20, LRN22, MKK20, NFP21, WKL^{+}24, YY23, ZZLL20, ZLL24, ZFZ23]. quantiles [Chr22, ESA23]. quantitative [KRK^{+}21, OSG20, Sha23b]. quasi [AJ21, LG20, Li21, Li24]. quasi-likelihood [AJ21, Li21, Li24]. quasi-maximum [LG20].

R [Sha22b, Sha23d, Sha23e, AFM22, CvNR23, LSM20, MT20, MAY23, Sha21f, Sha22b, SMWP20, Che20a, Sha21d, Sha22b, Sha23c, Sha23e, Sta21]. R-INLA [CvNR23]. Raja [Sha23b]. Ran [Sha20d]. random [ADA22, ANRS20, AL20, BA21, BP21a, CIPCR23, Cha23, CLY23, CZ23, Dab24, DWWW23, DXY21, DS22, FWC24, GSH20, GMJ23, IMAI20, JP21, KR21, PMW22, RCPA21, RDSG21, Sul22, TdL22, TLGW20, WW21, WW23, YLYD24, YDL23, Zha21b, ZYW22]. Randomized [SSS20, GGA22, JAZ22, MSH20, ZSS23, ZG21]. range [JIK22]. rank [ANA^{+}22, BLZ21, KM22b, Tac24, ZLLB21, ZYW23]. rank-sum [BLZ21]. Ranked [SC21, AE23, AAA^{+}22, AI23, BLA20, ESA23, KB20, KIS20, MB22, MT20, uAT20, PTG21, SSN20, TdL22, Tac24]. ranking [TWZ23]. ranks [PPCC21, RDSG21]. rapidly [BT24]. rare [BSH22, SBH21]. rate [Akg23, DK22, HCN20, NET^{+}24, PP24, PYF20, SA24, SZS21, WW22b, ZLFW24]. rater [AH22]. raters [AH22]. rates [Kel21, Sha21b]. ratio [ASOH23, ATH20, BL20a, BSH22, IMAI20, JT24, KLC23, KKZ23, MJGAF21, NAV23, uA20, SKSS22, SBH21, YLHZ20]. ratio-cum-product [KKZ23]. ratio-plot [MJGAF21]. ratio-type [SKSS22]. ratios [ZXZ^{+}22]. Raton [Sha20a, Sha20e, Sha21b, Sha21d, Sha21c, Sha22b, Sha22a, Sha23b, Sha23c, Sha23d, Sha23f, Sha23e, Sta21]. Rayleigh [AMD21, NJ20]. Raymond [Sha20a]. RBF [FNY21]. re [LWS20]. re-estimation [LWS20]. Real [AC20, NSK21]. real-coded [NSK21]. Real-time [AC20]. recall [PY24]. recall-based [PY24]. recapture [CB20]. record [CM21, PYF20, SSN20, SYG24]. records [AKA22, JS22b, KRK^{+}21]. Recurrent [MT23, BHAK21, Lee22, LYZ20, LGK20]. recurrent-event [LYZ20]. Recursive [VS20]. Redrawing [LL21]. Redrawing-resampling [LL21]. Reduced [GCF^{+}20, ZLLB21]. Reduced-bias [GCF^{+}20]. reduced-rank [ZLLB21]. reduction [BP24b, CLL21, Chr22, ZZLL20]. redundancy [DJA22, MNAP22, NAJ21, ZADA22, tB22]. Reference [Sha20b, HNK24]. reference-free [HNK24]. regime [LWWD20]. regime-switching [LWWD20]. regimes [SMWP20]. regional [LTL20]. regions [GAB21]. Regression [AAQ20, GMJ23, Abo23, AC20, AFS21, AALA22, AKAA22, AZS^{+}24, APAAC21, AP20b, AFAK24, ACS21b, AÖ21, AL20, Avc24, AFM22, BAM22, CLL21, CMMP20, Cha23, Che22b, CC24, CZ23, DDL^{+}20, DA21, DWXW22, DXF^{+}23, DSSNL20, DFLT21, DXY21, DDD23, DA23, ES22, ENK^{+}23, FZMZ20, FSLW21, FFS23, GJP24, HYWW21, HM23, HLK24, HZ23, IMAI20, Jan23, JZ22a, JS24, KKLK22, KLL^{+}23, KAQ23, KSAR21, Kim22, Li20, LZ20, LAW20, LW21b, LRN22, LW23, LZCF23, MGB21, MBM20, MAA22, MSMTK22, DRG23, MJG21, MVFFC22, MW24, NN23a, OCNN20, PZWL23, PEMCPRC22, PTdS^{+}23, QMK21, QAAM22, RMM24, SLM^{+}22, SBA23, SMGAV23, SKH21, SL23a, SCK23, dSSL23, SM21, SBAC21, TLGW20, TÖSY21, VLRZ21, VR20, WZ21, WKL^{+}24, WKS23, WWW23, XQ21, YY23, ZSWN21, ZLLB21, ZH22, ZY22, ZYW22, ZLY23, ZZLL20, Zhe21, ZZY22, ZFZ23, dSPG20, tB22]. regressions [Cha20, GWNW22]. regressive [MWMCR20]. regularization [AZL21, KLL^{+}23]. regularized [HLK24]. reinspection [MJSB23]. rejection [LL21]. related [GH23, SP21]. Relationship [QJFY20, Mah23]. relative [SYVV20]. relaxation [KB22]. relevancies [FFNM20]. Reliability [Çet21, LTW^{+}23b, LTS^{+}23, Akg23, AJD21, AKA22, Asa22, BLA20, BCM21, CHEH21, DK22, DJA22, EES24, FASuA21, Han23, Jia21, KKK23, KGK24, MNAP22, NAJ21, SSN20, SCG21, STG22, Sha21g, SYG24, TA20, ZADA22]. reliability-redundancy [DJA22, MNAP22, ZADA22]. removals [EN23, GSH20]. Renyi [KN22, CMMP20]. repairable [GWL21, KEC20]. repetitive [JB22, KJB22]. replacement [BKM21]. replicability [HD22b]. reported [SR23]. representation [WYBW22]. reproducibility [Kel21]. resampling [GSRY23, LL21, SC21, SRG22]. research [Ano22, Kel21, Sha20t, Sha23a, Ano22, Sha23a]. residual [FPMC20, NFP21, PYF20]. Response [Sha20f, ASB^{+}23, BP21a, DDD23, FL23, GGA22, JAZ22, KK22, KKZ23, LLZ20, MVFFC22, MSH20, Nik20, SAA^{+}22, SKSS22, Sha20h, Sha20g, SSS20, SMZ20, Won22, XSHC20, ZSS23, ZKG21, ZLZ24, dSAF22, tB22]. responses [JS24, KRK^{+}21, TC21]. restricted [DXF^{+}23, ED22, GB22a, LAW20, QAAM22, RMM24]. results [Dav21]. retrospective [ATH20]. reversible [MJK20]. Review [Ano23a, Che20a, Sha20a, Sha20c, Sha20b, Sha20d, Sha20f, Sha20h, Sha20g, Sha20e, Sha21b, Sha21d, Sha21c, Sha21e, Sha21h, Sha21i, Sha21j, Sha22a, Sha23b, Sha23c, Sha23a, Sha23d, Sha23f, Sha23e, Sta21, Ano22, Ber22, Sha22b, Sha23e]. revisit [CGGE20]. reweighted [CZ23, Zha21b]. Richard [Sha20e]. ridge [Aça23, AALA22, AZ20, Ana20, AGY21, ENK^{+}23, MGA21, PEMCPRC22, QMK21, QAAM22, RMM24, WKS23]. Riemann [ZQAH23]. Riesz [KZ21]. right [Che20c, GSH20, Jia21, JZY24, JRM20, SA24, SRU23, ZHF21]. right-censored [ZHF21]. right-censoring [Che20c]. Risk [KAN21, KL22b, BSH22, ES22, GCF^{+}20, Han23, SYVV20, SK21a, SK24, SBH21, SAMA21, TR22, WTZ20, ZY21]. Risk-adjusted [KAN21, SAMA21]. Risks [IS23a, AHB24, AMO20, DLK24, Lee22, LTW21, MMH20, MMH22, PY24, PMKP23, SCK23, TG21]. Robert [Sha21c]. Robotic [MHKS22].

Robust [AF20, ASOH23, BC23, CMMP20, CPX22, DA21, ED22, FFS23, GJP24, JNNTMS24, KJT21, KL24, LCZ^{+}20, LJ23, LXH20, MP20, MW24, SL23a, STM24, SDML20, SBAC21, SL23b, TWZ23, TR22, TH22, YKM23, YK20, YZZW23, ZW23, AFS21, AZS^{+}24, BB22, FFMZ22, PP24, RMM24, SLM^{+}22, WKS23, WQZS23, XYCZ23]. robustness [BP24b]. ROC [DCMCCB20, FLPV21, LLX24, NET^{+}24]. ROC-AUC [LLX24]. role [MC22]. root [LLBW23]. rotation [DFLT21]. row [CC24]. row-sparse [CC24]. RRT [ZKG21]. RSS [AE23, BLZ21]. ruin [Hus22]. rule [Üns22]. rules [ZSWN21]. run [AFR23, De 20, PPCC21, PPC^{+}23, YLK^{+}22].

S [SLM^{+}22]. safe [ZSWN21]. Sahu [Sha23c, Sha23e]. Samford [Sha21e]. Sample [LWS20, BA21, CB20, DF20, HSK22, HMS21, ISZ^{+}23, Kel21, KM22b, Li24, Mar21, MP21, PK22b, PB23, SM21, SÇK24, WLS^{+}21, WFY22, ZG21]. sampler [GPH23, KH23, MJK20, VLRZ21, ZTP21]. samplers [DN22]. samples [BLA20, CHEH21, JRM20, KGK24, LA20, MB22, STG22]. Sampling [Ano21, AE23, AAA^{+}22, AI23, BLA20, BP24a, CLY20, CLY22, CLY23, DSM22, ES21, ESA23, EGM20, HAV20, HSK22, HZXS24, IMAI20, JB22, KJB22, KB20, KK21, KIS20, KSGS20, LTL20, LL21, Mir21, NAV23, uAT20, Ozg21, PTG21, PMKP23, SSN20, SS22, SKSS22, Sha20s, SGY23, SSS20, SBB21, SJB23, TdL22, Tac24, TR24, TK24, TH21, VPMA23, WR22, XCXZ22, YLYD24, YLK^{+}22, ZKG21]. Sanjay [Sha23e]. SAS [Sha20d, Sta21, Sha22b]. Saunders [DDL^{+}20, KJB22, LBHX23, LCY22, PTG21, PNN22]. save [Sha20r]. scalar [MW24, XSHC20]. scale [BS21, CdSL22, GK20, HDP21, KL22b, Lee20, LK22, NM20, PK22b, QL20, Ros22, SG23, SMZ20, YKM23, ZY21]. scale-based [Lee20]. scale-space [HDP21]. scaled [AAY21]. scan [FMH^{+}22]. scenarios [BM23]. scheme [EGM20, KSGS20, LK23, MKK20, NAV23, SCA20, SSN20, SÇK24, TR24, TMMCS21, Zhe21]. schemes [AAA^{+}22, ES21, MC21, MJSB23, SMZ20]. Science [Sha20b, Ano23a, HGH22, Sha20q]. Scientific [Sha20c]. score [LWS^{+}23, SAB21, TH22]. scorecard [OSG20]. scores [Nik20]. scrambling [MSH20]. screening [HD22a, JF20, LLWZ20, LCZ^{+}20, LLZ20, MXYY22, SDML20, WB22, XSHC20, ZSWN21, ZLZ24]. search [DJA22, SG23]. seasonal [dSMA22, dSdNB22]. seasonality [BBC21, BL20b]. Second [ZY20]. Second-order [ZY20]. Seeing [Sha21e]. seemingly [ENK^{+}23]. segmentation [PK22a]. segmented [LWWD20]. selected [JP21]. Selecting [OV21, ZFLL23]. Selection [ZCD22, AGY21, BHA22, BP21b, CP22, Che20c, CG21, DK21, FFS23, FSGMM23, GW22, GJP24, HPP23, JIK22, KKLK22, KKS20, KB22, KHK20, LC20, LGM20, LYP20, MLG23, MH21, NAJ21, NI23, NF23, SM21, STM24, SBAC21, Ten20, WLZ21, WKL^{+}24, ZLY23, ZLL24]. selections [CHW21]. self [BL20a, LTW^{+}23b, YLYD24]. self-exciting [YLYD24]. self-similarity [BL20a]. self-stratifying [LTW^{+}23b]. SEM} [RNSR20, RN21]. semi [RK22, YZZW23, YZ21]. semi-parametric [YZZW23]. semi-varying [YZ21]. Semiparametric [Che20b, FSLW21, Lee22, WXWS21, ALAS21, AML22, AZS^{+}24, CMZ24, ED22, LW23, LE21, SKH21, TAWL21, ZYW22, ZXDD22]. sensitive [KK22, KKZ23, PS24, SSS^{+}22, SS22]. Sensitivity [HH21, AMO20]. sequences [DWXW22, GWL21]. Sequential [SK20, WW22b, GC20, LL21, TWS23, TK22]. Sequentially [DN22]. serial [dSAF22]. Series [Sha23b, AF20, BBC21, BHK23, BGV20, FPM22, GWL21, HTX23, KL22b, LKS21, LK22, LK23, LWWD20, MSBS22, RM21, Sha21i, SDP24, Ten20, XQ21]. series-based [Ten20]. service [AA22a]. set [AE23, AAA^{+}22, AI23, BLA20, DXF^{+}23, ESA23, KB20, KIS20, MB22, uAT20, PTG21, SSN20, TdL22, Tac24]. set-sampling [AAA^{+}22]. sets [OKF^{+}21, WR22]. setting [OSG20]. several [CGGE20, KC22, MKK20, RDSG21]. Shannon [DSH24]. shape [AA22b, DXF^{+}23, DA23, HCY21]. shape-restricted [DXF^{+}23]. shaped [BLA20, FSLW21]. shared [Dab24, JLZ21]. shift [ANRS20, GB22b, SL23a]. shifts [HNK24, ZWW21]. shiny [MN22]. shocks [KEC20]. shortfall [WYBW22]. Shoukri [Sta21]. shrinkage [AGY21, FH22, dSSL23, ZLY23]. sided [Haq20, MGT21]. Sign [HM23, KM22b, PPC^{+}23]. signal [KCC22]. signals [YXC21]. signed [ANA^{+}22, KM22b, PPCC21, RM21]. signed-rank [ANA^{+}22, KM22b]. significance [tB22]. similarity [BL20a]. Simple [ASW24, CLY22, CLY23, HBB21, PRG22, SB20, SCW21, TdL22, WWW23, ZYW23]. simulated [SC21]. simulating [BC23]. Simulation [Kur24, NN22, WR22, AKAA22, BP21a, BP24a, CJK21, DSSNL20, DT23, Gra20, KAQ23, LG21, LTW^{+}23b, OSG20, PS21, QMK21, TLGW20, TÖSY21, XG22, ZG21]. simulation* [SP21]. simulation-based [ZG21]. simulations [Sha20j]. Simultaneous [Li21, MKK20, ZXZ^{+}22, HR20, JS22b, KC22, Li24, SCA20, ZKG21, ZH22]. simultaneously [RN21]. sine [SBB21]. sine-type [SBB21]. Singapore [Sha20c]. Singhing [WGF22]. Single [HNK24, HJL22, HDP21, SDML20, SBAC21, Tak21, ZZLL20]. single-index [HJL22, HDP21, SBAC21, ZZLL20]. SIR [KJL23]. situations [AAY21]. size [BA21, CB20, CM21, HWF22, ISZ^{+}23, LWS20, WFY22, ZG21]. sizes [HSK22, Sul21]. skew [CdSL22, CIPCR23, DA23, FNY21, YDL23, ZGO24]. skew-normal [CdSL22, CIPCR23, YDL23, ZGO24]. skewed [ASOH23, Fer23, GTMP22, HPP23, MLA^{+}20, MAKRK22, MGA21]. skewness [MLA^{+}20]. skip [MJSB23]. skip-lot [MJSB23]. SLASSO [AAY21]. slice [LTL20]. Small [Sha20p, SKO20, CLL21, Li24, LLP^{+}23, RMN21, Sul21, WFY22]. small-sample [Li24]. smallest [THL^{+}22]. Smith [Sha20e]. Smooth [FZMZ20, SH23]. smoothing [FFS23, OV21, RLACE20, SCK23]. sold [CLY20]. solutions [Sha20j]. solving [SYA23]. Some [KM21, SBB21, AKAA22, PA23, QMK21]. Sons [Che20a]. Sound [Sha21b, Sha21d, Sha23c]. space [GC20, HVAVD20, HDP21, Kaz24, KJL23, LCP23, SAMA21]. spaced [Zha21a]. spacekime [Ano23a]. spaces [GPH23]. Sparse [ZLLB21, AR21, BPT23, CHJS20, CC24, FZMZ20, Kim22, Li20, XYZ20, YXC21, YLS22, ZLY23, ZWW21]. spatial [BC23, CIPCR23, FMH^{+}22, FSGMM23, GWNW22, KL22a, Kim22, LSM20, LG20, LYP20, NS22, Sha21c, Sha23e, SCW21, SO20, TWN22, YK20, Zha21a, Sha22b]. spatial-temporal [Sha21c, Sha23e]. spatially [SCW21, SKO20]. spatio [LSM20, Sha23c]. spatio-temporal [LSM20, Sha23c]. special [NAV23]. specific [SB20]. specification [vdNPCvB22]. speckled [AJN21]. spectral [JIK22, Zha21a]. Spiegelhalter [Sha20i]. spline [CIPCR23, GWNW22]. splines [OV21]. split [DN22]. Springer [Ano22, Sha20d, Sha23a]. SPRT [MGT21]. square [CZ23, IS23a, KKLL24, MXYY22]. squared [NN22]. squares [CLL23, De 20, GA20, Jia21, Zha21b]. Srivastava [Sha23e]. Stability [BM23]. stable [BT24, CCM20, Gra20, KLdS21, XG22]. stage [BS22b, BPT23, Das22, GGA22, LWS20, MYM23, SK24]. Stamford [Sha21e]. standard [SDP24]. standards [IS23a]. standby [DJA22, LTW^{+}23b]. state [GC20, HVAVD20, KJB22, KJL23, LCP23, SAMA21, TK24]. state-space [GC20, HVAVD20, KJL23, LCP23, SAMA21]. states [GC20]. stationary [AJ21, CCM20, GWL21, HTX23]. statistic [BBSM23, CLUX21, NM20, Pho23, WW22a]. Statistical [Asa22, GAB21, HJL22, KEC20, MTW21, MTSR22, MK20, NET^{+}24, Sha20b, Sha20q, Ano22, CF21, DXF^{+}23, DS22, EGM20, KK21, LG21, LCY22, MN21, PEMCPRC22, Sha20d, Sha20h, Sha20j, Sha21k, Sha23a, WLZ21, Sha22b]. Statistics [Sha20r, AMO22, AA23, BVC24, CS23, Das22, FSB20, MP20, MK20, Sha21a, Sha20i, Sha20l, Sha21h, Sha21j, Sha22b, Sha23e, XP20, dSdNB22, Sha20e, Sha21e]. status [AML22, YDS^{+}20]. steganography [KV23]. Stein [MBM20, SZS21]. Stein-type [MBM20]. step [AA22a, CLY22, CLY23, DRG23, SK21b, SK24, XCXZ22]. step-stress [AA22a, CLY22, CLY23, SK21b, SK24]. stepwise [CHW21]. Stewart [Sha20i]. stigmatized [SSS20]. Stochastic [KSL22, YSJL20, AC20, ACS21a, DJA22, JC20, LAW20, LCH20, Sha23d, SYA23, ZQAH23, Sha22b]. strategic [OSG20]. strategies [FH22, Sha20j, Sha23b]. strategy [DJA22, LCH20, MNAP22, NAJ21, ZADA22]. stratified [CB20, ES21, GGA22, HAV20, IMAI20, Ozg21, SS22, SCK23, ZKG21]. stratifying [LTW^{+}23b]. streaming [MLG23]. strength [Akg23, AKA22, BLA20, BCM21, Çet21, DK22, KKK23, KGK24, SSN20, SCG21, STG22, SÇK24]. strengths [Çet21]. Stress [AKA22, DK22, Akg23, AA22a, BLA20, BCM21, Çet21, CLY22, CLY23, KKK23, KGK24, MTW21, SSN20, SCG21, STG22, SK21b, SK24, SLER21, SÇK24]. stress-change [CLY23]. Stress-strength [AKA22, DK22, Akg23, BLA20, BCM21, Çet21, KKK23, KGK24, SSN20, SCG21, STG22, SÇK24]. strong [WW23, WWW23]. Structural [LLMY20, LLM23, YWW23b]. structural-change [YWW23b]. structure [Che22b, FFNM20, MLA^{+}20, PZWL23, PFMT20]. structured [PDD^{+}21]. structures [BM23, Che22a]. Student [HVAVD20]. Student- [HVAVD20]. studies [ATH20, KR21, MWS23, SBH21, Sha21f, SCK23, WR22, YZK22]. study [AAN^{+}23, DT23, DDD23, ENK^{+}23, FPM22, GMJ23, HD22b, HMS21, LCH20, PS21, SB20, SM21, TÖSY21]. study-specific [SB20]. subgroup [WXWS21, YH20]. subgrouping [HBKB22]. subinterval [OKF^{+}21]. subject [Che20c, KEC20, TR24]. subsampling [WW22a]. subsampling-based [WW22a]. Subset [GW22, LTW^{+}23b, WLZ21, ZLY23]. subspace [XYZ20]. subsystems [NAJ21]. successive [BP24a]. such [GH23]. Sudhir [Sha22a]. Sufficient [Chr22, LZ20, WWW23]. Suite [Sha21b, Sha21d, Sha23c]. Sujit [Sha23c, Sha23e]. sum [AFR23, BLZ21, CLL23, YLK^{+}22]. summands [Sul22]. sums [WW23]. superior [PK22b]. supervised [CLL21, HFM21]. Support [Kaz24, Zhe21]. surrogate [LCZ^{+}20]. survey [Cha23, SSS20, SBB21, XCXZ22]. surveys [uA20]. survival [AR21, BAM22, CG21, DF20, KHK20, LSL21, MT23, SKH21, SZ24, SMWP20, TZ22]. susceptibility [Dab24]. SVAR [MN23]. swarm [CHW21, GC20]. switching [LWWD20, YWW23a]. Switzerland [Ano22, Sha23a]. Symmetry [JS22b]. System [DJA22, MHKS22, AKA22, CM21, DK22, GWL21, HCN20, KEC20, MJSB23, TA20]. systematic [KSGS20]. Systems [MHKS22, LTW^{+}23b].

table [Sul21]. tackle [RMM24]. Tai [Sha20d]. Tail [ZY21, PB23, PB21, WL21, XP20]. tail-inflated [PB21]. tailed [AP20a, ATH20, BPT23, HPP23, MWMCR20]. target [OSG20]. target-setting [OSG20]. tasks [LLX24]. tau [XYCZ23]. taxicab [GPH23]. technique [EES24, JAZ22]. techniques [GMJ23, SBB21]. Ted [Che20a]. temperature [dSdNB22]. tempered [BT24, Gra20, XG22]. template [BBBS23, IAO^{+}23]. temporal [LSM20, Sha21c, Sha23c, Sha23e]. tensor [LZ20]. tessellation [GAB21]. Test [SK21c, ALAS21, AA22b, BS21, BKM21, BL20a, CLY22, Che22b, CLY23, De 20, KB22, KM22b, Lee20, LS21, LLM23, LWS^{+}23, MBM20, MTW21, MP20, Mar21, MK20, NM20, PK22b, PS21, QZ22, SB20, SBH21, SCW21, Sul21, SU20, SH23, TK22, TTWG20, TAVP20, TH22, WL21, XYCZ23, YLHZ20, YWW23b, Zha21a]. tested [JS22b]. Testing [CAA20, CMZ24, FPMC20, JT24, ZLFW24, ACS21a, BS22b, BS21, CGGE20, CPX22, GM22, HRB21, HTF21, KL22a, KB20, Li24, NJ20, NN23b, QL20, SK24, SSV21, SW20, TAVP20, YKM23, ZW23, ZYW23, dSCdC22, tB22]. Tests [FFNM20, RDSG21, AA22a, ANRS20, APAAC21, ASM20, Ano23c, AFM22, BLZ21, CMMP20, CS23, DM22, DLO22, DF20, GK20, HM23, HMS21, Kel21, KM22b, MJG21, MVFFC22, NN23a, PA23, RNSR20, RN21, SA24, WLS^{+}21, YZK22]. text [Sha20o, Che20a]. TF [BGGH21]. TF-MIDAS [BGGH21]. their [AAR24, NN22, PS24, SP21]. Theodosia [Sha20b]. theorem [DWXW22]. theoretical [RLACE20]. Theory [Sha20f, Sha20s, AÖ21, LTS^{+}23, Sha20h, Sha20g, Sha20k]. there [RLACE20]. thermostat [ASAP21]. Thin [CIPCR23]. thinning [RM21]. Thomas [Sha23e]. three [DDL^{+}20, Hus22, SKSS22, Sha20g, Sha21i, Sha21j, Sha22b, Sul21, Won22]. three-phase [SKSS22]. three-player [Hus22]. three-way [Sul21]. threshold [FWC24, KKS20, LZZ22, YLYD24]. thresholding [SGY23]. thumb [Üns22]. ties [FLPV21]. tilting [GSRY23, SRG22]. Time [MHKS22, AF20, AC20, Ali20, Ano23a, BBC21, BKM21, BHK23, BGV20, CAA20, CPX22, CLY23, CC21, FFMZ22, FPM22, FH22, GM22, HCZ^{+}21, HTX23, KJB22, KL22b, KCC22, KHK20, LKS21, LK22, Lee22, LK23, LSL21, LWWD20, LCP23, MWS23, MMH20, MSBS22, uA20, RM21, Sha21i, Sha21k, SDP24, TA20, WXWS21, XQ21, ZHF21]. time-based [uA20]. time-between-events [Ali20]. time-to-event [CC21, KHK20, ZHF21]. time-varying [FFMZ22, LCP23]. tobacco [WKS23]. Tobit [SM21, TÖSY21, AÖ21, AGY21]. tolerance [KC22, LBHK22]. tool [Pan22, vdNPCvB22]. tools [HGH22, Sha20h]. Topp [Lon22]. total [WS20]. tracking [Li20]. Tractable [ZGO24, GH23]. trading [Sha23b]. traditional [QHCB20]. Tran [Ano21]. transaction [FNY21]. transfer [BGGH21]. transform [KV23]. transformation [AML22, Che20b, VR20]. Transformed [JZ22b]. transition [BHK23]. transplant [LBHX23]. trapezoidal [ACHS23]. treatment [MWS23, SMWP20, TH22, WQZS23]. treatments [SAB21]. tree [CHEH21, GPH23, JS24, XQ21]. tree-patterns [CHEH21]. trees [Kim22]. trend [BYH23, CPX22, MG23, PA23, SA24, ZFLL23]. trial [LWS20]. trials [Sha20j, ZG21]. Triple [HZXS24]. trivariate [LBHK22, Sul22]. Truncated [TL21, HMB22, Kur24, NET^{+}24, SRU23, ZHF21]. truncation [Che20c, Gho20, PB23]. Trustworthy [HLK24]. Tsallis [KN22]. tuning [LC20, PK22a]. Turkey [Akg23]. tutorial [DDL^{+}20]. Tweedie [LT23, RK22]. Two [ATH20, GK20, SBA23, Sul22, XCXZ22, Abo23, AMD21, AH22, BLZ21, BS22b, BS21, BLA20, CAA20, ÇSK22, CB20, Das22, DF20, EGM20, GT22, GGA22, Haq20, HMS21, JT24, JZY24, Kel21, KLC23, KÖ21, LWS20, MGT21, Mah23, MKK20, DRG23, Mar21, MYM23, NFP21, Ozg21, PTG21, PS24, RNSR20, RN21, San23, Sha20h, STÖ21, SÇK24, SZT^{+}23, TAVP20, Won22, ZLFW24]. two-parameter [Abo23, AMD21, BLA20, JZY24, MKK20, PTG21, San23, STÖ21]. two-phase [Ozg21]. two-sample [CB20, DF20, HMS21, Kel21, Mar21, SÇK24]. two-sided [MGT21]. two-stage [BS22b, Das22, GGA22, LWS20, MYM23]. Two-step [XCXZ22, DRG23]. Two-tailed [ATH20]. Type [NET^{+}24, AAN^{+}23, ANA^{+}23, AAQ20, AALA22, AKA22, AGY21, BM22, CMMP20, Çet21, EES24, EN23, GvRS23, GMFB21, HRB21, JZY24, JRM20, MBM20, MTW21, MTSR22, MP20, MKK20, MC21, NN22, uA20, PPCC21, Sab23, SS22, SKSS22, SLER21, SBB21, SN20, WYBW22, WW23, BA21, Dav21, GSH20, Kel21, Lon22, TG21]. type-I [SLER21, Dav21, Lon22]. Type-II [NET^{+}24, BM22, Çet21, EES24, EN23, HRB21, JZY24, MTW21, MTSR22, MC21, BA21, GSH20, TG21]. types [Chr22, SDML20]. Tzu [Sha20d].

U [FSLW21]. U-shaped [FSLW21]. Uhlenbeck [Gra20, SP21]. ultra [GJP24, HD22a, ZLZ24]. ultra-high [GJP24, HD22a, ZLZ24]. Ultrahigh [LLZ20, Kaz24, LLWZ20, LCZ^{+}20, MXYY22]. ultrahigh-dimensional [Kaz24]. unbalanced [ACHS23, Kel21, SO20, YDL23]. unbiased [HCY21]. unbounded [NC23]. uncertain [PDD^{+}21]. uncertainties [LTS^{+}23]. uncertainty [Ano23a, JNNTMS24, LTS^{+}23, NJR22, NN22, Sha20i, WW22a, Sha20i]. unconventional [GAN^{+}20]. under-reported [SR23]. under/over [VR20]. under/over-dispersion [VR20]. underdispersed [dSLP^{+}23]. undergraduate [Sha21h]. Understanding [Sha20t]. unequal [BLA20]. unification [ZTP21]. unified [DLK24]. uniform [ASP22]. uniformity [HRB21, NN23b]. uniformly [PK22b]. unimodal [Kur24]. unimodality [Pan22]. Unit [PTdS^{+}23, Akg23, LLBW23]. units [WR22]. univariate [Ber22, DLO22, RDSG21]. unknown [CPX22, CLUX21, EH22, HNK24, MLA^{+}20, QH21]. unrelated [ENK^{+}23]. unreported [SB20]. unstable [LLBW23]. updating [MLG23]. upper [JS22b, SYG24]. urine [AFAK24]. USA [Sha20b]. use [Sha20c, XSHC20]. Using [ZG21, ANA^{+}23, ACHS23, AZS^{+}24, ASS20, AA22a, ARA^{+}21, Ano21, AHB24, BAM22, BP21a, BP24a, CdSL22, CHG21, DS20, DFLT21, DSH24, DLK24, ES21, EES24, EGA21, FL23, GSRY23, GB22b, GAN^{+}20, HBB21, ISZ^{+}23, IS23b, JC20, KAQ21, KV23, KK20, KH20, KK22, KCC22, LSL21, LT23, LTW23a, LTW^{+}23b, Lon22, LA20, Mah23, MAKRK22, Mir21, MMH22, MJK20, MSH20, NI23, uAT20, NF23, OSG20, PP24, QHCB20, RMF^{+}21, RDSG21, SSN20, STG22, SSS^{+}22, SS22, SRG22, SKSS22, Sha21d, Sha22b, TR24, TH21, THL^{+}22, YZZW23, ZKG21, dSdNB22]. utilizing [KKZ23]. Utts [Sha21e].

validation [BGV20, IAO^{+}23, LLMY20, OCNN20, QJFY20]. value [DDD23, GCF^{+}20, MRNLB20, OCNN20, SYVV20, Sha20o, TTWG20, THL^{+}22, WFY22]. value-at-risk [GCF^{+}20]. valued [AK23, BL20b, GMCL20, KCO20, KNM23, KL24, LJ23, RM21, SOBC23, XQ21, YLYD24]. values [BHA22, HD22b, LK23, LWS^{+}23, MGA21, RN21, SYG24, WX21, Won22]. values-inflated [LK23]. Varentropy [NN23b]. variability [AFR23, LG21, SCA20]. Variable [Ano21, Che20c, LYP20, NF23, ARA^{+}21, ASB^{+}23, BHA22, CP22, CHW21, DK21, FFS23, GJP24, GAN^{+}20, HSK22, HZXS24, ISZ^{+}23, JIK22, JF20, KRK^{+}21, KKLK22, KK21, KK22, KHK20, MLG23, SCA20, SSS^{+}22, SS22, STM24, TH21, WKL^{+}24, YLK^{+}22, YLS22, ZLL24]. variable-parameters [SCA20]. Variables [ASW24, ACS21a, DSM22, DS22, FL23, JB22, KKLL24, MSH20, Ozg21, PMW22, QL22, SSS^{+}22, SZ21, WW23, XSHC20, XCXZ22, YK21]. variance [AL20, BP21a, CS23, GAN^{+}20, HR20, HA23, HWF22, JZ22a, JHW24, LZZ22, NC23, PNN22]. variances [JT24, PK22b, PNN22, SB20]. variate [GTMP22]. Variation [Ano21, AAA^{+}22, HBKB22, HZXS24, San23, TH21, YLK^{+}22]. variational [HYWW21]. variations [LG21]. various [ES22, LLP^{+}23]. varying [AML22, FFMZ22, HTX23, HJL22, HZ23, LYP20, LCP23, SCW21, SKO20, SBAC21, SL23b, YDS^{+}20, YZ21, ZLLB21, ZZLL20, ZFZ23]. varying-coefficient [AML22, HTX23, YDS^{+}20, ZLLB21, ZZLL20, ZFZ23]. VBA [Sha20d]. VCS [CTB20]. Vecchia [NS22]. vector [GL22, Kaz24, KNM23, KB24, LLM23, LWS^{+}23, MWMCR20, XSHC20, Zhe21]. vector-valued [KNM23]. vectors [CTB20, RDSG21]. Velchev [Ano23a]. Velev [Ano23a]. Velu [Sha23b]. verification [RNSR20]. Verlag [Sha20d]. version [FNY21, Hus22]. versus [Sha20i, tB22]. via [DC23, DK21, GTMP22, GA20, LLM23, LZ20, LXC23, MXYY22, MN21, WKL^{+}24, WYBW22, XYZ20, ZSWN21, ZGO24]. viii [Sha23e]. visualising [WGF22]. Visualization [Sha20b, HH21]. visualizing [Sha21a]. volatility [BHK23, DT23, JC20]. Volume [Sha20f, Sha20h, Sha20g]. Voronoi [GAB21]. vs [FPM22, KEC20].

W [Sha21d]. Wald [CMMP20, MP20]. Wald-type [CMMP20, MP20]. Walter [Ano23a]. Wang [Sha22a]. Warner [ZSS23]. warning [SAR20]. warranty [CLY20]. Wasserstein [LA20, PDD^{+}21]. Wavelet [SGY23, dSSL23, KV23, SYA23]. wavelet-based [SYA23]. way [Sul21, YDL23]. Weak [ZYW22, Li24]. web [MN22]. web-based [MN22]. Weibull [ADA22, AJD21, AHB24, Çet21, CA20, FASuA21, Jia21, KJAK23, SK21b, TdL22]. Weighted [LZCF23, Nik20, YZK22, AAA^{+}22, ASAA23, CLL23, GTMP22, HR20, HA23, HZXS24, ISZ^{+}23, JZ22b, Mir21, MSR23, uAT20, uAR21, RNSR20, SMZ20, SH23, TMMCS21, TG21, WW23, YLHZ20, YZ21, tB22]. weighting [SAB21]. Wilcoxon [KM22b]. Wiley [Che20a]. Wilks [Pho23]. wise [JF20]. within [LT23, RK22]. within-profile [LT23]. without [HBKB22, HZXS24, PEMCPRC22, Üns22]. working [XLF22]. World [Sha20c].

x [Sha23d, MTW21, QHCB20]. X-bar [QHCB20]. XII [Akg23, STG22]. xiii [Sha21d]. Xiong [Sha23f]. xix [Sha20c, Sha22b]. xv [Sha22b, Sta21]. xvi [Sha20a, Sha20i, Sha23b]. xvii [Sha21b]. xx [Sha20d]. xxi [Sha20e]. xxii [Sha23c, Sha23e]. xxiv [Sha21e]. xxv [Ano22, Sha22a, Sha23a, Sha23f]. xxvi [Ano23a, Sha21c].

Yichuan [Ano22]. York [Sha20i]. You-Gan [Sha22a]. Yu [Sha22b].

Zero [DDD23, HMB22, MSBS22, ACS21b, GMCL20, GMFB21, LKS21, LE21, LTY^{+}24, MAY23, RCPA21, SR23, SZT^{+}23, ZXZ^{+}22]. zero- [ACS21b]. zeroaltered [MAY23]. Zero-and-one [MSBS22]. Zero-inflated [DDD23, GMCL20, GMFB21, LKS21, LE21, MAY23, ZXZ^{+}22]. zero-modified [RCPA21]. zeroone-two [SZT^{+}23]. zeros [HMB22, MSBS22, SN20]. Zhang [Sha20a]. Zhao [Ano22, Sha23a]. Zygmund [WW23].

References

Alam:2022:IMS

[AA22a] Intekhab Alam and Aquil Ahmed. Inference on maintenance service policy under step-stress partially accelerated life tests using progressive censoring. Journal of Statistical Computation and Simulation, 92(4):813–829, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arnroth:2022:NTS

[AA22b] Lukas Arnroth and Rauf Ahmad. Neyman’s C(α) test for the shape parameter of the exponential power class. Journal of Statistical Computation and Simulation, 92(9):1823–1850, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

AlAita:2023:ACU

[AA23] Abdulrahman AlAita and Muhammad Aslam. Analysis of covariance under neutrosophic statistics. Journal of Statistical Computation and Simulation, 93(3):397–415, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alabi:2022:EWM

[AAA^{+}22] Oluwapelumi Alabi, Nurudeen A. Adegoke, Femi Barnabas Adebola, Bamidele Mustapha Oseni, and Saddam Akber Abbasi. Exponentially weighted moving average control charts for monitoring coefficient of variation under ranked set-sampling schemes. Journal of Statistical Computation and Simulation, 92(8):1567–1589, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akram:2022:PCR

[AALA22] Muhammad Nauman Akram, Muhammad Amin, Adewale F. Lukman, and Saima Afzal. Principal component ridge type estimator for the inverse Gaussian regression model. Journal of Statistical Computation and Simulation, 92(10):2060–2089, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbas:2023:CSN

[AAN^{+}23] Zameer Abbas, Saber Ali, Hafiz Zafar Nazir, Muhammad Riaz, Yuan Li, and Xingfa Zhang. A comparative study on the nonparametric memory-type charts for monitoring process location. Journal of Statistical Computation and Simulation, 93(14):2450–2470, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akram:2020:NLT

[AAQ20] Muhammad Nauman Akram, Muhammad Amin, and Muhammad Qasim. A new Liu-type estimator for the inverse Gaussian regression model. Journal of Statistical Computation and Simulation, 90(7):1153–1172, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asgharzadeh:2024:ILD

[AAR24] A. Asgharzadeh, M. Alizadeh, and M. Z. Raqab. Inverse Lindley distribution: different methods for estimating their PDF and CDF. Journal of Statistical Computation and Simulation, 94(3):604–623, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arashi:2021:SSL

[AAY21] Mohammad Arashi, Yasin Asar, and Bahadir Yüzbasi. SLASSO: a scaled LASSO for multicollinear situations. Journal of Statistical Computation and Simulation, 91(15):3170–3183, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andrews:2023:NGA

[AB23] Divya Kuttenchalil Andrews and N. Balakrishna. A novel geometric AR(1) model and its estimation. Journal of Statistical Computation and Simulation, 93(16):2906–2935, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abonazel:2023:NMT

[Abo23] Mohamed R. Abonazel. New modified two-parameter Liu estimator for the Conway–Maxwell Poisson regression model. Journal of Statistical Computation and Simulation, 93(12):1976– 1996, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aboubacar:2020:RTE

[AC20] Amir Aboubacar and Mohamed Chaouch. Real-time estimation for functional stochastic regression models. Journal of Statistical Computation and Simulation, 90(9):1705–1732, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acar:2023:NAM

[Aça23] T. S. Açar. A new approach in modelling the circular data: circular ridge estimator. Journal of Statistical Computation and Simulation, 93(4):671–683, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:2023:CFC

[ACHS23] Mohammad Ahmad, Weihu Cheng, Ahteshamul Haq, and Sayed Kifayat Shah. Construction of fuzzy X̄ − S control chart using trapezoidal fuzzy number with unbalanced data. Journal of Statistical Computation and Simulation, 93(4):634–645, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arboretti:2021:PTG

[ACS21a] Rosa Arboretti, Riccardo Ceccato, and Luigi Salmaso. Permutation testing for goodness-of-fit and stochastic ordering with multivariate mixed variables. Journal of Statistical Computation and Simulation, 91(5):876–896, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arora:2021:FRM

[ACS21b] Monika Arora, N. Rao Chaganty, and Kimberly F. Sellers. A flexible regression model for zero- and k-inflated count data. Journal of Statistical Computation and Simulation, 91(9):1815– 1845, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ajmal:2022:OBA

[ADA22] Maria Ajmal, Muhammad Yameen Danish, and Irshad Ahmad Arshad. Objective Bayesian analysis for Weibull distribution with application to random censorship model. Journal of Statistical Computation and Simulation, 92(1):43–59, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aghabozorgi:2023:ENG

[AE23] Hamid Haji Aghabozorgi and Farzad Eskandari. Estimation of nonparanormal graphical models based on ranked set sampling (RSS). Journal of Statistical Computation and Simulation, 93 (6):932–955, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbas:2020:RCC

[AF20] Sermad Abbas and Roland Fried. Robust control charts for the mean of a locally linear time series. Journal of Statistical Computation and Simulation, 90(15):2741–2765, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amin:2024:IOD

[AFAK24] Muhammad Amin, Azka Fatima, Muhammad Nauman Akram, and Mustafa Kamal. Influential observation detection in the logistic regression under different link functions: an application to urine calcium oxalate crystals data. Journal of Statistical Computation and Simulation, 94(2):346–359, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Azais:2022:CRP

[AFM22] R. Azaïs, S. Ferrigno, and M.-J. Martinez. cvmgof: an R package for Cramér–von Mises goodness-of-fit tests in regression models. Journal of Statistical Computation and Simulation, 92(6):1246–1266, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Antzoulakos:2023:PIR

[AFR23] D. L. Antzoulakos, K. G. Fountoukidis, and A. C. Rakitzis. A phase II run sum chart for monitoring process mean and variability. Journal of Statistical Computation and Simulation, 93(13):2276–2296, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acitas:2021:RAM

[AFS21] Sukru Acitas, Peter Filzmoser, and Birdal Senoglu. A robust adaptive modified maximum likelihood estimator for the linear regression model. Journal of Statistical Computation and Simulation, 91(7):1394–1414, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aydin:2021:OSP

[AGY21] Dursun Aydin, Öznur Isçi Güneri, and Ersin Yilmaz. Optimum shrinkage parameter selection for ridge type estimator of Tobit model. Journal of Statistical Computation and Simulation, 91 (5):952–975, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andres:2022:MRD

[AH22] A. Martín Andrés and M. Álvarez Hernández. Multi-rater delta: extending the delta nominal measure of agreement between two raters to many raters. Journal of Statistical Computation and Simulation, 92(9):1877–1897, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arabi:2024:BJM

[AHB24] Zahra Sadat Arabi, Reza Hashemi, and Taban Baghfalaki. Bayesian joint modelling of multiple longitudinal outcomes and dependent competing risks using bivariate Marshall–Olkin Weibull distribution. Journal of Statistical Computation and Simulation, 94(1):142–163, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aldrabseh:2023:NMR

[AI23] Mahmoud Zuhier Aldrabseh and Mohd Tahir Ismail. New modification of ranked set sampling for estimating population mean. Journal of Statistical Computation and Simulation, 93(16):2843–2855, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmmed:2021:GQL

[AJ21] Foyez Ahmmed and Ahsan Rahman Jamee. Generalized quasi-likelihood estimation procedure for non-stationary overdispersed longitudinal counts. Journal of Statistical Computation and Simulation, 91(9):1802–1814, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amirzadi:2021:CEM

[AJD21] A. Amirzadi, E. Baloui Jamkhaneh, and E. Deiri. A comparison of estimation methods for reliability function of inverse generalized Weibull distribution under new loss function. Journal of Statistical Computation and Simulation, 91(13):2595–2622, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Almeida-Junior:2021:APS

[AJN21] Pedro M. Almeida-Junior and Abraão D. C. Nascimento. ARMA process for speckled data. Journal of Statistical Computation and Simulation, 91(15):3125–3153, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aries:2023:PIV

[AK23] Nawel Aries and Naushad Mamode Khan. On periodic integervalued moving average (INMA(q)) models. Journal of Statistical Computation and Simulation, 93(3):366–396, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arshad:2022:SSR

[AKA22] Mohd. Arshad, Nancy Khandelwal, and Qazi J. Azhad. Stressstrength reliability estimation in a system with p-type nonidentical multicomponents from PRHR family based on records. Journal of Statistical Computation and Simulation, 92(13): 2740–2771, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akram:2022:PSB

[AKAA22] Muhammad Nauman Akram, B. M. Golam Kibria, Mohamed R. Abonazel, and Nimra Afzal. On the performance of some biased estimators in the gamma regression model: simulation and applications. Journal of Statistical Computation and Simulation, 92(12):2425–2447, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akgul:2023:EMS

[Akg23] Fatma Gül Akgül. Estimation of multicomponent stressstrength reliability based on unit Burr XII distribution: an application to dam occupancy rate of Istanbul, Turkey. Journal of Statistical Computation and Simulation, 93(18):3217–3234, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Austin:2020:BIV

[AL20] Peter C. Austin and George Leckie. Bootstrapped inference for variance parameters, measures of heterogeneity and random effects in multilevel logistic regression models. Journal of Statistical Computation and Simulation, 90(17):3175–3199, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Labadi:2021:BSG

[ALAS21] Luai Al-Labadi, Forough Fazeli Asl, and Zahra Saberi. A Bayesian semiparametric Gaussian copula approach to a multivariate normality test. Journal of Statistical Computation and Simulation, 91(3):543–563, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ali:2020:PBA

[Ali20] Sajid Ali. A predictive Bayesian approach to EWMA and CUSUM charts for time-between-events monitoring. Journal of Statistical Computation and Simulation, 90(16):3025–3050, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Mosawi:2021:IBP

[AMD21] Riyadh Rustam Al-Mosawi and Sanku Dey. Inference based on partly interval censored data from a two-parameter Rayleigh distribution. Journal of Statistical Computation and Simulation, 91(12):2527–2550, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Mosawi:2022:EES

[AML22] Riyadh Al-Mosawi and Xuewen Lu. Efficient estimation of semiparametric varying-coefficient partially linear transformation model with current status data. Journal of Statistical Computation and Simulation, 92(2):416–435, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Azizi:2020:BSA

[AMO20] Fariba Azizi, Samaneh Eftekhari Mahabadi, and Elham Mosayebi Omshi. Bayesian sensitivity analysis to the non-ignorable missing cause of failure for hybrid censored competing risks data. Journal of Statistical Computation and Simulation, 90(12): 2228–2255, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akhter:2022:OSE

[AMO22] Zuber Akhter, S. M. T. K. MirMostafaee, and E. Ormoz. On the order statistics of exponentiated moment exponential distribution and associated inference. Journal of Statistical Computation and Simulation, 92(6):1322–1346, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anatolyev:2020:RHL

[Ana20] Stanislav Anatolyev. A ridge to homogeneity for linear models. Journal of Statistical Computation and Simulation, 90(13): 2455–2472, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbas:2022:NPP

[ANA^{+}22] Zameer Abbas, Hafiz Zafar Nazir, Noureen Akhtar, Muhammad Abid, and Muhammad Riaz. Non-parametric progressive signed-rank control chart for monitoring the process location. Journal of Statistical Computation and Simulation, 92(12):2596–2622, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbas:2023:DEM

[ANA^{+}23] Zameer Abbas, Hafiz Zafar Nazir, Saddam Akber Abbasi, Muhammad Riaz, and Dongdong Xiang. On designing efficient memory-type charts using multiple auxiliary-information. Journal of Statistical Computation and Simulation, 93(4):646–670, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2021:CMC

[Ano21] Anonymous. Correction: [Monitoring the Coefficient of Variation using Variable Sampling Interval CUSUM control charts, P. H. Tran and C. Heuchenne]. Journal of Statistical Computation and Simulation, 91(7):1482–1483, 2021. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [TH21].

Anonymous:2022:BRB

[Ano22] Anonymous. Book review of statistical methods for health research. Modern statistical methods for health research, edited by Yichuan Zhao and Ding-Geng Chen, Cham, Switzerland, Springer Nature Press, 2021, 496 + xxv pp., $131.04 (hardcover), ISBN: 978-3-030-72436-8. Journal of Statistical Computation and Simulation, 92(13):2862, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2023:BRD

[Ano23a] Anonymous. Book review: Data science: time complexity, inferential uncertainty, and spacekime analytics, by Ivo D. Dinov, Milen Velchev Velev, Berlin, Germany, Walter de Gruyter GmbH Press, 2022, 463 + xxvi pp., $94.99 (hardcover), ISBN: 978-3-11-069780-3. Journal of Statistical Computation and Simulation, 93(1):209–210, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2023:C

[Ano23b] Anonymous. Correction. Journal of Statistical Computation and Simulation, 93(6):1030, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2023:CET

[Ano23c] Anonymous. Correction [Exact tests for offline change point detection in multichannel binary and count data with application to networks]. Journal of Statistical Computation and Simulation, 93(16):3008–3009, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [DM22].

Allison:2020:NTE

[ANRS20] J. S. Allison, Ya. Yu. Nikitin, I. A. Ragozin, and L. Santana. New tests for exponentiality based on a characterization with random shift. Journal of Statistical Computation and Simulation, 90(15):2840–2857, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asar:2021:NBE

[AÖ21] Yasin Asar and Esra Ögütcüoglu. A new biased estimation method in tobit regression: theory and application. Journal of Statistical Computation and Simulation, 91(6):1257–1273, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agraz:2020:LTG

[AP20a] Melih Agraz and Vilda Purutçuoglu. Long-tailed graphical model and frequentist inference of the model parameters for biological networks. Journal of Statistical Computation and Simulation, 90(9):1591–1605, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alonso:2020:AIE

[AP20b] R. Alonso and M. C. Pardo. Assessing influence on the estimated coefficients efficiency in a Cox regression model. Journal of Statistical Computation and Simulation, 90(7):1216–1229, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alonso-Pena:2021:NTC

[APAAC21] M. Alonso-Pena, J. Ameijeiras-Alonso, and R. M. Crujeiras. Nonparametric tests for circular regression. Journal of Statistical Computation and Simulation, 91(3):477–500, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adhikary:2021:FPM

[AR21] Avizit C. Adhikary and M. Shafiqur Rahman. Firth’s penalized method in Cox proportional hazard framework for developing predictive models for sparse or heavily censored survival data. Journal of Statistical Computation and Simulation, 91(3):445– 463, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amir:2021:IED

[ARA^{+}21] Muhammad Wasim Amir, Mishal Rani, Zameer Abbas, Hafiz Zafar Nazir, Muhammad Riaz, and Noureen Akhtar. Increasing the efficiency of double moving average chart using auxiliary variable. Journal of Statistical Computation and Simulation, 91(14):2880–2898, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asadzadeh:2022:SPC

[Asa22] Shervin Asadzadeh. Statistical process control based on gamma-frailty models for heterogeneous reliability observations. Journal of Statistical Computation and Simulation, 92 (2):337–351, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anwer:2023:IMH

[ASAA23] Fatima Anwer, Aamir Sanaullah, Azaz Ahmad, and Amber Asghar. An improved mixed-homogeneously weighted moving average-CUSUM control chart for efficient monitoring of a process mean. Journal of Statistical Computation and Simulation, 93(10):1644–1666, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adegoke:2021:MCW

[ASAP21] Nurudeen A. Adegoke, Adam N. H. Smith, Marti J. Anderson, and Matthew D. M. Pawley. MEWMA charts when parameters are estimated with applications in gene expression and bimetal thermostat monitoring. Journal of Statistical Computation and Simulation, 91(1):37–57, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ayilara:2023:ELC

[ASB^{+}23] Olawale F. Ayilara, Tolulope T. Sajobi, Ruth Barclay, Mohammad Jafari Jozani, and Lisa M. Lix. Estimating longitudinal change in latent variable means: a comparison of non-negative matrix factorization and other item non-response methods. Journal of Statistical Computation and Simulation, 93(2):211– 230, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amini-Seresht:2020:NNP

[ASM20] Ebrahim Amini-Seresht and Bojana Milosević. New nonparametric tests for independence. Journal of Statistical Computation and Simulation, 90(7):1301–1314, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2023:RRE

[ASOH23] Azaz Ahmed, Aamir Sanaullah, Evrim Oral, and Muhammad Hanif. Robust ratio estimators of population mean for skewed and contaminated population. Journal of Statistical Computation and Simulation, 93(5):800–817, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Sharadqah:2022:EBU

[ASP22] Ali A. Al-Sharadqah and Alexandre G. Patriota. On estimating the boundaries of a uniform distribution under additive measurement errors. Journal of Statistical Computation and Simulation, 92(10):2112–2140, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alam:2020:CEU

[ASS20] Shameem Alam, Sarjinder Singh, and Javid Shabbir. Calibrated estimators using non-linear calibration constraints. Journal of Statistical Computation and Simulation, 90(3):489– 514, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Sharadqah:2024:SLF

[ASW24] Ali Al-Sharadqah and Nicholas Woolsey. Simple linear functional errors-in-variables models with correlated errors. Journal of Statistical Computation and Simulation, 94(3):642–663, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andres:2020:TTA

[ATH20] A. Martín Andrés, I. Herranz Tejedor, and M. Álvarez Hernández. Two-tailed asymptotic inferences for the odds ratio in prospective and retrospective studies: evaluation of methods of inference. Journal of Statistical Computation and Simulation, 90(1):138–156, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Avci:2024:ACP

[Avc24] Esin Avci. Adaptation of CERES plot to meta-regression. Journal of Statistical Computation and Simulation, 94(2):400–423, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arnold:2023:BIP

[AVMS23] Barry C. Arnold, Banoth Veeranna, B. G. Manjunath, and B. Shobha. Bayesian inference for pseudo-Poisson data. Journal of Statistical Computation and Simulation, 93(6):904–931, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aldahmani:2020:GGR

[AZ20] Saeed Aldahmani and Taoufik Zoubeidi. Graphical group ridge. Journal of Statistical Computation and Simulation, 90(18): 3422–3432, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2021:MLE

[AZIE21] Amani A. Ahmed, Alyaa R. Zahran, Moshira A. Ismail, and Abd El Naser Saad. On maximum likelihood estimation of the general projected normal distribution. Journal of Statistical Computation and Simulation, 91(16):3453–3472, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amira:2021:FMC

[AZL21] Ouafa Amira, Jiang-She Zhang, and Junmin Liu. Fuzzy cmeans clustering with conditional probability based K-L information regularization. Journal of Statistical Computation and Simulation, 91(13):2699–2716, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alaei:2024:RCN

[AZS^{+}24] Bita Alaei, Karim Zare, Soheil Shokri, Mohsen Maleki, and Arezo Hajrajabi. A robust class of nonlinear autoregressive models with regression function and dependent innovations using semiparametric kernel estimation. Journal of Statistical Computation and Simulation, 94(1):50–68, 2024. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basiri:2021:ORS

[BA21] Elham Basiri and Akbar Asgharzadeh. Optimal random sample size in progressively Type-II censoring based on cost constraint for the proportional hazards family. Journal of Statistical Computation and Simulation, 91(15):3154–3169, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhatti:2021:MME

[BAAR21] Sajjad Haider Bhatti, Muhammad Azeem, Tanvir Ahmad, and Muhammad Ali Raza. Modified moment estimators based on non-conventional measures for the power function distribution. Journal of Statistical Computation and Simulation, 91(15):3048–3068, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balay:2021:EGP

[Bal21] Iklim Gedik Balay. Estimation of the generalized process capability index C_{pyk} based on bias-corrected maximum-likelihood estimators for the generalized inverse Lindley distribution and bootstrap confidence intervals. Journal of Statistical Computation and Simulation, 91(10):1960–1979, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bentoumi:2022:DML

[BAM22] Rachid Bentoumi, Mayer Alvo, and Mhamed Mesfioui. Dependence measure for length-biased survival data using kernel density estimation with a regression procedure. Journal of Statistical Computation and Simulation, 92(2):211–231, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beyaztas:2022:DDR

[BB22] Beste Hamiye Beyaztas and Soutir Bandyopadhyay. Data driven robust estimation methods for fixed effects panel data models. Journal of Statistical Computation and Simulation, 92(7):1401–1425, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bacri:2023:CIP

[BBBS23] Timothée Bacri, Geir D. Berentsen, Jan Bulla, and Bård Støve. Computational issues in parameter estimation for hidden Markov models with template model builder. Journal of Statistical Computation and Simulation, 93(18):3421–3457, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baragona:2021:PAM

[BBC21] Roberto Baragona, Francesco Battaglia, and Domenico Cucina. Periodic autoregressive models for time series with integrated seasonality. Journal of Statistical Computation and Simulation, 91(4):694–712, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barros:2023:IGS

[BBSM23] Fabiana U. Barros, Denise A. Botter, Mônica C. Sandoval, and Tiago M. Magalhães. Improved gradient statistic in heteroskedastic generalized linear models. Journal of Statistical Computation and Simulation, 93(12):2052–2066, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baddeley:2023:RAS

[BC23] Adrian Baddeley and Ya-Mei Chang. Robust algorithms for simulating spatial cluster processes. Journal of Statistical Computation and Simulation, 93(12):1950–1975, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Biswas:2021:ESS

[BCM21] Aniket Biswas, Subrata Chakraborty, and Meghna Mukherjee. On estimation of stress-strength reliability with log-Lindley distribution. Journal of Statistical Computation and Simulation, 91(1):128–150, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernstein:2022:CRU

[Ber22] David H. Bernstein. A critical review of univariate nonparametric estimation of first derivatives. Journal of Statistical Computation and Simulation, 92(16):3511–3528, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bonino-Gayoso:2021:TMT

[BGGH21] Nicolás Bonino-Gayoso and Alfredo Garcia-Hiernaux. TF-MIDAS: a transfer function based mixed-frequency model. Journal of Statistical Computation and Simulation, 91(10): 1980–2017, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burkner:2020:ALF

[BGV20] Paul-Christian Bürkner, Jonah Gabry, and Aki Vehtari. Approximate leave-future-out cross-validation for Bayesian time series models. Journal of Statistical Computation and Simulation, 90(14):2499–2523, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bar-Hen:2022:ELM

[BHA22] Avner Bar-Hen and Vincent Audigier. An ensemble learning method for variable selection: application to high-dimensional data and missing values. Journal of Statistical Computation and Simulation, 92(16):3488–3510, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bedair:2021:CFM

[BHAK21] Khaled F. Bedair, Yili Hong, and Hussein R. Al-Khalidi. Copula-frailty models for recurrent event data based on Monte Carlo EM algorithm. Journal of Statistical Computation and Simulation, 91(17):3530–3548, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boussaha:2023:AVF

[BHK23] Nadia Boussaha, Fayçal Hamdi, and Abderaouf Khalfi. On the asymmetry in the volatility of financial time series: a buffered transition approach. Journal of Statistical Computation and Simulation, 93(14):2471–2493, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhattacharyya:2021:GFT

[BKM21] Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. A goodness of fit test for mean time to failure function in age replacement. Journal of Statistical Computation and Simulation, 91(17):3637–3652, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bianchi:2020:NES

[BL20a] Sergio Bianchi and Qiushi Li. A new estimator of the selfsimilarity exponent through the empirical likelihood ratio test. Journal of Statistical Computation and Simulation, 90(11): 1982–2001, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Buteikis:2020:IVA

[BL20b] Andrius Buteikis and Remigijus Leipus. An integer-valued autoregressive process for seasonality. Journal of Statistical Computation and Simulation, 90(3):391–411, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Basikhasteh:2020:BES

[BLA20] Mehdi Basikhasteh, Fazlollah Lak, and Mahmoud Afshari. Bayesian estimation of stress-strength reliability for twoparameter bathtub-shaped lifetime distribution based on maximum ranked set sampling with unequal samples. Journal of Statistical Computation and Simulation, 90(16):2975–2990, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2021:ARD

[BLZ21] N. Balakrishnan, Tao Li, and Jiawei Zhang. Application of RSS design with rank-sum precedence tests for early decision on equality of two distributions. Journal of Statistical Computation and Simulation, 91(4):667–693, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bedbur:2022:CBE

[BM22] Stefan Bedbur and Fabian Mies. Confidence bands for exponential distribution functions under progressive type-II censoring. Journal of Statistical Computation and Simulation, 92(1):60– 80, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bispo:2023:SPC

[BM23] Regina Bispo and Filipe Marques. Stability of principal components under normal and non-normal parent populations and different covariance structures scenarios. Journal of Statistical Computation and Simulation, 93(7):1060–1076, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhushan:2021:OEP

[BP21a] Shashi Bhushan and Abhay Pratap Pandey. Optimal estimation of population variance in the presence of random non-response using simulation approach. Journal of Statistical Computation and Simulation, 91(18):3814–3827, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bulbul:2021:NMS

[BP21b] Gül Bahar Bülbül and Vilda Purutçuoglu. Novel model selection criteria for LMARS: MARS designed for biological networks. Journal of Statistical Computation and Simulation, 91(9):1749–1761, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhushan:2024:ECE

[BP24a] Shashi Bhushan and Shailja Pandey. An effective class of estimators for population mean estimation in successive sampling using simulation approach. Journal of Statistical Computation and Simulation, 94(6):1204–1235, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bortolato:2024:BRR

[BP24b] Elena Bortolato and Euloge Clovis Kenne Pagui. Bias reduction and robustness for Gaussian longitudinal data analysis. Journal of Statistical Computation and Simulation, 94(2):360–375, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bernardini:2023:SEN

[BPT23] Davide Bernardini, Sandra Paterlini, and Emanuele Taufer. A 2-stage elastic net algorithm for estimation of sparse networks with heavy-tailed data. Journal of Statistical Computation and Simulation, 93(7):1031–1059, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barale:2021:TBD

[BS21] M. S. Barale and D. T. Shirke. A test based on data depth for testing location-scale of the two multivariate populations. Journal of Statistical Computation and Simulation, 91(4):768– 785, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bagmar:2022:CIM

[BS22a] Md. Shaddam Hossain Bagmar and Hua Shen. Causal inference with missingness in confounder. Journal of Statistical Computation and Simulation, 92(18):3917–3930, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balay:2022:EHT

[BS22b] Iklim Gedik Balay and Birdal Senoglu. Estimation and hypothesis testing in the two-stage nested design under nonnormality. Journal of Statistical Computation and Simulation, 92 (5):998–1014, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bohning:2022:CIE

[BSH22] Dankmar Böhning, Patarawan Sangnawakij, and Heinz Holling. Confidence interval estimation for the Mantel–Haenszel estimator of the risk ratio and risk difference in rare event metaanalysis with emphasis on the bootstrap. Journal of Statistical Computation and Simulation, 92(6):1267–1291, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bianchi:2024:EMN

[BT24] Michele Leonardo Bianchi and Gian Luca Tassinari. Estimation for multivariate normal rapidly decreasing tempered stable distributions. Journal of Statistical Computation and Simulation, 94(1):103–125, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bello:2024:TSC

[BVC24] Debora Zava Bello, Marcio Valk, and Gabriela Bettella Cybis. Towards U-statistics clustering inference for multiple groups. Journal of Statistical Computation and Simulation, 94(1):204– 222, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bao:2023:CTF

[BYH23] Ruoyi Bao, Hiroshi Yamada, and Kazuhiko Hayakawa. l_{1} common trend filtering: an extension. Journal of Statistical Computation and Simulation, 93(4):493–512, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheema:2020:BAC

[CA20] Ammara Nawaz Cheema and Muhammad Aslam. Bayesian analysis for 3-component mixture of exponentiated Weibull distribution assuming non-informative priors. Journal of Statistical Computation and Simulation, 90(4):586–605, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cebrian:2020:TIB

[CAA20] Ana C. Cebrián, Jesús Abaurrea, and Jesús Asín. Testing independence between two nonhomogeneous point processes in time. Journal of Statistical Computation and Simulation, 90(16):2878–2901, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2020:EPS

[CB20] Kiranmoy Chatterjee and Prajamitra Bhuyan. On the estimation of population size from a post-stratified two-sample capture-recapture data under dependence. Journal of Statistical Computation and Simulation, 90(5):819–838, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2021:FAA

[CC21] Taehwa Choi and Sangbum Choi. A fast algorithm for the accelerated failure time model with high-dimensional time-to-event data. Journal of Statistical Computation and Simulation, 91(16):3385–3403, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2024:FOM

[CC24] Bingzhen Chen and Canyi Chen. Fast optimization methods for high-dimensional row-sparse multivariate quantile linear regression. Journal of Statistical Computation and Simulation, 94(1):69–102, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:2020:HOM

[CCL^{+}20] Lok Hang Chan, Kun Chen, Chunxue Li, Chung Wang Wong, and Chun Yip Yau. On higher-order moment and cumulant estimation. Journal of Statistical Computation and Simulation, 90 (4):747–771, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chou-Chen:2020:IIL

[CCM20] Shu Wei Chou-Chen and Pedro A. Morettin. Indirect inference for locally stationary ARMA processes with stable innovations. Journal of Statistical Computation and Simulation, 90(17):3106–3134, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cabral:2022:BME

[CdSL22] Celso Rômulo Barbosa Cabral, Nelson Lima de Souza, and Jeremias Leão. Bayesian measurement error models using finite mixtures of scale mixtures of skew-normal distributions. Journal of Statistical Computation and Simulation, 92(3):623– 644, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cetinkaya:2021:RES

[Çet21]  Çagatay Çetinkaya. Reliability estimation of the stress-strength

model with non-identical jointly type-II censored Weibull component strengths. Journal of Statistical Computation and Simulation, 91(14):2917–2936, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Costa:2021:ESD

[CF21] A. Costa and S. Fichera. Economic-statistical design of adaptive ARMA control chart for autocorrelated data. Journal of Statistical Computation and Simulation, 91(3):623–647, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chu:2021:GOM

[CG21] Guotao Chu and Gyuhyeong Goh. Global optimal model selection for high-dimensional survival analysis. Journal of Statistical Computation and Simulation, 91(18):3850–3863, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cakmak:2020:RTE

[CGGE20] Mehmet Cakmak, Esra Gokpinar, Fikri Gokpinar, and Meral Ebegil. A revisit to testing the equality of means for several lognormal distributions. Journal of Statistical Computation and Simulation, 90(6):961–1000, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2021:JEB

[CH21] Jen-Hao Chen and Wen-Liang Hung. A jackknife entropy-based clustering algorithm for probability density functions. Journal of Statistical Computation and Simulation, 91(5):861–875, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2020:BCI

[Cha20] Seong Yeon Chang. Bootstrap confidence intervals for a break date in linear regressions. Journal of Statistical Computation and Simulation, 90(13):2438–2454, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:2023:MML

[Cha23] Tony M. T. Chan. Multilevel multinomial logit regression model with random effects: application to flash EuroBarometer euro survey data. Journal of Statistical Computation and Simulation, 93(1):58–76, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2020:BRT

[Che20a] Li-Pang Chen. Book review: Text mining in practice with R, Ted Kwartler, Hoboken, NJ, John Wiley & Sons, 2017, 320 pp., CDN $67.15 (hardback), ISBN 1-119-28201-2. Journal of Statistical Computation and Simulation, 90(13):2497–2498, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2020:SET

[Che20b] Li-Pang Chen. Semiparametric estimation for the transformation model with length-biased data and covariate measurement error. Journal of Statistical Computation and Simulation, 90 (3):420–442, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2020:VSE

[Che20c] Li-Pang Chen. Variable selection and estimation for the additive hazards model subject to left-truncation, right-censoring and measurement error in covariates. Journal of Statistical Computation and Simulation, 90(18):3261–3300, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2022:NDA

[Che22a] Li-Pang Chen. Nonparametric discriminant analysis with network structures in predictor. Journal of Statistical Computation and Simulation, 92(18):3836–3861, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2022:TTH

[Che22b] Mingjing Chen. F -test and z-test for high-dimensional regression models with a factor structure. Journal of Statistical Computation and Simulation, 92(15):3202–3221, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chou:2021:ERT

[CHEH21] Elizabeth Chou, Yin-Chen Hsieh, Sabrina Enriquez, and Fushing Hsieh. Evaluating reliability of tree-patterns in extreme-K categorical samples problems. Journal of Statistical Computation and Simulation, 91(18):3828–3849, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cherati:2021:CDF

[CHG21] Mahdiye Rahmani Cherati, Abdorrahman Haeri, and Seyed Farid Ghannadpour. Cryptocurrency direction forecasting using deep learning algorithms. Journal of Statistical Computation and Simulation, 91(12):2475–2489, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2020:SGL

[CHJS20] Hao Chen, Yong He, Jiadong Ji, and Yufeng Shi. The sparse group lasso for high-dimensional integrative linear discriminant analysis with application to Alzheimer’s disease prediction. Journal of Statistical Computation and Simulation, 90 (17):3218–3231, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Christou:2022:SDR

[Chr22] Eliana Christou. Sufficient dimension reduction for conditional quantiles with alternative types of data. Journal of Statistical Computation and Simulation, 92(2):300–317, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2021:PSS

[CHW21] Ray-Bing Chen, Chien-Chih Huang, and Weichung Wang. Particle swarm stepwise (PaSS) algorithm for information criteriabased variable selections. Journal of Statistical Computation and Simulation, 91(11):2211–2226, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cavieres:2023:TPS

[CIPCR23] Joaquin Cavieres, German Ibacache-Pulgar, and Javier E. Contreras-Reyes. Thin plate spline model under skew-normal random errors: estimation and diagnostic analysis for spatial data. Journal of Statistical Computation and Simulation, 93 (1):25–45, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2021:BSM

[CJK21] Michael Chiu, Kenneth R. Jackson, and Alexander Kreinin. Backward simulation of multivariate mixed Poisson processes. Journal of Statistical Computation and Simulation, 91(17): 3549–3572, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cai:2021:BIR

[CLL21] Xin Cai, Guang Lin, and Jinglai Li. Bayesian inverse regression for supervised dimension reduction with small datasets. Journal of Statistical Computation and Simulation, 91(14):2817–2832, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2023:ESS

[CLL23] Jen-Hsiang Chen, Shin-Li Lu, and Chien-Tzu Liao. An enhanced sum of squares generally weighted moving average chart based on auxiliary information for process monitoring. Journal of Statistical Computation and Simulation, 93(1):188–208, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chu:2021:EDS

[CLUX21] Jianghao Chu, Tae-Hwy Lee, Aman Ullah, and Haifeng Xu. Exact distribution of the F -statistic under heteroskedasticity of unknown form for improved inference. Journal of Statistical Computation and Simulation, 91(9):1782–1801, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2020:DBS

[CLY20] Lee-Shen Chen, TaChen Liang, and Ming-Chung Yang. Designing Bayesian sampling plans for censored data under products sold with warranty. Journal of Statistical Computation and Simulation, 90(7):1230–1250, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2022:DBS

[CLY22] Lee-Shen Chen, TaChen Liang, and Ming-Chung Yang. Designing Bayesian sampling plans for simple step-stress of accelerated life test on censored data. Journal of Statistical Computation and Simulation, 92(2):395–415, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2023:DEB

[CLY23] Lee-Shen Chen, TaChen Liang, and Ming-Chung Yang. Designing efficient Bayesian sampling plans based on the simple step-stress test under random stress-change time for censored data. Journal of Statistical Computation and Simulation, 93(7): 1104–1129, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2021:EPS

[CM21] Kiranmoy Chatterjee and Diganta Mukherjee. On the estimation of population size under dependent dual-record system: an adjusted profile-likelihood approach. Journal of Statistical Computation and Simulation, 91(13):2740–2763, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Castilla:2020:RWT

[CMMP20] E. Castilla, N. Martín, S. Muñoz, and L. Pardo. Robust Waldtype tests based on minimum Rényi pseudodistance estimators for the multiple linear regression model. Journal of Statistical Computation and Simulation, 90(14):2655–2680, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2024:TSM

[CMZ24] Feifei Chen, Simos G. Meintanis, and Lixing Zhu. Testing semiparametric model-equivalence hypotheses based on the characteristic function. Journal of Statistical Computation and Simulation, 94(6):1158–1190, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cavalaro:2022:PVS

[CP22] Lucas L. Cavalaro and Gustavo H. A. Pereira. A procedure for variable selection in double generalized linear models. Journal of Statistical Computation and Simulation, 92(13):2703–2720, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2022:RTT

[CPX22] Seong Yeon Chang, Pierre Perron, and Jiawen Xu. Robust testing of time trend and mean with unknown integration order errors. Journal of Statistical Computation and Simulation, 92(17):3561–3582, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2023:BTB

[CS23] Ji-Eun Choi and Dong Wan Shin. Bootstrapping tests for breaks in mean or variance based on U-statistics. Journal of Statistical Computation and Simulation, 93(14):2357–2369, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cetinkaya:2022:ELI

[ÇSK22]  Çagatay Çetinkaya, Farha Sultana, and Debasis Kundu. Ex-

act likelihood inference for two exponential populations under jointly generalized progressive hybrid censoring. Journal of Statistical Computation and Simulation, 92(17):3605–3629, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Costa:2020:MMV

[CTB20] Antonio Fernando Branco Costa, Alexandre Fonseca Torres, and Pedro Paulo Balestrassi. The monitoring of mean vectors with VCS charts for multivariate processes. Journal of Statistical Computation and Simulation, 90(10):1897–1920, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiang:2022:NMC

[CTPY22] Jyun-You Chiang, Tzong-Ru Tsai, Hoang Pham, and Kaizhi Yu. A new multivariate control chart for monitoring the quality of a process with the aid of auxiliary information. Journal of Statistical Computation and Simulation, 92(3):645–666, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2023:PEL

[CTY23] Xia Chen, Xiaoyan Tan, and Li Yan. Penalized empirical likelihood for high-dimensional generalized linear models with longitudinal data. Journal of Statistical Computation and Simulation, 93(10):1515–1531, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiuchiolo:2023:JPI

[CvNR23] Cristian Chiuchiolo, Janet van Niekerk, and Håvard Rue. Joint posterior inference for latent Gaussian models with R-INLA. Journal of Statistical Computation and Simulation, 93(5):723– 752, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cao:2021:CDF

[CXH21] Kai Cao, Ying Xiao, and Muzhou Hou. Correlation-driven framework based on graph convolutional network for clinical disease classification. Journal of Statistical Computation and Simulation, 91(15):3108–3124, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cui:2023:IRL

[CZ23] Yue Cui and Songfeng Zheng. Iteratively reweighted least square for kernel expectile regression with random features. Journal of Statistical Computation and Simulation, 93(14): 2370–2389, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dawoud:2021:RDK

[DA21] Issam Dawoud and Mohamed R. Abonazel. Robust Dawoud– Kibria estimator for handling multicollinearity and outliers in the linear regression model. Journal of Statistical Computation and Simulation, 91(17):3678–3692, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dogru:2023:MRM

[DA23] Fatma Zehra Dogru and Olcay Arslan. Mixture regression modelling based on the shape mixtures of skew Laplace normal distribution. Journal of Statistical Computation and Simulation, 93(18):3403–3420, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dabade:2024:CNB

[Dab24] A. D. Dabade. Compound negative binomial shared frailty model with random probability of susceptibility. Journal of Statistical Computation and Simulation, 94(4):843–861, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Das:2022:ODT

[Das22] Radhakanta Das. An optimal design in a two-stage ethical allocation based on U-statistics. Journal of Statistical Computation and Simulation, 92(8):1763–1780, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Davies:2021:PCR

[Dav21] Katherine F. Davies. Pitman closeness results for Type-I hybrid censored data from exponential distribution. Journal of Statistical Computation and Simulation, 91(1):58–80, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deng:2023:EDL

[DC23] Wei Q. Deng and Radu V. Craiu. Exploring dimension learning via a penalized probabilistic principal component analysis. Journal of Statistical Computation and Simulation, 93(2):266– 297, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Diaz-Coto:2020:CDR

[DCMCCB20] Susana Díaz-Coto, Pablo Martínez-Camblor, and Norberto Octavio Corral-Blanco. Cumulative/dynamic ROC curve estimation under interval censorship. Journal of Statistical Computation and Simulation, 90(9):1570–1590, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Diop:2023:ZIG

[DDD23] Aba Diop, El Hadji Deme, and Aliou Diop. Zero-inflated generalized extreme value regression model for binary response data and application in health study. Journal of Statistical Computation and Simulation, 93(1):1–24, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dasilva:2020:ITB

[DDL^{+}20] Alan Dasilva, Renata Dias, Victor Leiva, Carolina Marchant, and Helton Saulo. [invited tutorial] Birnbaum–Saunders regression models: a comparative evaluation of three approaches. Journal of Statistical Computation and Simulation, 90(14): 2552–2570, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeGaetano:2020:BBC

[De 20] Davide De Gaetano. A bootstrap bias correction of long run fourth order moment estimation in the CUSUM of squares test. Journal of Statistical Computation and Simulation, 90(5):907– 924, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ditzhaus:2020:MPL

[DF20] Marc Ditzhaus and Sarah Friedrich. More powerful logrank permutation tests for two-sample survival data. Journal of Statistical Computation and Simulation, 90(12):2209–2227, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DiMarzio:2021:PNR

[DFLT21] Marco Di Marzio, Stefania Fensore, Giovanni Lafratta, and Charles C. Taylor. The package: nonparametric regression using local rotation matrices in [????]. Journal of Statistical Computation and Simulation, 91(11):2289–2306, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duan:2022:ECA

[DGMN22] Jiangtao Duan, Wei Gao, Yanyuan Ma, and Hon Keung Tony Ng. Efficient computational algorithms for approximate optimal designs. Journal of Statistical Computation and Simulation, 92(4):764–793, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Darabi:2022:CCF

[DHN22] Nadiyeh Darabi and S. Mohammad E. Hosseini-Nasab. A centroid classifier for functional/longitudinal data. Journal of Statistical Computation and Simulation, 92(5):1034–1046, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doi:2024:IIE

[DHS24] Jimmy A. Doi, Bret Holladay, and Mark F. Schilling. Improved interval estimation of negative binomial parameters: a coverage probability approach. Journal of Statistical Computation and Simulation, 94(1):126–141, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dobani:2022:SRR

[DJA22] Ehsan Ramezani Dobani, Mohammad N. Juybari, and Mostafa Abouei Ardakan. System reliability-redundancy optimization with cold-standby strategy by fitness-distance balance stochastic fractal search algorithm. Journal of Statistical Computation and Simulation, 92(10):2156–2183, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doo:2021:BVS

[DK21] Woojin Doo and Heeyoung Kim. Bayesian variable selection in clustering high-dimensional data via a mixture of finite mixtures. Journal of Statistical Computation and Simulation, 91 (12):2551–2568, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demiray:2022:SSR

[DK22] Duygu Demiray and Fatih Kizilaslan. Stress-strength reliability estimation of a consecutive k-out-of-n system based on proportional hazard rate family. Journal of Statistical Computation and Simulation, 92(1):159–190, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutta:2024:PIU

[DLK24] Subhankar Dutta, Yuhlong Lio, and Suchandan Kayal. Parametric inferences using dependent competing risks data with partially observed failure causes from MOBK distribution under unified hybrid censoring. Journal of Statistical Computation and Simulation, 94(2):376–399, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Desgagne:2022:CEP

[DLO22] Alain Desgagné, Pierre Lafaye de Micheaux, and Frédéric Ouimet. A comprehensive empirical power comparison of univariate goodness-of-fit tests for the Laplace distribution. Journal of Statistical Computation and Simulation, 92(18):3743– 3788, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

De:2022:ETO

[DM22] Shyamal K. De and Soumendu Sundar Mukherjee. Exact tests for offline changepoint detection in multichannel binary and count data with application to networks. Journal of Statistical Computation and Simulation, 92(17):3659–3678, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See figure correction [Ano23c].

Dahl:2022:SAM

[DN22] David B. Dahl and Spencer Newcomb. Sequentially allocated merge-split samplers for conjugate Bayesian nonparametric models. Journal of Statistical Computation and Simulation, 92(7):1487–1511, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dedduwakumara:2021:EEP

[DPS21] Dilanka S. Dedduwakumara, Luke A. Prendergast, and Robert G. Staudte. An efficient estimator of the parameters of the generalized lambda distribution. Journal of Statistical Computation and Simulation, 91(1):197–215, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mari:2023:LPC

[DRG23] Roberto Di Mari, Roberto Rocci, and Stefano Antonio Gattone. LASSO-penalized clusterwise linear regression modelling: a two-step approach. Journal of Statistical Computation and Simulation, 93(18):3235–3258, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:2020:BCI

[DS20] Sanku Dey and Mahendra Saha. Bootstrap confidence intervals of process capability index S_{pmk} using different methods of estimation. Journal of Statistical Computation and Simulation, 90(1):28–50, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dung:2022:CMC

[DS22] Le Van Dung and Ta Cong Son. Complete moment convergence for m-ANA random variables and statistical applications. Journal of Statistical Computation and Simulation, 92(2):370–394, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

dosSantos:2022:BLI

[dSAF22] José Roberto Silva dos Santos, Caio Lucidius Naberezny Azevedo, and Jean-Paul Fox. Bayesian longitudinal item response modeling with multivariate asymmetric serial dependencies. Journal of Statistical Computation and Simulation, 92(3): 488–523, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daSilva:2022:NBA

[dSCdC22] Josimara Tatiane da Silva, Juliana Cobre, and Mário de Castro. New Bayesian approaches to equivalence testing. Journal of Statistical Computation and Simulation, 92(5):957–973, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daSilva:2022:DLS

[dSdNB22] Renato Santos da Silva, Fernando Ferraz do Nascimento, and Marcelo Bourguignon. Dynamic linear seasonal models applied to extreme temperature data: a Bayesian approach using the r-larger order statistics distribution. Journal of Statistical Computation and Simulation, 92(4):705–723, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dong:2024:BAS

[DSH24] Guoqing Dong, Mohammed K. Shakhatreh, and Daojiang He. Bayesian analysis for the Shannon entropy of the Lomax distribution using noninformative priors. Journal of Statistical Computation and Simulation, 94(6):1317–1338, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daSilva:2023:MGL

[dSLP^{+}23] Guilherme Parreira da Silva, Henrique Aparecido Laureano, Ricardo Rasmussen Petterle, Paulo Justiniano Ribeiro Jr, and Wagner Hugo Bonat. Multivariate generalized linear mixed models for underdispersed count data. Journal of Statistical Computation and Simulation, 93(14):2410–2427, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duarte:2022:ODM

[DSM22] Belmiro P. M. Duarte, Satya P. Singh, and Maria J. Moura. Optimal design of multivariate acceptance sampling plans by variables. Journal of Statistical Computation and Simulation, 92(15):3129–3149, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Melo:2022:CMP

[dSMA22] Moizés da Silva Melo and Airlane Pereira Alencar. Conway– Maxwell–Poisson seasonal autoregressive moving average model. Journal of Statistical Computation and Simulation, 92 (2):283–299, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daSilva:2020:BLR

[dSPG20] Nívea B. da Silva, Marcos O. Prates, and Flávio B. Gonçalves. Bayesian linear regression models with flexible error distributions. Journal of Statistical Computation and Simulation, 90(14):2571–2591, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sousa:2023:WSN

[dSSL23] Alex Rodrigo dos S. Sousa and Nancy Lopes Garcia. Wavelet shrinkage in nonparametric regression models with positive noise. Journal of Statistical Computation and Simulation, 93 (17):3011–3033, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Desousa:2020:NMB

[DSSNL20] Mário F. Desousa, Helton Saulo, Manoel Santos-Neto, and Víctor Leiva. On a new mixture-based regression model: simulation and application to data with high censoring. Journal of Statistical Computation and Simulation, 90(16):2861–2877, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ding:2020:BAA

[DSW20] Shu Ding, Xiaoping Shi, and Yuehua Wu. Beta approximation and its applications. Journal of Statistical Computation and Simulation, 90(7):1251–1266, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DiIorio:2023:EVC

[DT23] F. Di Iorio and U. Triacca. Estimate of a volatility’s common component in ARSV models: a simulation study. Journal of Statistical Computation and Simulation, 93(11):1734–1742, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deng:2023:APE

[DWWW23] Xin Deng, Yi Wu, Wei Wang, and Xuejun Wang. Asymptotic properties of M estimators in classical linear models with ϕmixing random errors. Journal of Statistical Computation and Simulation, 93(18):3259–3284, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deng:2022:CLT

[DWXW22] Xin Deng, Yi Wu, Mengmei Xi, and Xuejun Wang. The central limit theorem for ANA sequences and its application to nonparametric regression models. Journal of Statistical Computation and Simulation, 92(14):2891–2907, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deng:2023:ASA

[DXF^{+}23] Geng Deng, Guangning Xu, Qiang Fu, Xindong Wang, and Jing Qin. Active-set algorithm-based statistical inference for shaperestricted generalized additive Cox regression models. Journal of Statistical Computation and Simulation, 93(3):416–441, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ding:2021:MAM

[DXY21] Xianwen Ding, Jinhan Xie, and Xiaodong Yan. Model averaging for multiple quantile regression with covariates missing at random. Journal of Statistical Computation and Simulation, 91(11):2249–2275, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:2021:CBE

[DZS21] Sanku Dey, Chunfang Zhang, and Mahendra Saha. Classical and Bayesian estimation of the index C_{pmk} and its confidence intervals for normally distributed quality characteristic. Journal of Statistical Computation and Simulation, 91(10): 1911–1934, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Emami:2022:RRL

[ED22] Hadi Emami and Kourosh Dadkhah. Robust restricted Liu estimator in censored semiparametric linear models. Journal of Statistical Computation and Simulation, 92(9):1851–1876, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eissa:2024:LMM

[EES24] Fathy H. Eissa and Rashad M. El-Sagheer. Lindley’s method and MCMC technique for estimating the exponentiated Kies lifetime model parameters and reliability characteristics using adaptive type-II progressive censoring with application. Journal of Statistical Computation and Simulation, 94(2):223–247, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Erfanian:2021:NAM

[EGA21] Mahdiyeh Erfanian, Bahram Sadeghpour Gildeh, and Mahmoud Reza Azarpazhooh. A new approach for monitoring healthcare performance using generalized additive profiles. Journal of Statistical Computation and Simulation, 91(1):167– 179, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eizi:2020:CBT

[EGM20] Azamsadat Eizi, Bahram Sadeghpour Gildeh, and Sayyed Ehsan Monabbati. Comparison between two methods of the economic– statistical design of profile monitoring under the double sampling scheme. Journal of Statistical Computation and Simulation, 90(18):3400–3421, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Engmann:2022:AOC

[EH22] Gideon Mensah Engmann and Dong Han. Asymptotic optimized CUSUM and EWMA multi-charts for jointly detecting and diagnosing unknown change. Journal of Statistical Computation and Simulation, 92(3):524–543, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Emami:2020:IEE

[Ema20] Hadi Emami. Improved estimation in elliptical linear mixed measurement error models. Journal of Statistical Computation and Simulation, 90(9):1681–1704, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Elshahhat:2023:AAT

[EN23] Ahmed Elshahhat and Mazen Nassar. Analysis of adaptive type-II progressively hybrid censoring with binomial removals. Journal of Statistical Computation and Simulation, 93(7):1077– 1103, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Esfanjani:2023:MCS

[ENK^{+}23] Robab Mehdizadeh Esfanjani, Dariush Najarzadeh, Hossein Jabbari Khamnei, Farshin Hormozinejad, and Mahnaz Talebi. A Monte Carlo study on the ridge parameter of the seemingly unrelated ridge regression models. Journal of Statistical Computation and Simulation, 93(13):2176–2195, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eftekharian:2021:KBQ

[ES21] Abbas Eftekharian and Hani Samawi. On kernel-based quantile estimation using different stratified sampling schemes with optimal allocation. Journal of Statistical Computation and Simulation, 91(5):1040–1056, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Erdemir:2022:MPC

[ES22] Övgücan Karadag Erdemir and Meral Sucu. A modified pseudo-copula regression model for risk groups with various dependency levels. Journal of Statistical Computation and Simulation, 92(5):1092–1112, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eftekharian:2023:KBE

[ESA23] Abbas Eftekharian, Hani Samawi, and Mehdi Amiri. On kernelbased estimation of distribution function and its quantiles based on ranked set sampling. Journal of Statistical Computation and Simulation, 93(11):1772–1798, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feroze:2021:MWD

[FASuA21] Navid Feroze, Muhammad Aslam, Tabassum Naz Sindhu, and Muhammad Noor ul Amin. Mixed Weibull distributions for the Bayesian analysis of reliability when failures are progressively censored. Journal of Statistical Computation and Simulation, 91(17):3505–3529, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feroze:2023:IEA

[Fer23] Navid Feroze. Improved estimators for analyzing highly skewed environmental datasets with detection limits. Journal of Statistical Computation and Simulation, 93(3):475–492, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2022:NAR

[FFMZ22] Kuangnan Fang, Xinyan Fan, Shuangge Ma, and Qingzhao Zhang. Network-adaptive robust penalized estimation of timevarying coefficient models with longitudinal data. Journal of Statistical Computation and Simulation, 92(14):3045–3065, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2020:TRH

[FFNM20] Dário Ferreira, Sandra S. Ferreira, Célia Nunes, and João T. Mexia. Tests and relevancies for the hypotheses of an orthogonal family in a model with orthogonal block structure. Journal of Statistical Computation and Simulation, 90(3):412–419, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2023:RSV

[FFS23] Z. C. Fu, L. Y. Fu, and Y. N. Song. Robust and smoothing variable selection for quantile regression models with longitudinal data. Journal of Statistical Computation and Simulation, 93(15):2600–2624, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Forzley:2022:PSE

[FH22] Quinn Forzley and Shakhawat Hossain. Pretest and shrinkage estimation strategies in accelerated failure time model. Journal of Statistical Computation and Simulation, 92(7):1512–1529, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2023:BIM

[FL23] Zhihui Fu and Meilan Lu. Bayesian inference for multidimensional graded response model using Pólya–Gamma latent variables. Journal of Statistical Computation and Simulation, 93(16):2856–2887, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feng:2021:HUR

[FLPV21] Qunqiang Feng, Jialiang Li, Xingrun Ping, and Ben Van Calster. Hypervolume under ROC manifold for discrete biomarkers with ties. Journal of Statistical Computation and Simulation, 91(18):3864–3879, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

French:2022:CSS

[FMH^{+}22] Joshua P. French, Mohammad Meysami, Lauren M. Hall, Nicholas E. Weaver, Minh C. Nguyen, and Lee Panter. A comparison of spatial scan methods for cluster detection. Journal of Statistical Computation and Simulation, 92(16):3343–3372, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farshadmoghadam:2021:EOU

[FNY21] Farnaz Farshadmoghadam, Ali Reza Najafi, and Mohammad Reza Yaghouti. European option under a skew version of the GBM model with transaction costs by an RBF method. Journal of Statistical Computation and Simulation, 91(14): 2986–3004, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Farimani:2022:EPT

[FPM22] Nasser Motahari Farimani, Pedram Parsafar, and Shiva Mohammadi. Evaluation performance of time series methods in demand forecasting: Box–Jenkins vs artificial neural network (case study: Automotive parts industry). Journal of Statistical Computation and Simulation, 92(17):3639–3658, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Franco-Pereira:2020:TAN

[FPMC20] Alba M. Franco-Pereira and Pablo Martínez-Camblor. Testing ageing notions through percentiles of the residual life. Journal of Statistical Computation and Simulation, 90(5):869–888, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fuglsby:2020:SEP

[FSB20] Cami Fuglsby, Christopher P. Saunders, and JoAnn Buscaglia. U-statistics for estimating performance metrics in forensic handwriting analysis. Journal of Statistical Computation and Simulation, 90(6):1082–1117, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fuentes-Santos:2023:BBS

[FSGMM23] I. Fuentes-Santos, W. González-Manteiga, and J. Mateu. Bootstrap bandwidth selection for the pair correlation function of inhomogeneous spatial point processes. Journal of Statistical Computation and Simulation, 93(18):3329–3361, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fani:2021:SRU

[FSLW21] Shabnam Fani, Hua Shen, Xuewen Lu, and Jingjing Wu. Semiparametric regression with the U-shaped baseline hazard function in the additive hazards model under general censoring mechanisms. Journal of Statistical Computation and Simulation, 91(16):3255–3282, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2024:NTI

[FWC24] Yixuan Fan, Dehui Wang, and Jianhua Cheng. A new threshold INAR(1) model based on modified negative binomial operator with random coefficient. Journal of Statistical Computation and Simulation, 94(6):1236–1268, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2020:SLS

[FZMZ20] Kuangnan Fang, Xiaochen Zhang, Shuangge Ma, and Qingzhao Zhang. Smooth and locally sparse estimation for multipleoutput functional linear regression. Journal of Statistical Computation and Simulation, 90(2):341–354, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goktas:2020:CPL

[GA20] Atila Göktas and Özge Akkus. Comparison of partial least squares with other prediction methods via generated data. Journal of Statistical Computation and Simulation, 90(16): 3009–3024, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gezer:2021:SPP

[GAB21] Fatih Gezer, Robert G. Aykroyd, and Stuart Barber. Statistical properties of Poisson-Voronoi tessellation cells in bounded regions. Journal of Statistical Computation and Simulation, 91 (5):915–933, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gulzar:2020:EEP

[GAN^{+}20] Muhammad Awais Gulzar, Muhammad Abid, Hafiz Zafar Nazir, Faisal Maqbool Zahid, and Muhammad Riaz. On enhanced estimation of population variance using unconventional measures of an auxiliary variable. Journal of Statistical Computation and Simulation, 90(12):2180–2197, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghapani:2022:DMR

[GB22a] Fatemeh Ghapani and Babak Babadi. Diagnostic measures for the restricted Liu estimator in linear measurement error models. Journal of Statistical Computation and Simulation, 92(11): 2257–2272, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guler:2022:DIM

[GB22b] Zümre Özdemir Güler and Mehmet Akif Bakir. Detection and identification of mean shift using independent component analysis in multivariate processes. Journal of Statistical Computation and Simulation, 92(9):1920–1940, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guzman:2020:NMS

[GC20] I. R. E. Guzman and P. C. C. Ccori. A new method for sequential learning of states and parameters for state-space models: the particle swarm learning optimization. Journal of Statistical Computation and Simulation, 90(11):2057–2079, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2020:RBP

[GCF^{+}20] M. Ivette Gomes, Frederico Caeiro, Fernanda Figueiredo, Lígia Henriques-Rodrigues, and Dinis Pestana. Reduced-bias and partially reduced-bias mean-of-order-p value-at-risk estimation: a Monte-Carlo comparison and an application. Journal of Statistical Computation and Simulation, 90(10):1735–1752, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garcia-Fernandez:2022:MAC

[GFPG22] Rosa M. García-Fernández and Federico Palacios-González. Multiresolution approximation and consistent estimation of a multivariate density function. Journal of Statistical Computation and Simulation, 92(12):2623–2641, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2022:MCS

[GGA22] Neha Gupta, Srikant Gupta, and Mohd. Tanwir Akhtar. Multichoice stratified randomized response model with two-stage classification. Journal of Statistical Computation and Simulation, 92(5):895–910, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gorecki:2023:CPL

[GH23] J. Górecki and M. Hofert. Composite pseudo-likelihood estimation for pair-tractable copulas such as Archimedean, Archimax and related hierarchical extensions. Journal of Statistical Computation and Simulation, 93(13):2321–2355, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Ghosh:2020:BIH

[Gho20] Indranil Ghosh. Bayesian inference for hidden truncation Pareto (IV) models. Journal of Statistical Computation and Simulation, 90(12):2136–2155, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghosh:2024:RAV

[GJP24] Abhik Ghosh, María Jaenada, and Leandro Pardo. Robust adaptive variable selection in ultra-high dimensional linear regression models. Journal of Statistical Computation and Simulation, 94(3):571–603, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2022:AMD

[GJQ22] Jia-Xing Gao, Da-Quan Jiang, and Min-Ping Qian. Adaptive manifold density estimation. Journal of Statistical Computation and Simulation, 92(11):2317–2331, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goyal:2020:TNC

[GK20] Manish Goyal and Narinder Kumar. Two new classes of nonparametric tests for scale parameters. Journal of Statistical Computation and Simulation, 90(17):3093–3105, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

G:2022:PCV

[GL22] Rubén Darío Guevara G. and Tania Alejandra López. Process capability vector for multivariate nonlinear profiles. Journal of Statistical Computation and Simulation, 92(6):1292–1321, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghosh:2022:MID

[GM22] Shyamal Ghosh and Priyanka Majumder. A moment inequality for decreasing mean time to failure distributions with hypothesis testing application. Journal of Statistical Computation and Simulation, 92(14):2875–2890, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garay:2020:BAO

[GMCL20] Aldo M. Garay, Francyelle L. Medina, Celso R. B. Cabral, and Tsung-I Lin. Bayesian analysis of the p-order integer-valued AR process with zero-inflated Poisson innovations. Journal of Statistical Computation and Simulation, 90(11):1943–1964, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gholiabad:2021:MZI

[GMFB21] Somayeh Ghorbani Gholiabad, Abbas Moghimbeigi, Javad Faradmal, and Ahmad Reza Baghestani. A multilevel zeroinflated Conway–Maxwell type negative binomial model for analysing clustered count data. Journal of Statistical Computation and Simulation, 91(9):1762–1781, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomez-Mendez:2023:RMD

[GMJ23] Irving Gómez-Méndez and Emilien Joly. Regression with missing data, a comparison study of techniques based on random forests. Journal of Statistical Computation and Simulation, 93(12):1924–1949, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Geels:2023:TSM

[GPH23] Vincent Geels, Matthew T. Pratola, and Radu Herbei. The taxicab sampler: MCMC for discrete spaces with application to tree models. Journal of Statistical Computation and Simulation, 93(5):753–774, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grabchak:2020:SGT

[Gra20] Michael Grabchak. On the simulation of general tempered stable Ornstein–Uhlenbeck processes. Journal of Statistical Computation and Simulation, 90(6):1057–1081, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghahramani:2020:APT

[GSH20] Masume Ghahramani, Maryam Sharafi, and Reza Hashemi. Analysis of the progressively Type-II right censored data with dependent random removals. Journal of Statistical Computation and Simulation, 90(6):1001–1021, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gohil:2023:MPE

[GSRY23] Kavita Gohil, Hani Samawi, Haresh Rochani, and Lili Yu. Model parameters estimation with non-ignorable missing data using influential exponential tilting resampling approach. Journal of Statistical Computation and Simulation, 93(1):163–174, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2022:IED

[GT22] Yi Gao and Lili Tian. Interval estimation for the difference of two correlated gamma means: a generalized inference method and hybrid methods. Journal of Statistical Computation and Simulation, 92(13):2772–2787, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallaugher:2022:MBC

[GTMP22] Michael P. B. Gallaugher, Salvatore D. Tomarchio, Paul D. McNicholas, and Antonio Punzo. Model-based clustering via skewed matrix-variate cluster-weighted models. Journal of Statistical Computation and Simulation, 92(13):2645–2666, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gasana:2023:ETE

[GvRS23] Emelyne Umunoza Gasana, Dietrich von Rosen, and Martin Singull. An Edgeworth-type expansion for the distribution of a likelihood-based discriminant function. Journal of Statistical Computation and Simulation, 93(17):3185–3202, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2022:SSN

[GW22] Mingyu Gao and Canhong Wen. Subset selection in networklinked data. Journal of Statistical Computation and Simulation, 92(11):2350–2371, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guo:2021:ELA

[GWL21] Yunfei Guo, Tao Wang, and Zhonghua Li. Estimation of the limiting availability of a repairable series system with stationary dependent sequences. Journal of Statistical Computation and Simulation, 91(17):3593–3602, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ge:2022:OBL

[GWNW22] Shufei Ge, Shijia Wang, Farouk S. Nathoo, and Liangliang Wang. Online Bayesian learning for mixtures of spatial spline regressions with mixed effects. Journal of Statistical Computation and Simulation, 92(7):1530–1566, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guo:2024:AHF

[GYL24] Dan Guo, Liang Yan, and Menghan Li. Asymptotically honest fiducial generalized inference: an application in autoregressive models. Journal of Statistical Computation and Simulation, 94 (3):447–459, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2023:NWA

[HA23] Abdul Haq and Anila Anis Abbasi. New weighted adaptive CUSUM charts for monitoring the generalized variance of a bivariate normal process. Journal of Statistical Computation and Simulation, 93(4):604–633, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hong:2020:EEC

[HAC20] Hanwoom Hong, Sung K. Ahn, and Sinsup Cho. Estimation of error correction model with measurement errors. Journal of Statistical Computation and Simulation, 90(9):1661–1680, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:2020:BEM

[Han20] Ming Han. The E-Bayesian estimation and its E-MSE of Pareto distribution parameter under different loss functions. Journal of Statistical Computation and Simulation, 90(10):1834–1848, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:2023:BEP

[Han23] Ming Han. E-Bayesian estimation and E-posterior risk of failure probability under different loss functions and its applications in reliability analysis. Journal of Statistical Computation and Simulation, 93(12):1879–1898, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2020:OST

[Haq20] Abdul Haq. One-sided and two one-sided MEWMA charts for monitoring process mean. Journal of Statistical Computation and Simulation, 90(4):699–718, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2020:CAP

[HAV20] Ahteshamul Haq, Irfan Ali, and Rahul Varshney. Compromise allocation problem in multivariate stratified sampling with flexible fuzzy goals. Journal of Statistical Computation and Simulation, 90(9):1557–1569, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2021:MMS

[HBB21] Abdul Haq, Mehwish Bibi, and Jennifer Brown. Monitoring multivariate simple linear profiles using individual observations. Journal of Statistical Computation and Simulation, 91(17):3573–3592, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2022:MPC

[HBKB22] Abdul Haq, Nazish Bibi, Michael B. C. Khoo, and Jennifer Brown. Monitoring the process coefficient of variation without subgrouping. Journal of Statistical Computation and Simulation, 92(9):1805–1822, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hermanns:2020:EAO

[HCN20] M. Hermanns, E. Cramer, and H. K. T. Ng. EM algorithms for ordered and censored system lifetime data under a proportional hazard rate model. Journal of Statistical Computation and Simulation, 90(18):3301–3337, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2021:MBL

[HCY21] Xiaofang He, Wangxue Chen, and Rui Yang. Modified best linear unbiased estimator of the shape parameter of log-logistic distribution. Journal of Statistical Computation and Simulation, 91(2):383–395, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2021:PAE

[HCZ^{+}21] XueLong Hu, Philippe Castagliola, JianLan Zhong, AnAn Tang, and YuLong Qiao. On the performance of the adaptive EWMA chart for monitoring time between events. Journal of Statistical Computation and Simulation, 91(6):1175–1211, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2022:GFS

[HD22a] Hanji He and Guangming Deng. Grouped feature screening for ultra-high dimensional data for the classification model. Journal of Statistical Computation and Simulation, 92(5):974–997, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hoang:2022:CIV

[HD22b] Anh-Tuan Hoang and Thorsten Dickhaus. Combining independent p-values in replicability analysis: a comparative study. Journal of Statistical Computation and Simulation, 92(10): 2184–2204, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huh:2021:NES

[HDP21] Jib Huh, Derek Dyal, and Cheolwoo Park. Nonparametric estimation of single-index models in scale-space. Journal of Statistical Computation and Simulation, 91(12):2414–2443, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Has:2021:KCS

[HFM21] Sothea Has, Aurélie Fischer, and Mathilde Mougeot. KFC: A clusterwise supervised learning procedure based on the aggregation of distances. Journal of Statistical Computation and Simulation, 91(11):2307–2327, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hancova:2022:PEC

[HGH22] Martina Hancová, Andrej Gajdos, and Jozef Hanc. A practical, effective calculation of gamma difference distributions with open data science tools. Journal of Statistical Computation and Simulation, 92(11):2205–2232, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsieh:2021:SAV

[HH21] I-Chung Hsieh and Yufen Huang. Sensitivity analysis and visualization for functional data. Journal of Statistical Computation and Simulation, 91(8):1593–1615, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2022:SIS

[HJL22] Zhensheng Huang, Zhiqiang Jiang, and Jing Li. Statistical inference for a single-index varying coefficient model with measurement errors in all covariates. Journal of Statistical Computation and Simulation, 92(5):1047–1061, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2021:EAM

[HKLA21] Abdul Haq, Michael B. C. Khoo, Ming Ha Lee, and Saddam Akber Abbasi. Enhanced adaptive multivariate EWMA and CUSUM charts for process mean. Journal of Statistical Computation and Simulation, 91(12):2361–2382, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2021:DAD

[HLC21] Daojiang He, Lei Liu, and Mingxiang Cao. A doubly accelerated degradation model based on the inverse Gaussian process and its objective Bayesian analysis. Journal of Statistical Computation and Simulation, 91(8):1485–1503, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2024:TRH

[HLK24] Mingwei Hu, Mei Li, and Lingchen Kong. Trustworthy regularized Huber regression for outlier detection. Journal of Statistical Computation and Simulation, 94(5):1121–1137, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horn:2023:SDT

[HM23] Melanie Horn and Christine H. Müller. Sign depth tests in multiple regression. Journal of Statistical Computation and Simulation, 93(7):1169–1191, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haque:2022:ZTP

[HMB22] M. Ershadul Haque, Taslim Sazzad Mallick, and Wasimul Bari. Zero truncated Poisson model: an alternative approach for analyzing count data with excess zeros. Journal of Statistical Computation and Simulation, 92(3):476–487, February 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2021:CST

[HMS21] Linhan Hu, Soutrik Mandal, and Samiran Sinha. A comparative study of two-sample tests for interval-censored data. Journal of Statistical Computation and Simulation, 91(18):3894–3916, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2024:SDR

[HNK24] Abdul Haq, Mehwish Naz, and Michael B. C. Khoo. Single and dual reference-free CUSCORE charts for detecting unknown patterned mean shifts. Journal of Statistical Computation and Simulation, 94(3):484–498, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Henderson:2023:BSK

[HPP23] Daniel J. Henderson, Alecos Papadopoulos, and Christopher F. Parmeter. Bandwidth selection for kernel density estimation of fat-tailed and skewed distributions. Journal of Statistical Computation and Simulation, 93(12):2110–2135, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:2023:GAM

[HPV23] Shakhawat Hossain, Sharandeep Pandher, and Andrei Volodin. Generalized autoregressive moving average models: an efficient estimation approach. Journal of Statistical Computation and Simulation, 93(4):556–580, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hunt:2023:BEG

[HPW23] R. Hunt, S. Peiris, and N. Weber. Bayesian estimation of Gegenbauer processes. Journal of Statistical Computation and Simulation, 93(9):1357–1377, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2020:MWA

[HR20] Abdul Haq and Faiqa Razzaq. Maximum weighted adaptive CUSUM charts for simultaneous monitoring of process mean and variance. Journal of Statistical Computation and Simulation, 90(16):2949–2974, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hazeb:2021:NPE

[HRB21] Raja Hazeb, Mohammad Z. Raqab, and Husam Awni Bayoud. Non-parametric estimation of the extropy and the entropy measures based on progressive type-II censored data with testing uniformity. Journal of Statistical Computation and Simulation, 91(11):2178–2210, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2022:PFA

[HSK22] Abdul Haq, Haleema Sadia, and Michael B. C. Khoo. A parameter-free adaptive EWMA chart with variable sample sizes and variable sampling intervals for the process mean. Journal of Statistical Computation and Simulation, 92(13):2802– 2828, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hemerik:2021:PTH

[HTF21] Jesse Hemerik, Magne Thoresen, and Livio Finos. Permutation testing in high-dimensional linear models: an empirical investigation. Journal of Statistical Computation and Simulation, 91 (5):897–914, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2023:GVC

[HTX23] Lixia Hu, Yiming Tang, and Qunfang Xu. Generalized varyingcoefficient additive model for locally stationary time series. Journal of Statistical Computation and Simulation, 93(8):1337– 1354, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hussain:2022:NVT

[Hus22] Abid Hussain. A novel version for three-player gambler’s ruin problem. Journal of Statistical Computation and Simulation, 92(14):2863–2874, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hernandez-Velasco:2020:MES

[HVAVD20] Lina L. Hernandez-Velasco, Carlos A. Abanto-Valle, and Dipak K. Dey. Mixed effects state-space models with studentt errors. Journal of Statistical Computation and Simulation, 90(17):3157–3174, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2020:IDG

[HW20] San-Teng Huang and Wei-Ying Wu. The information detection for the generalized additive model. Journal of Statistical Computation and Simulation, 90(12):2267–2290, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2022:PVE

[HWF22] Shuwen Hu, You-Gan Wang, and Liya Fu. Performance of variance estimators in the analysis of longitudinal data with a large cluster size. Journal of Statistical Computation and Simulation, 92(1):1–18, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2021:ALP

[HYWW21] Yunfei He, Lianqiang Yang, Xuejun Wang, and Shijie Wang. Adaptive local polynomial estimations for heterogeneously variational regression functions. Journal of Statistical Computation and Simulation, 91(3):605–622, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huo:2023:EEV

[HZ23] Junfeng Huo and Xiuqing Zhou. Efficient estimation in varying coefficient regression models. Journal of Statistical Computation and Simulation, 93(13):2297–2320, 2023. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2023:MGP

[HZD23] Jie Huang, Fukang Zhu, and Dianliang Deng. A mixed generalized Poisson INAR model with applications. Journal of Statistical Computation and Simulation, 93(11):1851–1878, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2024:TEW

[HZXS24] XueLong Hu, SuYing Zhang, FuPeng Xie, and Zhi Song. Triple exponentially weighted moving average control charts without or with variable sampling interval for monitoring the coefficient of variation. Journal of Statistical Computation and Simulation, 94(3):536–570, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ibrahim:2023:CVT

[IAO^{+}23] Mohd Salami Ibrahim, Myat Moe Thwe Aung, San San Oo, Mohd Ihsanuddin Abas, and Nur Azree Ferdaus Kamudin. Content validation template (CVT). Journal of Statistical Computation and Simulation, 93(7):1151–1168, 2023. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iqbal:2020:MRR

[IMAI20] Kanwal Iqbal, Muhammad Moeen, Asad Ali, and Anam Iqbal. Mixture regression cum ratio estimators of population mean under stratified random sampling. Journal of Statistical Computation and Simulation, 90(5):854–868, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iqbal:2023:BEM

[IS23a] Azeem Iqbal and Muhammad Yousuf Shad. E-Bayesian estimation of Maxwell distribution and its evaluation standards: E-posterior risks and e-MSEs (expected mean square errors). Journal of Statistical Computation and Simulation, 93(7):1192– 1206, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iqbal:2023:BAH

[IS23b] Azeem Iqbal and Muhammad Yousuf Shad. On E-Bayesian analysis of the hierarchical normal and inverse gamma model using different loss functions and its application. Journal of Statistical Computation and Simulation, 93(11):1743–1771, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Imran:2023:DEM

[ISZ^{+}23] Muhammad Imran, Jinsheng Sun, Fatima Sehar Zaidi, Zameer Abbas, and Hafiz Zafar Nazir. On designing efficient multivariate exponentially weighted moving average control chart for compositional data using variable sample size. Journal of Statistical Computation and Simulation, 93(10):1622–1643, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jantre:2023:BQR

[Jan23] Sanket Jantre. Bayesian quantile regression for longitudinal count data. Journal of Statistical Computation and Simulation, 93(1):103–127, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jabeen:2022:EDC

[JAZ22] Riffat Jabeen, Muhammad Aslam, and Azam Zaka. Effects of different calibration constraints on calibration estimators under the randomized response technique. Journal of Statistical Computation and Simulation, 92(10):1995–2017, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jeyadurga:2022:DVR

[JB22] P. Jeyadurga and Saminathan Balamurali. Designing of variables repetitive group sampling plan for consumer protection. Journal of Statistical Computation and Simulation, 92(10): 2141–2155, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jerbi:2020:ECP

[JC20] Yacin Jerbi and Samira Chaabene. European call price modelling using neural networks in considering volatility as stochastic with comparison to the Heston model. Journal of Statistical Computation and Simulation, 90(10):1793–1810, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2020:CVS

[JF20] Hangjin Jiang and Xiaodan Fan. A consistent variable screening procedure with family-wise error control. Journal of Statistical Computation and Simulation, 90(7):1267–1285, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jana:2022:IEC

[JG22] Nabakumar Jana and Meenakshi Gautam. Interval estimation of the common mean and difference of medians for a bivariate lognormal distribution. Journal of Statistical Computation and Simulation, 92(15):3249–3274, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2024:BEV

[JHW24] Jie Jiang, Tian He, and Lichun Wang. Bayes estimates of variance components in mixed linear model. Journal of Statistical Computation and Simulation, 94(4):744–761, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jia:2021:CDL

[Jia21] Xiang Jia. A comparison of different least-squares methods for reliability of Weibull distribution based on right censored data. Journal of Statistical Computation and Simulation, 91 (5):976–999, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jeong:2022:CBV

[JIK22] Donghoon Jeong, Jongho Im, and Young Min Kim. Cosinebased variable bandwidth selection for nonparametric spectral density estimation under long-range dependence. Journal of Statistical Computation and Simulation, 92(6):1158–1174, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2021:NPF

[JLZ21] Xuejun Jiang, Wei Liu, and Bo Zhang. A note on the prediction of frailties with misspecified shared frailty models. Journal of Statistical Computation and Simulation, 91(2):219–241, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jafarian-Namin:2021:IQM

[JNFTM^{+}21] Samrad Jafarian-Namin, Mohammad Saber Fallahnezhad, Reza Tavakkoli-Moghaddam, Ali Salmasnia, and Seyyed Mohammad Taghi Fatemi Ghomi. An integrated quality, maintenance and production model based on the delayed monitoring under the ARMA control chart. Journal of Statistical Computation and Simulation, 91(13):2645–2669, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Jafarian-Namin:2024:RDA

[JNNTMS24] Samrad Jafarian-Namin, Mohammad Saber Fallah Nezhad, Reza Tavakkoli-Moghaddam, and Ali Salmasnia. Robust design of ARMA and ACC charts for imperfect and autocorrelated processes under uncertainty. Journal of Statistical Computation and Simulation, 94(4):762–786, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jobe:2021:DSN

[JP21] J. Marcus Jobe and Michael Pokojovy. Detecting selected non-random patterns with individuals control charts. Journal of Statistical Computation and Simulation, 91(17):3615–3636, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Joukar:2020:EPL

[JRM20] A. Joukar, M. Ramezani, and S. M. T. K. MirMostafaee. Estimation of P (X > Y) for the power Lindley distribution based on progressively type II right censored samples. Journal of Statistical Computation and Simulation, 90(2):355–389, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ji:2022:CBD

[JS22a] Yonggang Ji and Haifang Shi. Constrained Bayesian doubly elastic net lasso for linear quantile mixed models. Journal of Statistical Computation and Simulation, 92(3):579–609, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jose:2022:SBT

[JS22b] Jitto Jose and E. I. Abdul Sathar. Symmetry being tested through simultaneous application of upper and lower k-records in extropy. Journal of Statistical Computation and Simulation, 92(4):830–846, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jing:2024:RTM

[JS24] Wenbo Jing and Jeffrey S. Simonoff. A regression tree method for longitudinal and clustered data with multivariate responses. Journal of Statistical Computation and Simulation, 94(4):820– 842, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jena:2024:TPR

[JT24] Pravash Jena and Manas Ranjan Tripathy. Testing powers of the ratio of variances of two normal populations with a common mean. Journal of Statistical Computation and Simulation, 94 (5):903–922, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2022:BJE

[JZ22a] Hongyan Jiang and Yichuan Zhao. Bayesian jackknife empirical likelihood for the error variance in linear regression models. Journal of Statistical Computation and Simulation, 92(16): 3400–3413, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2022:TJE

[JZ22b] Hongyan Jiang and Yichuan Zhao. Transformed jackknife empirical likelihood for probability weighted moments. Journal of Statistical Computation and Simulation, 92(8):1618–1639, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2024:AEC

[JZY24] Ruizhe Jiang, Jiujun Zhang, and Zhuoxi Yu. Adaptive EWMA control chart for monitoring two-parameter exponential distribution with type-II right censored data. Journal of Statistical Computation and Simulation, 94(4):787–819, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Keshavarz:2021:RAF

[KAN21] Maryam Keshavarz, Shervin Asadzadeh, and Seyed Taghi Akhavan Niaki. Risk-adjusted frailty-based CUSUM control chart for phase i monitoring of patients’ lifetime. Journal of Statistical Computation and Simulation, 91(2):334–352, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khanam:2021:CMA

[KAQ21] Nisha Khanam, A. H. Ansari, and Abdul Quddoos. Compromise mixed allocation using fuzzy programming approach. Journal of Statistical Computation and Simulation, 91(3):435– 444, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kausar:2023:IDC

[KAQ23] Tehzeeb Kausar, Atif Akbar, and Muhammad Qasim. Influence diagnostics for the Cox proportional hazards regression model: method, simulation and applications. Journal of Statistical Computation and Simulation, 93(10):1580–1600, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kazemi:2024:SVM

[Kaz24] Mohammad Kazemi. Support vector machine in ultrahighdimensional feature space. Journal of Statistical Computation and Simulation, 94(3):517–535, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karadag:2020:HTI

[KB20] Özge Karadag and Sevil Bacanli. Hypothesis testing for the inverse Gaussian distribution mean based on ranked set sampling. Journal of Statistical Computation and Simulation, 90 (13):2384–2394, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2022:ART

[KB22] Nam-Hwui Kim and Ryan P. Browne. Anderson relaxation test for intrinsic dimension selection in model-based clustering. Journal of Statistical Computation and Simulation, 92(16): 3468–3487, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2024:BVH

[KB24] Young Geun Kim and Changryong Baek. Bayesian vector heterogeneous autoregressive modelling. Journal of Statistical Computation and Simulation, 94(6):1139–1157, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2022:CST

[KC22] K. Krishnamoorthy and Saptarshi Chakraberty. Construction of simultaneous tolerance intervals for several normal distributions. Journal of Statistical Computation and Simulation, 92 (1):101–114, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2022:PIE

[KCC22] Nirpeksh Kumar, Subha Chakraborti, and Philippe Castagliola. Phase II exponential charts for monitoring time between events data: performance analysis using exact conditional average time to signal distribution. Journal of Statistical Computation and Simulation, 92(7):1457–1486, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2020:FBI

[KCO20] N. Mamode Khan, Hatice Oncel Cekim, and Gamze Ozel. The family of the bivariate integer-valued autoregressive process (BINAR(1)) with Poisson–Lindley (PL) innovations. Journal of Statistical Computation and Simulation, 90(4):624–637, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2020:IMC

[KD20] Xiaoning Kang and Xinwei Deng. An improved modified Cholesky decomposition approach for precision matrix estimation. Journal of Statistical Computation and Simulation, 90 (3):443–464, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kamranfar:2020:SIR

[KEC20] H. Kamranfar, J. Etminan, and M. Chahkandi. Statistical inference for a repairable system subject to shocks: classical vs. Bayesian. Journal of Statistical Computation and Simulation, 90(1):112–137, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelter:2021:TIE

[Kel21] Riko Kelter. Type I and II error rates of Bayesian two-sample tests under preliminary assessment of normality in balanced and unbalanced designs and its influence on the reproducibility of medical research. Journal of Statistical Computation and Simulation, 91(15):3184–3206, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumari:2024:BLE

[KGK24] Anita Kumari, Indranil Ghosh, and Kapil Kumar. Bayesian and likelihood estimation of multicomponent stress-strength reliability from power Lindley distribution based on progressively censored samples. Journal of Statistical Computation and Simulation, 94(5):923–964, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kara:2022:EPG

[KGSA22] Mahmut Kara, Gamze Güven, Birdal Senoglu, and Halil Aydogdu. Estimation of the parameters of the gamma geometric process. Journal of Statistical Computation and Simulation, 92(12):2525–2535, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2020:MDD

[KH20] C. Satheesh Kumar and S. Harisankar. Modeling over-dispersed datasets using generalized geometric distributions. Journal of Statistical Computation and Simulation, 90(4):606–623, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kawakami:2023:AGS

[KH23] Jun Kawakami and Shintaro Hashimoto. Approximate Gibbs sampler for Bayesian Huberized lasso. Journal of Statistical Computation and Simulation, 93(1):128–162, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kwon:2020:PVS

[KHK20] Sookhee Kwon, Il Do Ha, and Jong-Min Kim. Penalized variable selection in copula survival models for clustered time-to-event data. Journal of Statistical Computation and Simulation, 90(4):657–675, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2022:BAR

[Kim22] Chanmin Kim. Bayesian additive regression trees in spatial data analysis with sparse observations. Journal of Statistical Computation and Simulation, 92(15):3275–3300, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2020:MRS

[KIS20] Zahid Khan, Muhammad Ismail, and Hani Samawi. Mixture ranked set sampling for estimation of population mean and median. Journal of Statistical Computation and Simulation, 90 (4):573–585, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2023:NCF

[KJAK23] Hyoung-Moon Kim, Yu-Hyeong Jang, Barry C. Arnold, and Yu-Kwang Kim. New closed-form efficient estimators for a bivariate Weibull distribution. Journal of Statistical Computation and Simulation, 93(11):1716–1733, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kannan:2022:DMD

[KJB22] G. Kannan, P. Jeyadurga, and S. Balamurali. Determination of multiple deferred state repetitive group sampling plan for life time assurance under Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 92(1):145–158, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2023:BCM

[KJL23] Jaeoh Kim, Seongil Jo, and Kyoungjae Lee. Bayesian computational methods for state-space models with application to SIR model. Journal of Statistical Computation and Simulation, 93(7):1207–1223, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ke:2021:RMD

[KJT21] Rui Ke, Jing Jia, and Changchun Tan. Robust minimum distance estimators for the CARR(1,1) model. Journal of Statistical Computation and Simulation, 91(3):564–580, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kordasiabi:2020:BNE

[KK20] S. Rahnamay Kordasiabi and S. Khazaei. Bayesian nonparametric estimation of bandwidth using mixtures of kernel estimators for length-biased data. Journal of Statistical Computation and Simulation, 90(10):1849–1874, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Katebi:2021:OES

[KK21] Mehdi Katebi and Michael B. C. Khoo. Optimal economic statistical design of combined double sampling and variable sampling interval multivariate T^{2} control charts. Journal of Statistical Computation and Simulation, 91(10):2094–2115, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2022:JIE

[KK22] Sunil Kumar and Sanam Preet Kour. The joint influence of estimation of sensitive variable under measurement error and non-response using ORRT models. Journal of Statistical Computation and Simulation, 92(17):3583–3604, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumari:2023:IRM

[KKK23] Anita Kumari, Shrawan Kumar, and Kapil Kumar. Inference for reliability in a multicomponent stress-strength model from generalized inverted exponential lifetime distribution under progressive first failure censoring. Journal of Statistical Computation and Simulation, 93(6):863–887, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2022:OBV

[KKLK22] Sang Gil Kang, Dal Ho Kim, Woo Dong Lee, and Yongku Kim. Objective Bayesian variable selection in linear regression model. Journal of Statistical Computation and Simulation, 92(6):1133– 1157, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2024:PHS

[KKLL24] Joungyoun Kim, Youngrae Kim, Johan Lim, and Sungim Lee. Post Hotelling’s T -square procedure to identify fault variables. Journal of Statistical Computation and Simulation, 94(1):1–28, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2020:ETS

[KKS20] Kipoong Kim, Jajoon Koo, and Hokeun Sun. An empirical threshold of selection probability for analysis of highdimensional correlated data. Journal of Statistical Computation and Simulation, 90(9):1606–1617, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2023:ERP

[KKZ23] Sunil Kumar, Sanam Preet Kour, and Qi Zhang. An enhanced ratio-cum-product estimator with non-response and observational error by utilizing ORRT models: a sensitive estimation approach. Journal of Statistical Computation and Simulation, 93(5):818–836, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2022:ETH

[KL22a] Xiaojuan Kang and Tizheng Li. Estimation and testing of a higher-order partially linear spatial autoregressive model. Journal of Statistical Computation and Simulation, 92(15):3167– 3201, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2022:RMC

[KL22b] Minjo Kim and Sangyeol Lee. Risk measurement for conditionally heteroscedastic location-scale time series models with ASTD and AEPD innovations. Journal of Statistical Computation and Simulation, 92(8):1673–1695, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2024:REG

[KL24] Byungsoo Kim and Sangyeol Lee. Robust estimation for general integer-valued autoregressive models based on the exponentialpolynomial divergence. Journal of Statistical Computation and Simulation, 94(6):1300–1316, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2023:NCI

[KLC23] K. Krishnamoorthy, Shanshan Lv, and Saptarshi Chakraberty. A new confidence interval for the ratio of two normal means and comparisons. Journal of Statistical Computation and Simulation, 93(5):708–722, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karling:2021:BAE

[KLdS21] M. J. Karling, S. R. C. Lopes, and R. M. de Souza. A Bayesian approach for estimating the parameters of an α-stable distribution. Journal of Statistical Computation and Simulation, 91(9): 1713–1748, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2023:CEM

[KLL^{+}23] Lican Kang, Xuerui Li, Yanyan Liu, Yuan Luo, and Zhikang Zou. A communication-efficient method for l_{0} regularization linear regression models. Journal of Statistical Computation and Simulation, 93(4):533–555, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kavya:2021:SPM

[KM21] P. Kavya and M. Manoharan. Some parsimonious models for lifetimes and applications. Journal of Statistical Computation and Simulation, 91(18):3693–3708, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karczewski:2022:DDK

[KM22a] Maciej Karczewski and Andrzej Michalski. A data-driven kernel estimator of the density function. Journal of Statistical Computation and Simulation, 92(17):3529–3541, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kitani:2022:OSL

[KM22b] Masato Kitani and Hidetoshi Murakami. One-sample location test based on the sign and Wilcoxon signed-rank tests. Journal of Statistical Computation and Simulation, 92(3):610–622, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khammar:2022:DRT

[KN22] Amirhamzeh Khammar and Hadi Alizadeh Noughabi. On differential Renyi’s–Tsallis divergence measure and its applications. Journal of Statistical Computation and Simulation, 92(15):3109–3128, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khorshidi:2023:VVG

[KNM23] H. R. Khorshidi, A. R. Nematollahi, and T. Manouchehri. On the vector-valued generalized autoregressive models. Journal of Statistical Computation and Simulation, 93(14):2428–2449, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuran:2021:IPM

[KÖ21] Özge Kuran and Nimet Özbay. Improving prediction by means of a two parameter approach in linear mixed models. Journal of Statistical Computation and Simulation, 91(18):3721–3743, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kamat:2021:LRA

[KR21] Gauri Kamat and Jerome P. Reiter. Leveraging random assignment to impute missing covariates in causal studies. Journal of Statistical Computation and Simulation, 91(7):1275–1305, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2021:BAV

[KRK^{+}21] Xiaoning Kang, Shyam Ranganathan, Lulu Kang, Julia Gohlke, and Xinwei Deng. Bayesian auxiliary variable model for birth records data with qualitative and quantitative responses. Journal of Statistical Computation and Simulation, 91(16):3283– 3303, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kazemi:2021:IPL

[KSAR21] M. Kazemi, D. Shahsvani, M. Arashi, and P. C. Rodrigues. Identification for partially linear regression model with autoregressive errors. Journal of Statistical Computation and Simulation, 91(7):1441–1454, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2020:OSS

[KSGS20] Zaheen Khan, Javid Shabbir, Sat Gupta, and Amjad Shamim. An optimal systematic sampling scheme. Journal of Statistical Computation and Simulation, 90(11):2023–2036, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2024:BIM

[KSJ24] Naushad Mamode Khan, Yuvraj Sunecher, and Vandna Jowaheer. Bivariate INAR(1) model under negative binomial innovations with non-homogeneous over-dispersed indices and application. Journal of Statistical Computation and Simulation, 94 (4):665–698, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2022:SGL

[KSL22] Sehwan Kim, Qifan Song, and Faming Liang. Stochastic gradient Langevin dynamics with adaptive drifts. Journal of Statistical Computation and Simulation, 92(2):318–336, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kurose:2024:STU

[Kur24] Yuta Kurose. Simulation of truncated and unimodal gamma distributions. Journal of Statistical Computation and Simulation, 94(5):996–1015, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kiran:2023:NMM

[KV23] Garikamukkala Vijaya Kiran and K. Vidhya. Novel multi-media steganography model using meta-heuristic and deep learning assisted adaptive lifting wavelet transform. Journal of Statistical Computation and Simulation, 93(17):3126–3155, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kessentini:2021:MLM

[KZ21] Sameh Kessentini and Raoudha Zine. Maximum likelihood and maximum a posteriori estimators for the Riesz probability distribution. Journal of Statistical Computation and Simulation, 91(14):2854–2879, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Luini:2020:DEM

[LA20] E. Luini and P. Arbenz. Density estimation of multivariate samples using Wasserstein distance. Journal of Statistical Computation and Simulation, 90(2):181–210, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:SRL

[LAW20] Yong Li, Yasin Asar, and Jibo Wu. On the stochastic restricted Liu estimator in logistic regression model. Journal of Statistical Computation and Simulation, 90(15):2766–2788, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Liu:2022:CTE

[LBHK22] W. Liu, F. Bretz, A. J. Hayter, and S. Kiatsupaibul. Computation of tolerance ellipses for bivariate and trivariate normal populations. Journal of Statistical Computation and Simulation, 92(17):3630–3638, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:LIB

[LBHX23] Kai Liu, N. Balakrishnan, Mu He, and Lingfang Xie. Likelihood inference for Birnbaum–Saunders frailty model with an application to bone marrow transplant data. Journal of Statistical Computation and Simulation, 93(13):2158–2175, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2020:MIC

[LC20] J. Lee and J. Chen. A modified information criterion for tuning parameter selection in 1d fused LASSO for inference on multiple change points. Journal of Statistical Computation and Simulation, 90(8):1496–1519, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2020:SDA

[LCH20] Ge Liu, Peter F. Craigmile, and Radu Herbei. A study of the data augmentation strategy for stochastic differential equations. Journal of Statistical Computation and Simulation, 90 (10):1753–1772, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:ETV

[LCP23] Shaowen Liu, Massimiliano Caporin, and Sandra Paterlini. Estimating time-varying proximity with a state-space model. Journal of Statistical Computation and Simulation, 93(14): 2534–2562, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2022:ESI

[LCY22] Ming-Che Lu, Dong-Shang Chang, and Su-Fen Yang. Exact statistical inferences for the median of the Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 92(3):567–578, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lai:2020:RMF

[LCZ^{+}20] Peng Lai, Yuanxing Chen, Jie Zhang, Bingying Dai, and Qingzhao Zhang. Robust model-free feature screening for ultrahigh dimensional surrogate data. Journal of Statistical Computation and Simulation, 90(3):550–569, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2021:LES

[LE21] Jianxuan Liu and Sara Eftekharnejad. Locally efficient semiparametric estimator for zero-inflated Poisson model with errorprone covariates. Journal of Statistical Computation and Simulation, 91(6):1092–1107, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2020:LSB

[Lee20] Sangyeol Lee. Location and scale-based CUSUM test with application to autoregressive models. Journal of Statistical Computation and Simulation, 90(13):2309–2328, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2022:SAR

[Lee22] Minjung Lee. Semiparametric analysis of recurrent discrete time data with competing risks. Journal of Statistical Computation and Simulation, 92(16):3301–3316, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:PPQ

[LG20] Tizheng Li and Yue Guo. Penalized profile quasi-maximum likelihood method of partially linear spatial autoregressive model. Journal of Statistical Computation and Simulation, 90(15): 2705–2740, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lesany:2021:EOS

[LG21] S. A. Lesany and S. M. T. Fatemi Ghomi. Extraction and organization of statistical distribution functions for simulation of variations and patterns in the variability control charts. Journal of Statistical Computation and Simulation, 91(12):2383–2401, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:NPB

[LGK20] Qing Li, Feng Guo, and Inyoung Kim. A non-parametric Bayesian change-point method for recurrent events. Journal of Statistical Computation and Simulation, 90(16):2929–2948, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:BMS

[LGM20] Qiong Li, Xin Gao, and Hélène Massam. Bayesian model selection approach for coloured graphical Gaussian models. Journal of Statistical Computation and Simulation, 90(14):2631–2654, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:ESP

[Li20] Ning Li. Efficient sparse portfolios based on composite quantile regression for high-dimensional index tracking. Journal of Statistical Computation and Simulation, 90(8):1466–1478, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2021:SCI

[Li21] Bo Li. Simultaneous confidence intervals of estimable functions based on quasi-likelihood in generalized linear models for overdispersed data. Journal of Statistical Computation and Simulation, 91(1):108–127, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2024:ADW

[Li24] Bo Li. An alternative derivation of weak convergence concerning quasi-likelihood estimation with a small-sample correction for simultaneous testing. Journal of Statistical Computation and Simulation, 94(4):862–890, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2023:REB

[LJ23] Sangyeol Lee and Minyoung Jo. Robust estimation for bivariate integer-valued autoregressive models based on minimum density power divergence. Journal of Statistical Computation and Simulation, 93(17):3156–3184, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:EPM

[LJW23] Xiufang Liu, Hao Jiang, and Dehui Wang. Estimation of parameters in the MDDRCINAR(p) model. Journal of Statistical Computation and Simulation, 93(6):983–1010, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2022:MPC

[LK22] Sangyeol Lee and Chang Kyeom Kim. Monitoring parameter change for time series models with application to location-scale heteroscedastic models. Journal of Statistical Computation and Simulation, 92(18):3885–3916, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2023:MVI

[LK23] Sangyeol Lee and Dongwon Kim. Multiple values-inflated time series of counts: modeling and inference based on INGARCH scheme. Journal of Statistical Computation and Simulation, 93(8):1297–1317, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2021:MIC

[LKS21] Sangyeol Lee, Dongwon Kim, and Seongwoo Seok. Modeling and inference for counts time series based on zero-inflated exponential family INGARCH models. Journal of Statistical Computation and Simulation, 91(11):2227–2248, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2021:RRR

[LL21] Xuhua Liu and Na Li. Redrawing-resampling rejection controlled sequential importance sampling. Journal of Statistical Computation and Simulation, 91(12):2345–2360, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:AIM

[LLBW23] Congmin Liu, Chang Liu, Yansong Bai, and Dehui Wang. Asymptotic inference for moderate deviations from a unit root of nearly unstable INAR(1) processes. Journal of Statistical Computation and Simulation, 93(8):1318–1336, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Lee:2023:CPT

[LLM23] Sangjo Lee, Sangyeol Lee, and Koichi Maekawa. Change point test for structural vector autoregressive model via independent component analysis. Journal of Statistical Computation and Simulation, 93(5):687–707, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2020:SMB

[LLMY20] Shenglin Lin, Wei Li, Ping Ma, and Ming Yang. Structural modelling and Bayesian inference for model validation and confidence extrapolation. Journal of Statistical Computation and Simulation, 90(2):211–233, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:ECV

[LLP^{+}23] Yuzi Liu, Xiaohui Liu, Yingying Pan, Jiming Jiang, and Peiwen Xiao. An empirical comparison of various MSPE estimators and associated prediction intervals for small area means. Journal of Statistical Computation and Simulation, 93(10):1532–1558, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lai:2020:FSQ

[LLWZ20] Peng Lai, Weijuan Liang, Fangjian Wang, and Qingzhao Zhang. Feature screening of quadratic inference functions for ultrahigh dimensional longitudinal data. Journal of Statistical Computation and Simulation, 90(14):2614–2630, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2024:IID

[LLX24] Yujian Liu, Yazhe Li, and Dejun Xie. Implications of imbalanced datasets for empirical ROC-AUC estimation in binary classification tasks. Journal of Statistical Computation and Simulation, 94(1):183–203, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2020:UDF

[LLZ20] Shishi Liu, Xiangjie Li, and Jingxiao Zhang. Ultrahigh dimensional feature screening for additive model with multivariate response. Journal of Statistical Computation and Simulation, 90 (5):775–799, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Long:2022:EPT

[Lon22] Bing Long. Estimation and prediction for Topp–Leone distribution using double Type-I hybrid censored data. Journal of Statistical Computation and Simulation, 92(14):2999–3019, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2022:ELB

[LRN22] Mei Li, Suthakaran Ratnasingam, and Wei Ning. Empiricallikelihood-based confidence intervals for quantile regression models with longitudinal data. Journal of Statistical Computation and Simulation, 92(12):2536–2553, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2021:OGF

[LS21] Sangyeol Lee and Byungtae Seo. Omnibus goodness of fit test based on quadratic distance. Journal of Statistical Computation and Simulation, 91(18):3771–3791, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2021:BSA

[LSL21] Yi Li, Sumi Seo, and Kyu Ha Lee. Bayesian survival analysis using gamma processes with adaptive time partition. Journal of Statistical Computation and Simulation, 91(14):2937–2952, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lasinio:2020:CRP

[LSM20] Giovanna Jona Lasinio, Mario Santoro, and Gianluca Mastrantonio. CircSpaceTime: an R package for spatial and spatiotemporal modelling of circular data. Journal of Statistical Computation and Simulation, 90(7):1315–1345, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2023:CCP

[LT23] Chung-I Li and Meng-Rong Tsai. Control charts for profile monitoring of within-profile correlations using the Tweedie exponential dispersion process model. Journal of Statistical Computation and Simulation, 93(4):513–532, 2023. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Li:2020:PGE

[LTL20] Song Li, Geoffrey K. F. Tso, and Jin Li. Parallel generalized elliptical slice sampling with adaptive regional pseudopriors. Journal of Statistical Computation and Simulation, 90(15):2789–2813, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:RMC

[LTS^{+}23] Xiaoping Liu, Xiao Tian, Jian Shi, Shunhai Xu, Linying Bai, and Shaoping Wang. Reliability modelling considering multidimensional cognitive uncertainties based on uncertainty theory. Journal of Statistical Computation and Simulation, 93(10): 1496–1514, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lodhi:2021:IGF

[LTW21] Chandrakant Lodhi, Yogesh Mani Tripathi, and Liang Wang. Inference for a general family of inverted exponentiated distributions with partially observed competing risks under generalized progressive hybrid censoring. Journal of Statistical Computation and Simulation, 91(12):2503–2526, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2023:PEG

[LTW23a] ShiXiang Li, Yubin Tian, and Dianpeng Wang. Parallel efficient global optimization by using the minimum energy criterion. Journal of Statistical Computation and Simulation, 93(17):3104–3125, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2023:RAC

[LTW^{+}23b] Zhixian Lin, Longlong Tao, Shaoxuan Wang, Chunhua Chen, Jianye Wang, and Daochuan Ge. Reliability analysis of cold-standby systems using a self-stratifying subset simulation method. Journal of Statistical Computation and Simulation, 93(13):2257–2275, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2024:CIG

[LTY^{+}24] Pengyi Liu, Guo-Liang Tian, Kam Chuen Yuen, Yuan Sun, and Chi Zhang. Compositional inverse Gaussian models with applications in compositional data analysis with possible zero observations. Journal of Statistical Computation and Simulation, 94(2):248–278, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2021:CIP

[LW21a] Chung-Han Lee and Hsiuying Wang. Confidence intervals for the proportion of conformance. Journal of Statistical Computation and Simulation, 91(8):1564–1579, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2021:BBA

[LW21b] Bohan Li and Juan Wu. Bayesian bootstrap adaptive lasso estimators of regression models. Journal of Statistical Computation and Simulation, 91(8):1651–1680, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2022:NEM

[LW22] Nan Lu and Lihong Wang. A nonparametric estimation method for the multivariate mixture models. Journal of Statistical Computation and Simulation, 92(17):3727–3742, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2023:MPH

[LW23] Jiang Li and Jingjing Wu. Minimum profile Hellinger distance estimation of semiparametric multiple linear regression models. Journal of Statistical Computation and Simulation, 93(14): 2494–2511, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2020:EME

[LWG20] Wanbo Lu, Yanfeng Wang, and Yuxuan Gao. M-estimates for the multiplicative error model. Journal of Statistical Computation and Simulation, 90(1):1–27, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:SSR

[LWS20] Yan Li, Guoqiao Wang, and Jeff M. Szychowski. Sample size re-estimation for confirmatory two-stage flexible multi-arm trial with normal outcomes. Journal of Statistical Computation and Simulation, 90(1):157–178, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2023:STD

[LWS^{+}23] Yonghui Liu, Jing Wang, Dawei Shi, Víctor Leiva, and Shuangzhe Liu. A score test for detecting extreme values in a vector autoregressive model. Journal of Statistical Computation and Simulation, 93(15):2751–2779, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyles:2020:CBC

[LWW20] Robert H. Lyles, Paul Weiss, and Lance A. Waller. Calibrated Bayesian credible intervals for binomial proportions. Journal of Statistical Computation and Simulation, 90(1):75–89, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2020:SGM

[LWWD20] Yufeng Lin, Yuehua Wu, Xiaogang Wang, and Hao Ding. A segmented generalized Markov regime-switching model with its application in financial time series data. Journal of Statistical Computation and Simulation, 90(5):839–853, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2023:FDC

[LXC23] Xinyu Li, Jianjun Xu, and Haoyang Cheng. Functional data clustering via information maximization. Journal of Statistical Computation and Simulation, 93(16):2982–3007, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2020:REC

[LXH20] Fei Lu, Liugen Xue, and Yuping Hu. Robust estimation for the correlation matrix of multivariate longitudinal data. Journal of Statistical Computation and Simulation, 90(13):2473–2496, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:VSP

[LYP20] Tizheng Li, Qingyan Yin, and Jialong Peng. Variable selection of partially linear varying coefficient spatial autoregressive model. Journal of Statistical Computation and Simulation, 90(15):2681–2704, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:CPD

[LYZ20] Qing Li, Kehui Yao, and Xinyu Zhang. A change-point detection and clustering method in the recurrent-event context. Journal of Statistical Computation and Simulation, 90(6):1131– 1149, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2020:SDF

[LZ20] Xiangjie Li and Jingxiao Zhang. Sufficient dimension folding via tensor inverse regression. Journal of Statistical Computation and Simulation, 90(8):1413–1429, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2023:WMB

[LZCF23] Jingmao Li, Qingzhao Zhang, Song Chen, and Kuangnan Fang. Weighted multiple blockwise imputation method for high-dimensional regression with blockwise missing data. Journal of Statistical Computation and Simulation, 93(3):459–474, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2022:IAH

[LZZ22] Xiang Li, Yi Zhang, and Jun Zhao. An improved algorithm for high-dimensional continuous threshold expectile model with variance heterogeneity. Journal of Statistical Computation and Simulation, 92(8):1590–1617, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Majid:2022:LEB

[MAA22] Abdul Majid, Muhammad Amin, and Muhammad Nauman Akram. On the Liu estimation of Bell regression model in the presence of multicollinearity. Journal of Statistical Computation and Simulation, 92(2):262–282, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahmoudi:2023:ERB

[Mah23] Mohammad Reza Mahmoudi. Evaluating the relationship between two periodically correlated processes with Mandelbrot– Van Ness fractional Brownian motion errors using periodic copula. Journal of Statistical Computation and Simulation, 93(1): 46–57, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meraou:2022:ASD

[MAKRK22] Mohammed A. Meraou, Noriah M. Al-Kandari, Mohammad Z. Raqab, and Debasis Kundu. Analysis of skewed data by using compound Poisson exponential distribution with applications to insurance claims. Journal of Statistical Computation and Simulation, 92(5):928–956, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marozzi:2021:CBT

[Mar21] Marco Marozzi. A combined bootstrap test for the two-sample location problem. Journal of Statistical Computation and Simulation, 91(1):180–196, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mousavi:2023:RPA

[MAY23] Niloufar Dousti Mousavi, Hani Aldirawi, and Jie Yang. An R package AZIAD for analysing zero-inflated and zero-altered data. Journal of Statistical Computation and Simulation, 93 (16):2801–2827, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Makhdoom:2022:LCB

[MB22] Iman Makhdoom and Mehdi Basikhasteh. Location charts based on concomitants of ranked set samples from Morgenstern family. Journal of Statistical Computation and Simulation, 92 (5):911–927, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moon:2023:BEN

[MB23] Chul Moon and Adel Bedoui. Bayesian elastic net based on empirical likelihood. Journal of Statistical Computation and Simulation, 93(11):1669–1693, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahmoudi:2020:CPT

[MBM20] Akram Mahmoudi, Reza Arabi Belaghi, and Saumen Mandal. A comparison of preliminary test, Stein-type and penalty estimators in gamma regression model. Journal of Statistical Computation and Simulation, 90(17):3051–3079, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Mohan:2021:EPK

[MC21] Rakhi Mohan and Manoj Chacko. Estimation of parameters of Kumaraswamy-exponential distribution based on adaptive type-II progressive censored schemes. Journal of Statistical Computation and Simulation, 91(1):81–107, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martinez-Camblor:2022:FRD

[MC22] Pablo Martínez-Camblor. The fundamental role of density functions in the binary classification problem. Journal of Statistical Computation and Simulation, 92(13):2846–2861, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Michael:2022:CBP

[MD22] Courtney Vanderford Michael and Xin Dang. Correlations in bivariate Pareto distributions. Journal of Statistical Computation and Simulation, 92(12):2501–2524, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohammed:2023:BIM

[MG23] Noura Ait Mohammed and Hafida Guerbyenne. Bayesian inference in a multiple contaminated autoregressive model with trend. Journal of Statistical Computation and Simulation, 93 (12):2067–2109, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mermi:2021:MPR

[MGA21] Selman Mermi, Atila Göktas, and Özge Akkus. Are most proposed ridge parameter estimators skewed and do they have any effect on MSE values? Journal of Statistical Computation and Simulation, 91(10):2074–2093, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Magalhaes:2021:IPE

[MGB21] Tiago M. Magalhães, Diego I. Gallardo, and Marcelo Bourguignon. Improved point estimation for inverse gamma regression models. Journal of Statistical Computation and Simulation, 91(12):2444–2456, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahadik:2021:TSS

[MGT21] Shashibhushan B. Mahadik, Dadasaheb G. Godase, and W. L. Teoh. A two-sided SPRT control chart for process dispersion. Journal of Statistical Computation and Simulation, 91 (17):3603–3614, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moslehi:2021:NFS

[MH21] Fateme Moslehi and Abdorrahman Haeri. A novel feature selection approach based on clustering algorithm. Journal of Statistical Computation and Simulation, 91(3):581–604, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mullah:2023:CEA

[MHB23] Muhammad Abu Shadeque Mullah, Zakir Hossain, and Andrea Benedetti. Comparing estimation approaches for generalized additive mixed models with binary outcomes. Journal of Statistical Computation and Simulation, 93(18):3362–3381, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maged:2022:TBE

[MHKS22] Ahmed Maged, Salah Haridy, Michael B. C. Khoo, and Mohammad Shamsuzzaman. Time between events monitoring for imperfect maintained systems with application to a robotic system. Journal of Statistical Computation and Simulation, 92(7): 1347–1372, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mirkamali:2021:IEW

[Mir21] Sayed Jamal Mirkamali. An improved exponentially weighted moving average chart for monitoring proportions using maxima nomination sampling. Journal of Statistical Computation and Simulation, 91(2):282–299, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martinez:2021:CEA

[MJG21] Gustavo-Ignacio Rivas Martínez and María-Dolores Jiménez-Gamero. Computationally efficient approximations for independence tests in non-parametric regression. Journal of Statistical Computation and Simulation, 91(6):1134–1154, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Milosevic:2021:QRP

[MJGAF21] Bojana Milosević, M. Dolores Jiménez-Gamero, and M. Virtudes Alba-Fernández. Quantifying the ratio-plot for the geometric distribution. Journal of Statistical Computation and Simulation, 91(11):2153–2177, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mulder:2020:BIM

[MJK20] Kees Mulder, Pieter Jongsma, and Irene Klugkist. Bayesian inference for mixtures of von Mises distributions using reversible jump MCMC sampler. Journal of Statistical Computation and Simulation, 90(9):1539–1556, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murugeswari:2023:EOD

[MJSB23] N. Murugeswari, P. Jeyadurga, S. Sridevi, and S. Balamurali. Evaluation and optimal designing of a new system of skip-lot reinspection schemes. Journal of Statistical Computation and Simulation, 93(9):1420–1440, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meselidis:2020:SIM

[MK20] Christos Meselidis and Alex Karagrigoriou. Statistical inference for multinomial populations based on a double index family of test statistics. Journal of Statistical Computation and Simulation, 90(10):1773–1792, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Malekzadeh:2020:SCI

[MKK20] Ahad Malekzadeh and Mahmood Kharrati-Kopaei. Simultaneous confidence intervals for the quantile differences of several two-parameter exponential distributions under the progressive type II censoring scheme. Journal of Statistical Computation and Simulation, 90(11):2037–2056, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehmood:2020:GSC

[MLA^{+}20] Rashid Mehmood, Muhammad Hisyam Lee, Ambreen Afzel, Sheraz Bashir, and Muhammad Riaz. Generalized skewness correction structure of X̄ control chart for unknown process parameters and skewed probability distributions. Journal of Statistical Computation and Simulation, 90(8):1349–1372, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ma:2023:GFO

[MLG23] Xiaoyu Ma, Lu Lin, and Yujie Gai. A general framework of online updating variable selection for generalized linear models with streaming datasets. Journal of Statistical Computation and Simulation, 93(3):325–340, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Misaei:2020:AMC

[MMH20] Hasan Misaei, Samaneh Eftekhari Mahabadi, and Firoozeh Haghighi. Analysis of masked competing risks data with cause and time dependent masking mechanism. Journal of Statistical Computation and Simulation, 90(12):2256–2266, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Misaii:2022:MIM

[MMH22] Hasan Misaii, Samaneh Eftekhari Mahabadi, and Firoozeh Haghighi. Multiple imputation of masked competing risks data using machine learning algorithms. Journal of Statistical Computation and Simulation, 92(16):3317–3342, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mojirsheibani:2021:SCI

[MN21] Majid Mojirsheibani and My-Nhi Nguyen. On statistical classification with incomplete covariates via filtering. Journal of Statistical Computation and Simulation, 91(7):1342–1365, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Manouchehri:2022:PMA

[MN22] T. Manouchehri and A. R. Nematollahi. PAR(1) model analysis: a web-based shiny application for analysing periodic autoregressive models. Journal of Statistical Computation and Simulation, 92(10):2090–2111, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maekawa:2023:ENG

[MN23] Koichi Maekawa and Tadashi Nakanishi. Estimation of non-Gaussian SVAR models: a pseudo-log-likelihood function approach. Journal of Statistical Computation and Simulation, 93(11):1830–1850, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahdavi-Nasab:2022:NMR

[MNAP22] Narges Mahdavi-Nasab, Mostafa Abouei Ardakan, and Abdossaber Peiravi. A new model for the reliability-redundancy allocation problem with the mixed redundancy strategy. Journal of Statistical Computation and Simulation, 92(14):2956–2979, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Modak:2022:NNI

[Mod22] Soumita Modak. A new nonparametric interpoint distancebased measure for assessment of clustering. Journal of Statistical Computation and Simulation, 92(5):1062–1077, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maji:2020:RWT

[MP20] Avijit Maji and Leandro Pardo. Robust Wald-type test statistics based on minimum C-divergence estimators. Journal of Statistical Computation and Simulation, 90(14):2592–2613, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mojirsheibani:2021:NNB

[MP21] Majid Mojirsheibani and William Pouliot. A nearest-neighbor-based ensemble classifier and its large-sample optimality. Journal of Statistical Computation and Simulation, 91(10):2034– 2050, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marquez:2020:EEV

[MRNLB20] Hector Gomez Marquez, Adrian Ramirez-Nafarrate, and Rafael López-Bracho. Estimation of the expected value of the arrival epochs given a bounded period of observation for the homogeneous and non-homogeneous Poisson processes. Journal of Statistical Computation and Simulation, 90(17):3250–3259, 2020.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohammadi:2022:ZOI

[MSBS22] Z. Mohammadi, Z. Sajjadnia, H. S. Bakouch, and M. Sharafi. Zero-and-one inflated Poisson–Lindley INAR(1) process for modelling count time series with extra zeros and ones. Journal of Statistical Computation and Simulation, 92(10):2018–2040, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murtaza:2020:IOR

[MSH20] Maryam Murtaza, Sarjinder Singh, and Zawar Hussain. An innovative optimal randomized response model using correlated scrambling variables. Journal of Statistical Computation and Simulation, 90(15):2823–2839, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maleki:2022:PIM

[MSMTK22] Mohammad Reza Maleki, Ali Salmasnia, Edgard M. Maboudou-Tchao, and Parnaz Khanbeygi. Phase II monitoring of logistic regression profiles with estimated parameters. Journal of Statistical Computation and Simulation, 92(13):2721–2739, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mustafa:2023:NEW

[MSR23] Fakhar Mustafa, Rehan Ahmad Khan Sherwani, and Muhammad Ali Raza. A new exponentially weighted moving average control chart to monitor count data with applications in healthcare and manufacturing. Journal of Statistical Computation and Simulation, 93(18):3308–3328, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mollica:2020:PPR

[MT20] Cristina Mollica and Luca Tardella. PLMIX: an R package for modelling and clustering partially ranked data. Journal of Statistical Computation and Simulation, 90(5):925–959, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marthin:2023:RNN

[MT23] Pius Marthin and N. Ata Tutkun. Recurrent neural network for complex survival problems. Journal of Statistical Computation and Simulation, 93(13):2232–2256, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahto:2022:SIG

[MTSR22] Amulya Kumar Mahto, Yogesh Mani Tripathi, Farha Sultana, and Manoj Kumar Rastogi. Statistical inference for a Gumbel type-II distribution under hybrid censoring. Journal of Statistical Computation and Simulation, 92(11):2290–2316, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahto:2021:SIB

[MTW21] Amulya Kumar Mahto, Yogesh Mani Tripathi, and Shuo-Jye Wu. Statistical inference based on progressively type-II censored data from the Burr X distribution under progressivestress accelerated life test. Journal of Statistical Computation and Simulation, 91(2):368–382, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meilan-Vila:2022:GFT

[MVFFC22] A. Meilán-Vila, M. Francisco-Fernández, and R. M. Crujeiras. Goodness-of-fit tests for multiple regression with circular response. Journal of Statistical Computation and Simulation, 92(9):1941–1963, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Miao:2024:REF

[MW24] Zi Miao and Lihong Wang. Robust estimation for functionon-scalar regression models. Journal of Statistical Computation and Simulation, 94(5):1035–1055, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maleki:2020:AHT

[MWMCR20] Mohsen Maleki, Darren Wraith, Mohammad R. Mahmoudi, and Javier E. Contreras-Reyes. Asymmetric heavy-tailed vector auto-regressive processes with application to financial data. Journal of Statistical Computation and Simulation, 90(2):324– 340, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ma:2023:ECC

[MWS23] Yuqing Ma, Peijie Wang, and Jianguo Sun. Estimation of complier causal treatment effects under the case-cohort studies with interval-censored failure time data. Journal of Statistical Computation and Simulation, 93(18):3285–3307, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ma:2022:MFF

[MXYY22] Weidong Ma, Jingsong Xiao, Ying Yang, and Fei Ye. Model-free feature screening for ultrahigh dimensional data via a Pearson chi-square based index. Journal of Statistical Computation and Simulation, 92(15):3222–3248, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mondal:2023:TSA

[MYM23] Anirban Mondal, Kai Yin, and Abhijit Mandal. A two-stage adaptive Metropolis algorithm. Journal of Statistical Computation and Simulation, 93(7):1130–1150, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Najmi:2021:ORR

[NAJ21] Ali Najmi, Mostafa Abouei Ardakan, and Youness Javid. Optimization of reliability redundancy allocation problem with component mixing and strategy selection for subsystems. Journal of Statistical Computation and Simulation, 91(10):1935–1959, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadeem:2023:CIC

[NAV23] Hira Nadeem, S. Ejaz Ahmed, and Andrei Volodin. Confidence intervals for the cross product ratio under the special case of direct-inverse sampling scheme and its applications. Journal of Statistical Computation and Simulation, 93(17):3203–3216, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nitithumbundit:2023:MLO

[NC23] Thanakorn Nitithumbundit and Jennifer S. K. Chan. Maximum leave-one-out likelihood method for the location parameter of variance gamma distribution with unbounded density. Journal of Statistical Computation and Simulation, 93(15):2642–2671, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadeb:2024:SIP

[NET^{+}24] Hossein Nadeb, Javad Estabraqi, Hamzeh Torabi, Yichuan Zhao, and Saeede Bafekri. Statistical inference for the partial area under ROC curve for the lower truncated proportional hazard rate models based on progressive type-II censoring. Journal of Statistical Computation and Simulation, 94(5):965–995, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Novais:2023:VSU

[NF23] Luísa Novais and Susana Faria. Variable selection using the EM and CEM algorithms in mixtures of linear mixed models. Journal of Statistical Computation and Simulation, 93(13): 2196–2231, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2021:ETO

[NFP21] M. Shafaei Noughabi and A. M. Franco-Pereira. Estimation of two ordered quantile residual life functions based on mixtures. Journal of Statistical Computation and Simulation, 91 (18):3792–3813, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Naqvi:2023:IPG

[NI23] Syed Yasir Abbas Naqvi and Zahid Iqbal. To improve the performance of genetic algorithms by using a novel selection operator. Journal of Statistical Computation and Simulation, 93(17): 3067–3081, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nikoloulopoulos:2020:WSE

[Nik20] Aristidis K. Nikoloulopoulos. Weighted scores estimating equations and CL1 information criteria for longitudinal ordinal response. Journal of Statistical Computation and Simulation, 90(11):2002–2022, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nikoloulopoulos:2022:BFN

[Nik22] Aristidis K. Nikoloulopoulos. The bivariate K-finite normal mixture ‘blanket’ copula. Journal of Statistical Computation and Simulation, 92(6):1224–1245, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nishida:2023:KDE

[Nis23] Kiheiji Nishida. Kernel density estimation by genetic algorithm. Journal of Statistical Computation and Simulation, 93(8):1263– 1281, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2020:NPD

[NJ20] Hadi Alizadeh Noughabi and Jalil Jarrahiferiz. Nonparametric probability density functions of entropy estimators applied to testing the Rayleigh distribution. Journal of Statistical Computation and Simulation, 90(14):2537–2551, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nguyen:2022:AUC

[NJR22] Thuan Nguyen, Jiming Jiang, and J. Sunil Rao. Assessing uncertainty for classified mixed model prediction. Journal of Statistical Computation and Simulation, 92(2):249–261, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nishino:2020:GMC

[NM20] Takuya Nishino and Hidetoshi Murakami. The generalized multisample Cucconi test statistic for the location and scale parameters. Journal of Statistical Computation and Simulation, 90(12):2291–2305, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nikita:2022:SMS

[NN22] Efthymia Nikita and Panos Nikitas. Simulation methods for squared Euclidean and Mahalanobis type distances for multivariate data and their application in assessing the uncertainty in hierarchical clustering. Journal of Statistical Computation and Simulation, 92(11):2403–2424, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2023:NEK

[NN23a] Hadi Alizadeh Noughabi and Mohammad Shafaei Noughabi. A new estimator of Kullback–Leibler information based on a local linear regression and its application in goodness-of-fit tests. Journal of Statistical Computation and Simulation, 93(16):2828–2842, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2023:VEA

[NN23b] Hadi Alizadeh Noughabi and Mohammad Shafaei Noughabi. Varentropy estimators with applications in testing uniformity. Journal of Statistical Computation and Simulation, 93(15): 2582–2599, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nascimento:2022:VAH

[NS22] Mauricio Nascimento and Benjamin A. Shaby. A Vecchia approximation for high-dimensional Gaussian cumulative distribution functions arising from spatial data. Journal of Statistical Computation and Simulation, 92(9):1977–1994, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Naqvi:2021:NLD

[NSK21] Fakhra Batool Naqvi, Muhammad Yousaf Shad, and Saima Khan. A new logistic distribution based crossover operator for real-coded genetic algorithm. Journal of Statistical Computation and Simulation, 91(4):817–835, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oliveira:2020:IDM

[OCNN20] José V. Oliveira Jr., Francisco Cribari-Neto, and Juvêncio S. Nobre. Influence diagnostics and model validation for the generalized extreme-value nonlinear regression model. Journal of Statistical Computation and Simulation, 90(3):515–549, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Olmez:2021:PKM

[OKF^{+}21] Fatih Olmez, Peter R. Kramer, John Fricks, Deena R. Schmidt, and Janet Best. Penalized KS method to fit data sets with power law distribution over a bounded subinterval. Journal of Statistical Computation and Simulation, 91(8):1524–1563, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ostadi:2020:QTS

[OSG20] Bakhtiar Ostadi, Zahra Honarmand Shahzileh, and Aboutaleb Gerami. A quantitative target-setting model using Monte Carlo simulation method: mapping in strategic management and balanced scorecard (BSC) context. Journal of Statistical Computation and Simulation, 90(13):2421–2437, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Orbe:2021:SSP

[OV21] Jesus Orbe and Jorge Virto. Selecting the smoothing parameter and knots for an extension of penalized splines to censored data. Journal of Statistical Computation and Simulation, 91(14):2953–2985, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozgul:2021:NIC

[Ozg21] Nilgun Ozgul. New improved calibration estimator based on two auxiliary variables in stratified two-phase sampling. Journal of Statistical Computation and Simulation, 91(6):1243– 1256, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pekalp:2023:CMT

[PA23] Mustafa Hilmi Pekalp and Halil Aydogdu. Comparison of monotonic trend tests for some counting processes. Journal of Statistical Computation and Simulation, 93(8):1282–1296, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pandolfo:2022:GPD

[Pan22] Giuseppe Pandolfo. The GLD-plot: a depth-based graphical tool to investigate unimodality of directional data. Journal of Statistical Computation and Simulation, 92(11):2372–2385, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2021:BII

[Par21] Jaewoo Park. Bayesian indirect inference for models with intractable normalizing functions. Journal of Statistical Computation and Simulation, 91(2):300–315, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Punzo:2021:MTI

[PB21] Antonio Punzo and Luca Bagnato. The multivariate tailinflated normal distribution and its application in finance. Journal of Statistical Computation and Simulation, 91(1):1–36, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Poudyal:2023:FSP

[PB23] Chudamani Poudyal and Vytaras Brazauskas. Finite-sample performance of the T - and W -estimators for the Pareto tail index under data truncation and censoring. Journal of Statistical Computation and Simulation, 93(10):1601–1621, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Papayiannis:2021:CUS

[PDD^{+}21] G. I. Papayiannis, G. N. Domazakis, D. Drivaliaris, S. Koukoulas, A. E. Tsekrekos, and A. N. Yannacopoulos. On clustering uncertain and structured data with Wasserstein barycenters and a geodesic criterion for the number of clusters. Journal of Statistical Computation and Simulation, 91(13):2569–2594, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perez-Elizalde:2022:HSP

[PEMCPRC22] Sergio Pérez-Elizalde, Blanca E. Monroy-Castillo, Paulino Pérez-Rodríguez, and José Crossa. HDBRR: a statistical package for high-dimensional Bayesian ridge regression without MCMC. Journal of Statistical Computation and Simulation, 92(17):3679–3705, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pho:2020:PPA

[PFMT20] Kim-Hung Pho, Zhenshuai Fu, Mohammad Reza Mahmoudi, and Bui Anh Tuan. Prediction for the processes with almost cyclostationary structure. Journal of Statistical Computation and Simulation, 90(15):2814–2822, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pal:2022:MPG

[PG22] Subhadip Pal and Jeremy Gaskins. Modified Pólya–Gamma data augmentation for Bayesian analysis of directional data.

Journal of Statistical Computation and Simulation, 92(16): 3430–3451, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Phong:2023:ADF

[Pho23] Duong Thanh Phong. Approximation for the distribution function of non-central complex Wilks statistic. Journal of Statistical Computation and Simulation, 93(17):3034–3051, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parker:2021:GBM

[PHW21] Paul A. Parker, Scott H. Holan, and Skye A. Wills. A general Bayesian model for heteroskedastic data with fully conjugate full-conditional distributions. Journal of Statistical Computation and Simulation, 91(15):3207–3227, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parpoula:2022:OSP

[PK22a] Christina Parpoula and Alex Karagrigoriou. On optimal segmentation and parameter tuning for multiple change-point detection and inference. Journal of Statistical Computation and Simulation, 92(18):3789–3816, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Patil:2022:USE

[PK22b] K. P. Patil and H. V. Kulkarni. An uniformly superior exact multi-sample test procedure for homogeneity of variances under location-scale family of distributions. Journal of Statistical Computation and Simulation, 92(18):3931–3957, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Prajapati:2023:OBS

[PMKP23] Deepak Prajapati, Sharmishtha Mitra, Debasis Kundu, and Ayan Pal. Optimal Bayesian sampling plan for censored competing risks data. Journal of Statistical Computation and Simulation, 93(5):775–799, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Popovic:2022:CED

[PMW22] Bozidar V. Popović, Andjela Mijanović, and Viktor Witkovský. Computing the exact distribution of a linear combination of generalized logistic random variables and its applications. Journal of Statistical Computation and Simulation, 92(5):1015– 1033, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Puggard:2022:CIV

[PNN22] Wisunee Puggard, Sa-Aat Niwitpong, and Suparat Niwitpong. Confidence intervals for the variance and difference of variances of Birnbaum–Saunders distributions. Journal of Statistical Computation and Simulation, 92(13):2829–2845, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2024:RFD

[PP24] Hoyoung Park and Junyong Park. A robust false discovery rate controlling procedure using the empirical likelihood with a fast algorithm. Journal of Statistical Computation and Simulation, 94(5):1097–1120, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perdikis:2023:ESC

[PPC^{+}23] Theodoros Perdikis, Stelios Psarakis, Philippe Castagliola, Vicent Giner-Bosch, Petros E. Maravelakis, and Athanasios C. Rakitzis. An EWMA sign chart for dispersion with exact run length properties. Journal of Statistical Computation and Simulation, 93(11):1799–1829, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perdikis:2021:ETC

[PPCC21] T. Perdikis, S. Psarakis, P. Castagliola, and G. Celano. An EWMA-type chart based on signed ranks with exact run length properties. Journal of Statistical Computation and Simulation, 91(4):732–751, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pakzad:2022:DLB

[PRG22] A. Pakzad, H. Razavi, and B. Sadeghpour Gildeh. Developing loss-based functional process capability indices for simple linear profile. Journal of Statistical Computation and Simulation, 92 (1):115–144, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pinto:2021:CME

[PS21] Vimukthini Pinto and Roshini Sooriyarachchi. Comparison of methods of estimation for a goodness of fit test — an analytical and simulation study. Journal of Statistical Computation and Simulation, 91(9):1846–1866, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pushadapu:2024:CMO

[PS24] Kavya Pushadapu and Sarjinder Singh. Chaudhuri and Mukerjee ORRT for two sensitive characteristics and their overlap. Journal of Statistical Computation and Simulation, 94(5): 1056–1072, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Petterle:2023:UGM

[PTdS^{+}23] Ricardo R. Petterle, César A. Taconeli, José L. P. da Silva, Guilherme P. da Silva, Henrique A. Laureano, and Wagner H. Bonat. Unit gamma mixed regression models for continuous bounded data. Journal of Statistical Computation and Simulation, 93(6):1011–1029, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pedroso:2021:EBR

[PTG21] Vinicius César Pedroso, Cesar Augusto Taconeli, and Suely Ruiz Giolo. Estimation based on ranked set sampling for the twoparameter Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 91(2):316–333, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Panwar:2024:MRB

[PY24] M. S. Panwar and C. P. Yadav. Modelling recall-based competing risks data with k different models. Journal of Statistical Computation and Simulation, 94(4):722–743, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Piriaei:2020:BEC

[PYF20] Hassan Piriaei, Gholamhossein Yari, and Rahman Farnoosh. E-Bayesian estimations for the cumulative hazard rate and mean residual life based on exponential distribution and record data.

Journal of Statistical Computation and Simulation, 90(2):271– 290, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2023:HDE

[PZWL23] Yingli Pan, Xiaoluo Zhao, Sha Wei, and Zhan Liu. Highdimensional expectile regression incorporating graphical structure among predictors. Journal of Statistical Computation and Simulation, 93(2):231–248, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qasim:2022:RGR

[QAAM22] Muhammad Qasim, Muhammad Nauman Akram, Muhammad Amin, and Kristofer Månsson. A restricted gamma ridge regression estimator combining the gamma ridge regression and the restricted maximum likelihood methods of estimation. Journal of Statistical Computation and Simulation, 92(8):1696–1713, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qiao:2021:CMC

[QH21] Lei Qiao and Dong Han. CUSUM multi-chart based on nonparametric likelihood approach for detecting unknown abrupt changes and its application for network data. Journal of Statistical Computation and Simulation, 91(17):3473–3491, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quinino:2020:CCM

[QHCB20] R. C. Quinino, L. L. Ho, F. R. B. Cruz, and L. F. Bessegato. A control chart to monitor the process mean based on inspecting attributes using control limits of the traditional X-bar chart. Journal of Statistical Computation and Simulation, 90(9):1639– 1660, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qin:2020:RVA

[QJFY20] Chunying Qin, Shuang Jia, Xingwu Fang, and Xiaofeng Yu. Relationship validation among items and attributes. Journal of Statistical Computation and Simulation, 90(18):3360–3375, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Queiroz:2020:HTI

[QL20] Francisco F. Queiroz and Artur J. Lemonte. On hypothesis testing inference in location-scale models under model misspecification. Journal of Statistical Computation and Simulation, 90(11):2080–2097, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qian:2022:CMI

[QL22] Lianyong Qian and Qi Li. A class of max-INAR(1) processes with explanatory variables. Journal of Statistical Computation and Simulation, 92(9):1898–1919, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qasim:2021:SBR

[QMK21] Muhammad Qasim, Kristofer Månsson, and B. M. Golam Kibria. On some beta ridge regression estimators: method, simulation and application. Journal of Statistical Computation and Simulation, 91(9):1699–1712, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qi:2022:ELM

[QZ22] Yongcheng Qi and Yingchao Zhou. Empirical likelihood method for complete independence test on high-dimensional data. Journal of Statistical Computation and Simulation, 92(11):2386– 2402, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodriguez:2020:ICC

[RANF20] William D. A. Rodríguez, Getúlio J. A. Amaral, Abraão D. C. Nascimento, and Jodavid A. Ferreira. Information criteria in classification: new divergence-based classifiers. Journal of Statistical Computation and Simulation, 90(17):3200–3217, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raquel:2021:ZMP

[RCPA21] Gabriela C. Raquel, Katiane S. Conceição, Marcos O. Prates, and Marinho G. Andrade. A zero-modified Poisson mixed model with generalized random effect. Journal of Statistical Computation and Simulation, 91(12):2457–2474, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Roy:2021:TMI

[RDSG21] Angshuman Roy, Kunal Das, Soham Sarkar, and Anil K. Ghosh. Tests of mutual independence among several random vectors using univariate and multivariate ranks of nearest neighbours. Journal of Statistical Computation and Simulation, 91(9):1890–1906, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rumsey:2021:FMA

[RH21] Kellin N. Rumsey and Gabriel Huerta. Fast matrix algebra for Bayesian model calibration. Journal of Statistical Computation and Simulation, 91(7):1331–1341, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahma:2022:DBW

[RK22] Abid Rahma and Célestin C. Kokonendji. Discriminating between and within (semi)continuous classes of both Tweedie and geometric Tweedie models. Journal of Statistical Computation and Simulation, 92(4):794–812, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rachdi:2020:FTP

[RLACE20] Mustapha Rachdi, Ali Laksaci, Ibrahim M. Almanjahie, and Zouaoui Chikr-Elmezouar. FDA: theoretical and practical efficiency of the local linear estimation based on the k NN smoothing of the conditional distribution when there are missing data. Journal of Statistical Computation and Simulation, 90(8):1479– 1495, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramezani:2021:IVB

[RM21] S. Ramezani and M. Mohammadpour. Integer-valued bilinear time series model with signed generalized power series thinning operator. Journal of Statistical Computation and Simulation, 91(2):242–259, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramos:2021:BAI

[RMF^{+}21] Pedro L. Ramos, Alex L. Mota, Paulo H. Ferreira, Eduardo Ramos, Vera L. D. Tomazella, and Francisco Louzada. Bayesian analysis of the inverse generalized gamma distribution using objective priors. Journal of Statistical Computation and Simulation, 91(4):786–816, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozbeh:2024:RCA

[RMM24] M. Roozbeh, M. Maanavi, and N. A. Mohamed. A robust counterpart approach for the ridge estimator to tackle outlier effect in restricted multicollinear regression models. Journal of Statistical Computation and Simulation, 94(2):279–296, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodas:2021:PYS

[RMN21] Paul Corral Rodas, Isabel Molina, and Minh Nguyen. Pull your small area estimates up by the bootstraps. Journal of Statistical Computation and Simulation, 91(16):3304–3357, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roldan-Nofuentes:2021:CMS

[RN21] J. A. Roldán-Nofuentes. Computational methods to simultaneously compare the predictive values of two diagnostic tests with missing data: EM–SEM algorithms and multiple imputation. Journal of Statistical Computation and Simulation, 91(16):3358–3384, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roldan-Nofuentes:2020:ESA

[RNSR20] J. A. Roldán-Nofuentes and S. B. Sidaty-Regad. EM and SEM algorithms to compare the weighted kappa coefficients of two diagnostic tests in the presence of partial verification and discrete covariates. Journal of Statistical Computation and Simulation, 90(18):3454–3476, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ross:2022:ASF

[Ros22] Joseph Ross. Asymmetric scale functions for t-digests. Journal of Statistical Computation and Simulation, 92(1):191–210, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Risser:2020:BIH

[RT20] Mark D. Risser and Daniel Turek. Bayesian inference for highdimensional nonstationary Gaussian processes. Journal of Statistical Computation and Simulation, 90(16):2902–2928, 2020.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saha:2024:FTT

[SA24] Aritra Saha and M. Z. Anis. A family of tests for trend change in failure rate function with right censored data. Journal of Statistical Computation and Simulation, 94(6):1191–1203, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sami:2022:MOP

[SAA^{+}22] Faiza Sami, Muhammad Amin, Muhammad Nauman Akram, Muhammad Moeen Butt, and Bushra Ashraf. A modified one parameter Liu estimator for Conway–Maxwell Poisson response model. Journal of Statistical Computation and Simulation, 92(12):2448–2466, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Silan:2021:EIP

[SAB21] M. Silan, B. Arpino, and G. Boccuzzo. Evaluating inverse propensity score weighting in the presence of many treatments. An application to the estimation of the neighbourhood effect. Journal of Statistical Computation and Simulation, 91(4):836– 859, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sabahno:2023:AMT

[Sab23] Hamed Sabahno. An adaptive max-type multivariate control chart by considering measurement errors and autocorrelation. Journal of Statistical Computation and Simulation, 93(16):2956–2981, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sogandi:2021:PRA

[SAMA21] Fatemeh Sogandi, Majid Aminnayeri, Adel Mohammadpour, and Amirhossein Amiri. Phase I risk-adjusted Bernoulli chart in multistage healthcare processes based on the state-space model. Journal of Statistical Computation and Simulation, 91(3):522– 542, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sangnawakij:2023:ACI

[San23] Patarawan Sangnawakij. Alternative confidence interval estimation for the mean and coefficient of variation in a twoparameter exponential distribution. Journal of Statistical Computation and Simulation, 93(16):2936–2955, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salmasnia:2020:JOD

[SAR20] Ali Salmasnia, Behnam Abdzadeh, and Aram Rahimi. Joint optimisation of double warning T^{2}-Hotelling chart and maintenance policy with multiple assignable causes. Journal of Statistical Computation and Simulation, 90(3):465–488, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sangnawakij:2020:STD

[SB20] Patarawan Sangnawakij and Dankmar Böhning. A simple test for the difference of means in meta-analysis when study-specific variances are unreported. Journal of Statistical Computation and Simulation, 90(14):2524–2536, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sami:2023:TPE

[SBA23] Faiza Sami, Muhammad Moeen Butt, and Muhammad Amin. Two parameter estimators for the Conway–Maxwell–Poisson regression model. Journal of Statistical Computation and Simulation, 93(13):2137–2157, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2021:RFC

[SBAC21] W. Sun, H. F. Bindele, A. Abebe, and H. E. Correia. Robust functional coefficient selection for the single-index varying coefficients regression model. Journal of Statistical Computation and Simulation, 91(8):1681–1697, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2021:SLS

[SBB21] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. Some logarithmic and sine-type imputation techniques for missing data in survey sampling in the presence of measurement errors. Journal of Statistical Computation and Simulation, 91 (4):713–731, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sangnawakij:2021:END

[SBH21] Patarawan Sangnawakij, Dankmar Böhning, and Heinz Holling. On the exact null-distribution of a test for homogeneity of the risk ratio in meta-analysis of studies with rare events. Journal of Statistical Computation and Simulation, 91(2):420–434, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samawi:2021:RSR

[SC21] Hani M. Samawi and Ding-Geng Chen. Ranked simulated resampling: a more efficient and accurate resampling approximations for bootstrap inference. Journal of Statistical Computation and Simulation, 91(18):3709–3720, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sabahno:2020:AVP

[SCA20] Hamed Sabahno, Philippe Castagliola, and Amirhossein Amiri. An adaptive variable-parameters scheme for the simultaneous monitoring of the mean and variability of an autocorrelated multivariate normal process. Journal of Statistical Computation and Simulation, 90(8):1430–1465, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saini:2021:ESS

[SCG21] Shubham Saini, Ajit Chaturvedi, and Renu Garg. Estimation of stress-strength reliability for generalized Maxwell failure distribution under progressive first failure censoring. Journal of Statistical Computation and Simulation, 91(7):1366–1393, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Son:2023:QRC

[SCK23] Dongjae Son, Sangbum Choi, and Sangwook Kang. Quantile regression for competing risks data from stratified case-cohort studies: an induced-smoothing approach. Journal of Statistical Computation and Simulation, 93(8):1225–1243, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Sultana:2024:ESS

[SÇK24] Farha Sultana, Çagatay Çetinkaya, and Debasis Kundu. Estimation of the stress-strength parameter under two-sample balanced progressive censoring scheme. Journal of Statistical Computation and Simulation, 94(6):1269–1299, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simola:2021:ABC

[SCKW21] Umberto Simola, Jessi Cisewski-Kehe, and Robert L. Wolpert. Approximate Bayesian computation for finite mixture models. Journal of Statistical Computation and Simulation, 91(6):1155– 1174, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2021:STS

[SCW21] Si-Lian Shen, Jian-Ling Cui, and Xin-Qian Wu. A simple test for spatial heteroscedasticity in spatially varying coefficient models. Journal of Statistical Computation and Simulation, 91(8):1580–1592, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2020:RFS

[SDML20] Jinhui Sun, Pang Du, Hongyu Miao, and Hua Liang. Robust feature screening procedures for single and mixed types of data. Journal of Statistical Computation and Simulation, 90(7):1173– 1193, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Silva:2024:CSL

[SDP24] H. P. T. N. Silva, G. S. Dissanayake, and T. S. G. Peiris. Comparison of standard long memory time series. Journal of Statistical Computation and Simulation, 94(4):891–901, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sevinc:2023:ACA

[SG23] Büsra Sevinç and Selma Gürler. An adaptive clustering algorithm by neighbourhood search for large-scale data. Journal of Statistical Computation and Simulation, 93(1):175–187, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shirazi:2023:WBT

[SGY23] Esmaeil Shirazi, Bahareh Ghanbari, and Masoud Yarmohammadi. Wavelet block thresholding for copula density estimation under biased sampling. Journal of Statistical Computation and Simulation, 93(14):2512–2533, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Susam:2023:WIT

[SH23] Selim Orhun Susam and Burcu Hudaverdi. A weighted independence test based on smooth estimation of Kendall distribution. Journal of Statistical Computation and Simulation, 93(17):3082–3103, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRA

[Sha20a] Ramalingam Shanmugam. Book review: Analysis of Integrated Data, edited by Li-Chun Zhang and Raymond L. Chambers, Boca Raton, CRC Press, 2019, pp. 256 + xvi, $87.70, ISBN 13: 978-1-4987-2798-3 (hardcover). Journal of Statistical Computation and Simulation, 90(15):2858, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRDb

[Sha20b] Ramalingam Shanmugam. Book review: Data Visualization and Statistical Literacy for Open and Big Data, by Theodosia Prodromou, 2017, IGI Global, Information Science Reference (an imprint of IGI Global), 701 E. Chocolate Avenue, Hershey PA, USA 17033, ISBN: 978-1-5225-2512-7 (hardcover), pp. 365 + ix, $164.00. Journal of Statistical Computation and Simulation, 90(15):2858–2859, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRDa

[Sha20c] Ramalingam Shanmugam. Book review: Discover probability: how to use it, how to avoid misusing it, and how it affects every aspect of your life, by Arieh Ben-Naim, Singapore, World Scientific Press, 2015, 318 + xix pp., $37.37 (paperback), ISBN: 978-981-4616-32-4. Journal of Statistical Computation and Simulation, 90(1):179, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRE

[Sha20d] Ramalingam Shanmugam. Book review: Essentials of Excel, Excel VBA, SAS, and Minitab for statistical and financial analysis, by Cheng-Few Lee, John Lee, John-Ran Chang and Tzu Tai, Chan, Springer Verlag Press, 2016, 1041 + xx pp., $109.00 (paperback), ISBN: 978-3-319-388865-6. Journal of Statistical Computation and Simulation, 90(1):179–180, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRHd

[Sha20e] Ramalingam Shanmugam. Book review: Handbook of environmental and ecological statistics, edited by Alan E. Gelfand, Montse Fuentes, Jennifer A. Hoeting, and Richard L. Smith, Boca Raton, FL, CRC Press, 2019, 853 + xxi pp., $208.94, ISBN-13: 978-1-498-75202-2 (hardcover). Journal of Statistical Computation and Simulation, 90(18):3478, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRHa

[Sha20f] Ramalingam Shanmugam. Book review: Handbook of Item Response Theory: Volume one, Models. Journal of Statistical Computation and Simulation, 90(10):1922, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRHc

[Sha20g] Ramalingam Shanmugam. Book review: Handbook of item response theory: volume three, applications. Journal of Statistical Computation and Simulation, 90(12):2306–2307, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRHb

[Sha20h] Ramalingam Shanmugam. Book review: Handbook of item response theory: volume two, statistical tools. Journal of Statistical Computation and Simulation, 90(12):2306, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BUV

[Sha20i] Ramalingam Shanmugam. Books of uncertainty versus statistics: Do Dice Play God? The Mathematics of Uncertainty, by Ian Stewart, New York, NY, Basic Books, Hachette Book Press, 2019, 293 + ix pp., $17.99 (hardcover), ISBN: 978-1-5416-9947-2. The Art of Statistics: How to Learn from Data, by David Spiegelhalter, New York, NY, Basic Books, Hachette Book Press, 2019, 426 + xvi pp., $18.99 (hardcover), ISBN: 978-1-5416-1851-0. Journal of Statistical Computation and Simulation, 90(18):3477, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:ISS

[Sha20j] Ramalingam Shanmugam. Innovative strategies, statistical solutions, and simulations for modern clinical trials. Journal of Statistical Computation and Simulation, 90(4):772, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:MPG

[Sha20k] Ramalingam Shanmugam. Mathematical programming and game theory. Journal of Statistical Computation and Simulation, 90(3):570, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:MCB

[Sha20l] Ramalingam Shanmugam. Measuring crime: behind the statistics. Journal of Statistical Computation and Simulation, 90 (4):772–773, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:MBG

[Sha20m] Ramalingam Shanmugam. Model-based geostatistics for global public health: methods and applications. Journal of Statistical Computation and Simulation, 90(6):1150, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:MAE

[Sha20n] Ramalingam Shanmugam. Multivariate analysis of ecological data with ade4. Journal of Statistical Computation and Simulation, 90(6):1150–1151, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:PTA

[Sha20o] Ramalingam Shanmugam. Practical text analytics: maximizing the value of text data. Journal of Statistical Computation and Simulation, 90(7):1346, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:SAE

[Sha20p] Ramalingam Shanmugam. Small area estimation and microsimulation modeling. Journal of Statistical Computation and Simulation, 90(7):1346–1347, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:SDS

[Sha20q] Ramalingam Shanmugam. Statistical data science. Journal of Statistical Computation and Simulation, 90(9):1733, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:SHC

[Sha20r] Ramalingam Shanmugam. Statistics and health care fraud: how to save billions. Journal of Statistical Computation and Simulation, 90(3):571, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:TSS

[Sha20s] Ramalingam Shanmugam. Theory of sampling and sampling practice. Journal of Statistical Computation and Simulation, 90(9):1733–1734, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:URM

[Sha20t] Ramalingam Shanmugam. Understanding research methods: an overview of the essentials. Journal of Statistical Computation and Simulation, 90(10):1921–1922, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shang:2021:DBV

[Sha21a] Han Lin Shang. Double bootstrapping for visualizing the distribution of descriptive statistics of functional data. Journal of Statistical Computation and Simulation, 91(10):2116–2132, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:BRA

[Sha21b] Ramalingam Shanmugam. Book review: Analysis of Incidence Rates, by Peter Cummings, 2019, CRC Press, 6000 Broken Sound Parkway NW, Suite 300, Boca Raton, FL 33487, ISBN: 978-0-367-15206-2 (hardcover), pp. xvii + 473, $131.49. Journal of Statistical Computation and Simulation, 91(1):216, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:BRMb

[Sha21c] Ramalingam Shanmugam. Book review: Modelling spatial and spatial-temporal data: a Bayesian approach, by Robert Haining and Guangquan Li, 2019, CRC Press, Boca Raton, FL, ISBN: 978-1-4822-3742-9 (hardcover), pp. 608 + xxvi, $129.95. Journal of Statistical Computation and Simulation, 91(7):1480– 1481, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:BRMa

[Sha21d] Ramalingam Shanmugam. Book review: Multilevel modeling using R, by W. Holmes Finch, Jocelyn E. Bolin and Ken Kelley, 6000 Broken Sound Parkway NW, Suite 300, Boca Raton, FL 33487, CRC Press, 2014, xiii + 216 pp., $50.99 (paperback), ISBN 13: 978-1-4665-1585-7. Journal of Statistical Computation and Simulation, 91(1):216–217, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:BRS

[Sha21e] Ramalingam Shanmugam. Book review: Seeing through statistics, by Jessica M. Utts, 200 First Samford Place, 4th Floor, Stamford, CT 06902, Cengage Learning Press, 2020, pp. 630 + xxiv, $30.99 (Paperback), ISBN-13: 978-1-285-05088-1. Journal of Statistical Computation and Simulation, 91(7):1480, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:OAS

[Sha21f] Ramalingam Shanmugam. Omic association studies with R and Bioconductor. Journal of Statistical Computation and Simulation, 91(3):648–649, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:RAL

[Sha21g] Ramalingam Shanmugam. Recent advances in lifetime and reliability models. Journal of Statistical Computation and Simulation, 91(5):1057, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:RFE

[Sha21h] Ramalingam Shanmugam. Review of four excellent undergraduate statistics books. Journal of Statistical Computation and Simulation, 91(5):1057–1059, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:RTE

[Sha21i] Ramalingam Shanmugam. Review of three excellent time series books. Journal of Statistical Computation and Simulation, 91(9):1907–1908, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:RTH

[Sha21j] Ramalingam Shanmugam. Review of three health/medical statistics books. Journal of Statistical Computation and Simulation, 91(9):1908–1909, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:SAM

[Sha21k] Ramalingam Shanmugam. The statistical analysis of multivariate failure time data: a marginal modeling approach. Journal of Statistical Computation and Simulation, 91(3):648, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2022:BRL

[Sha22a] Ramalingam Shanmugam. Book review: Longitudinal data Analysis of longitudinal data with examples, by You-Gan Wang, Liya Fu, and Sudhir Paul, 2022, Boca Raton, FL, CRC Press, 225 + xxv pp., $99.95 (hardcover), ISBN: 978-1-4987-6460-5. Journal of Statistical Computation and Simulation, 92(14): 3085, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2022:BRB

[Sha22b] Ramalingam Shanmugam. Book review of three excellent books on R language for statistics: 1. Stochastic processes with R: an introduction, by Olga Korosteleva, 2022, Boca Raton, FL, CRC Press, 190 + ix pp., $93.22 (hardcover), ISBN: 978-1-138-59152-3; 2 Statistical methods for mediation, confounding and moderation analysis using R and SAS, by Qingzhao Yu and Bin Li, 2022, Boca Raton, FL, CRC Press, 277 + xv pp., $144.49 (hardcover), ISBN: 978-0-367-36547-9; 3 Spatial predictive modeling with R, by Jin Li, 2022, Boca Raton, FL, CRC Press, 383 + xix pp., $118.10 (hardcover), ISBN: 978-0-367-55054-7. Journal of Statistical Computation and Simulation, 92(12):2642–2643, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRBb

[Sha23a] Ramalingam Shanmugam. Book review: Book on modern statistical methods for health research: Modern statistical methods for health research, edited by, Zhao, Y., Chen, D. G., 2021, Springer Press, 496 + xxv pp., $159.99 (hardback), ISBN: 978-3-030-72439-9, Springer Nature Switzerland AG. Part of Springer Nature. Journal of Statistical Computation and Simulation, 93(8):1355, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRA

[Sha23b] Ramalingam Shanmugam. Book review: Algorithmic trading and quantitative strategies by Raja Velu, Maxence Hardy, and Daniel Nehren, Boca Raton, FL, Chapman and Hall, 2020, CRC Financial Mathematics Series, 434 + xvi pp., $94.75 (hardback), ISBN: 978-1-4987-3716-6. Journal of Statistical Computation and Simulation, 93(9):1478, 2023. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRBa

[Sha23c] Ramalingam Shanmugam. Book review: Bayesian modeling of spatio-temporal data with R by Sujit K. Sahu, 2022, Chapman and Hall, CRC Press, 411 + xxii pp., $74.21 (hardback), ISBN: 978-0-367-27798-7, 6000 Broken Sound Parkway, NW, (Suite 300), Boca Raton, FL 33487. Journal of Statistical Computation and Simulation, 93(7):1224, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRBc

[Sha23d] Ramalingam Shanmugam. Book review: Book on stochastic processes with R: an introduction by Olga Korosteleva. (2022). Stochastic Processes with R: an Introduction, Boca Raton, FL: Chapman and Hall, CRC Press, 2022, x + 190 pp., $99.95 (hardback), ISBN 978-1-032-15373-5., pp. 190 + ix, $ 99.95 (hardback). Journal of Statistical Computation and Simulation, 93(10):1667, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRR

[Sha23e] Ramalingam Shanmugam. Book review: Review of Bayesian books in Bayesian statistics. 1. Bayesian modeling of spatialtemporal data with R, by Sujit K. Sahu, 2022, Boca Raton, FL, CRC Press, 411 + xxii pp., $110.00 (hardcover), ISBN: 978-0-367-27798-7. 2. Bayesian implementation, by Thomas R. Palfrey and Sanjay Srivastava, 2022, Boca Raton, FL, CRC Press, 105 + viii pp., $244.76 (hardcover), ISBN 13: 978-3-718-65314-0. Journal of Statistical Computation and Simulation, 93 (4):684–685, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRC

[Sha23f] Ramalingam Shanmugam. Book review: Causal inference: Artificial intelligence and causal inference, by Momiao Xiong, Boca Raton, FL, CRC Press, 2022, 368 + xxv pp., $130.00 (hardcover), ISBN: 978-0-367-85940-4. Journal of Statistical Computation and Simulation, 93(2):324, 2023. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Siray:2020:PCE

[Sir20] Gülesen Üstündag Siray. Principal components estimator for measurement error models. Journal of Statistical Computation and Simulation, 90(6):1022–1038, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sridevi:2023:DMD

[SJB23] S. Sridevi, P. Jeyadurga, and S. Balamurali. Determination of modified double sampling plan based on the process capability index. Journal of Statistical Computation and Simulation, 93(9):1441–1460, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2020:SCP

[SK20] Junmo Song and Jiwon Kang. Sequential change point detection in ARMA–GARCH models. Journal of Statistical Computation and Simulation, 90(8):1520–1538, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samanta:2021:BID

[SK21a] Debashis Samanta and Debasis Kundu. Bayesian inference of a dependent competing risk data. Journal of Statistical Computation and Simulation, 91(15):3069–3086, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samanta:2021:MAS

[SK21b] Debashis Samanta and Debasis Kundu. Meta-analysis of a stepstress experiment under Weibull distribution. Journal of Statistical Computation and Simulation, 91(9):1867–1889, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2021:TPC

[SK21c] Junmo Song and Jiwon Kang. Test for parameter change in the presence of outliers: the density power divergence-based approach. Journal of Statistical Computation and Simulation, 91(5):1016–1039, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samanta:2024:CRM

[SK24] Debashis Samanta and Debasis Kundu. On competing risk model under step-stress stage life testing. Journal of Statistical Computation and Simulation, 94(4):699–721, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Selingerova:2021:CPS

[SKH21] Iveta Selingerova, Stanislav Katina, and Ivanka Horova. Comparison of parametric and semiparametric survival regression models with kernel estimation. Journal of Statistical Computation and Simulation, 91(13):2717–2739, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sugasawa:2020:SAE

[SKO20] Shonosuke Sugasawa, Yuki Kawakubo, and Kota Ogasawara. Small area estimation with spatially varying natural exponential families. Journal of Statistical Computation and Simulation, 90(6):1039–1056, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanaullah:2022:GRT

[SKSS22] Aamir Sanaullah, Warda Atiq Khan, Husna Sadaf, and Iram Saleem. Generalized ratio-type and exponential ratio-type estimators for mean in presence of non-response using follow-up procedure based on three-phase and four-phase sampling. Journal of Statistical Computation and Simulation, 92(12):2554– 2595, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shin:2023:RLR

[SL23a] Bokyoung Shin and Seokho Lee. Robust logistic regression with shift parameter estimation. Journal of Statistical Computation and Simulation, 93(15):2625–2641, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2023:REL

[SL23b] Huihui Sun and Qiang Liu. Robust empirical likelihood inference for partially linear varying coefficient models with longitudinal data. Journal of Statistical Computation and Simulation, 93(10):1559–1579, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sief:2021:ICS

[SLER21] Mohamed Sief, Xinsheng Liu, and Abd El-Raheem M. Abd El-Raheem. Inference for a constant-stress model under progressive type-I interval censored data from the generalized halfnormal distribution. Journal of Statistical Computation and Simulation, 91(15):3228–3253, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salini:2022:CMR

[SLM^{+}22] S. Salini, F. Laurini, G. Morelli, M. Riani, and A. Cerioli. Covariance matrices of S robust regression estimators. Journal of Statistical Computation and Simulation, 92(4):724–747, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2021:CSM

[SM21] Yue Su and P. K. Mwanakatwe. The comparison study of the model selection criteria on the tobit regression model based on the bootstrap sample augmentation mechanisms. Journal of Statistical Computation and Simulation, 91(7):1415–1440, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sandoval-Moreno:2023:IMR

[SMGAV23] Gabriela Sandoval-Moreno, Manuel Galea, and Reinaldo Arellano-Valle. Inference in multivariate regression models with measurement errors. Journal of Statistical Computation and Simulation, 93(12):1997–2025, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simoneau:2020:ODT

[SMWP20] Gabrielle Simoneau, Erica E. M. Moodie, Michael P. Wallace, and Robert W. Platt. Optimal dynamic treatment regimes with survival endpoints: introducing DWSurv in the R package DTRreg. Journal of Statistical Computation and Simulation, 90(16):2991–3008, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2020:EAD

[SMZ20] Zhi Song, Amitava Mukherjee, and Jiujun Zhang. An efficient approach of designing distribution-free exponentially weighted moving average schemes with dynamic fast initial response for joint monitoring of location and scale. Journal of Statistical Computation and Simulation, 90(13):2329–2353, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stewart:2020:CID

[SN20] Patrick Stewart and Wei Ning. Confidence intervals for data containing many zeros and ones based on empirical likelihoodtype methods. Journal of Statistical Computation and Simulation, 90(18):3376–3399, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2020:UML

[SO20] Brajendra C. Sutradhar and Alwell J. Oyet. An unbalanced multidimensional latent effects-based logistic mixed model and GQL estimation for spatial binary data. Journal of Statistical Computation and Simulation, 90(18):3338–3359, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shirozhan:2023:FIV

[SOBC23] Masoumeh Shirozhan, Emmanuel W. Okereke, Hassan S. Bakouch, and Christophe Chesneau. A flexible integer-valued AR(1) process: estimation, forecasting and modeling COVID-19 data. Journal of Statistical Computation and Simulation, 93(9):1461–1477, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sabino:2021:GRO

[SP21] Piergiacomo Sabino and Nicola Cufaro Petroni. Gamma-related Ornstein–Uhlenbeck processes and their simulation*. Journal of Statistical Computation and Simulation, 91(6):1108–1133, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sengupta:2023:MZI

[SR23] Debjit Sengupta and Surupa Roy. Modelling zero inflated and under-reported count data. Journal of Statistical Computation and Simulation, 93(14):2390–2409, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samawi:2022:MFE

[SRG22] Hani Samawi, Haresh Rochani, and Kavita Gohil. Mean functional estimation with non-ignorable missing data using influential exponential tilting resampling approach. Journal of Statistical Computation and Simulation, 92(11):2332–2349, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shukla:2023:BAG

[SRU23] Asmita Shukla, Rakesh Ranjan, and Satyanshu K. Upadhyay. Bayes analysis of the generalized gamma AFT models for left truncated and right censored data. Journal of Statistical Computation and Simulation, 93(12):2026–2051, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saleem:2022:EMS

[SS22] Iram Saleem and Aamir Sanaullah. Estimation of mean of a sensitive variable using efficient exponential-type estimators in stratified sampling. Journal of Statistical Computation and Simulation, 92(2):232–248, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sadeghpour:2020:ESS

[SSN20] Amineh Sadeghpour, Mahdi Salehi, and Ahmad Nezakati. Estimation of the stress-strength reliability using lower record ranked set sampling scheme under the generalized exponential distribution. Journal of Statistical Computation and Simulation, 90(1):51–74, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2020:RRM

[SSS20] Garib Nath Singh, Chandraketu Singh, and Surbhi Suman. Randomized response model to alter the nuisance effect of nonresponse due to stigmatized issues in survey sampling. Journal of Statistical Computation and Simulation, 90(13):2395–2420, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saleem:2022:ECE

[SSS^{+}22] Iram Saleem, Aamir Sanullah, Javid Shabbir, Husna Sadaf, and Riffat Jabeen. Efficient class of estimators for the estimation of the mean of a sensitive variable using multi-auxiliary variables. Journal of Statistical Computation and Simulation, 92(14):2908–2940, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simushkin:2021:PAM

[SSV21] D. S. Simushkin, S. V. Simushkin, and I. N. Volodin. On the d-posterior approach to the multiple testing problem. Journal of Statistical Computation and Simulation, 91(4):651–666, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stadtlander:2021:BRA

[Sta21] Christian T. K.-H. Stadtländer. Book review: Analysis of correlated data with SAS and R (4th ed.), by Mohamed M. Shoukri, Boca Raton, FL, CRC Press, 2018, 497 + xv pp., $139.95 (hardback), ISBN: 978-1-1381-9745-9. Journal of Statistical Computation and Simulation, 91(7):1478–1479, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saini:2022:REM

[STG22] Shubham Saini, Sachin Tomer, and Renu Garg. On the reliability estimation of multicomponent stress-strength model for Burr XII distribution using progressively first-failure censored samples. Journal of Statistical Computation and Simulation, 92 (4):667–704, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2024:RVS

[STM24] Peng Su, Garth Tarr, and Samuel Muller. Robust variable selection under cellwise contamination. Journal of Statistical Computation and Simulation, 94(6):1371–1387, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Siray:2021:DTP

[STÖ21] Gülesen Üstündag Siray, Selma Toker, and Nimet Özbay. Defining a two-parameter estimator: a mathematical programming evidence. Journal of Statistical Computation and Simulation, 91(11):2133–2152, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stojanovic:2022:EAE

[Sto22] Milena S. Stojanović. An EM algorithm for estimation of the parameters of the geometric minification INAR model. Journal of Statistical Computation and Simulation, 92(14):2941–2955, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Susam:2020:GFT

[SU20] Selim Orhun Susam and Burcu Hudaverdi Ucer. A goodness-of-fit test based on Bézier curve estimation of Kendall distribution. Journal of Statistical Computation and Simulation, 90(7):1194– 1215, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sulewski:2021:LMT

[Sul21] Piotr Sulewski. Logarithmic minimum test for independence in three-way contingency table of small sizes. Journal of Statistical Computation and Simulation, 91(13):2780–2799, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sulewski:2022:TMC

[Sul22] Piotr Sulewski. Two methods of conjoint summands of generating bivariate and trivariate normal pseudo-random numbers.

Journal of Statistical Computation and Simulation, 92(8):1714– 1739, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2020:HMM

[SW20] Weizhe Su and Xia Wang. Hidden Markov model in multiple testing on dependent count data. Journal of Statistical Computation and Simulation, 90(5):889–906, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shariati:2023:CWB

[SYA23] Nasim Madah Shariati, Mohammadreza Yaghouti, and Amjad Alipanah. A convergent wavelet-based method for solving linear stochastic differential equations included 1D and 2D noise. Journal of Statistical Computation and Simulation, 93(5):837– 861, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:2024:BHB

[SYG24] Weihua Shi, Tianrui Ye, and Wenhao Gui. The E-Bayesian and hierarchical Bayesian estimations for the reliability analysis of Kumaraswamy generalized distribution based on upper record values. Journal of Statistical Computation and Simulation, 94(6):1339–1370, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salimov:2020:EMV

[SYVV20] Rustem Salimov, Su-Fen Yang, Andrei Volodin, and Igor Volodin. Estimation of mean value of a normal distribution with constraints on the relative error and d-risk. Journal of Statistical Computation and Simulation, 90(7):1286–1300, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2021:CBM

[SZ21] Bing Su and Fukang Zhu. Comparison of BINAR(1) models with bivariate negative binomial innovations and explanatory variables. Journal of Statistical Computation and Simulation, 91(8):1616–1634, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2024:COS

[SZ24] Yimin Shen and Shuhong Zhang. A Cox-optimized survival model based on GrowNet. Journal of Statistical Computation and Simulation, 94(3):624–641, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2021:EBE

[SZS21] Ji Sun, Ying-Ying Zhang, and Ya Sun. The empirical Bayes estimators of the rate parameter of the inverse gamma distribution with a conjugate inverse gamma prior under Stein’s loss function. Journal of Statistical Computation and Simulation, 91(8):1504–1523, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2023:LBM

[SZT^{+}23] Yuan Sun, Shishun Zhao, Guo-Liang Tian, Man-Lai Tang, and Tao Li. Likelihood-based methods for the zero-one-two inflated Poisson model with applications to biomedicine. Journal of Statistical Computation and Simulation, 93(6):956–982, 2023. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tavangar:2020:CRE

[TA20] Mahdi Tavangar and Majid Asadi. Component reliability estimation based on system failure-time data. Journal of Statistical Computation and Simulation, 90(17):3232–3249, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taconeli:2024:DRR

[Tac24] Cesar Augusto Taconeli. Dual-rank ranked set sampling. Journal of Statistical Computation and Simulation, 94(1):29–49, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tak:2021:MFF

[Tak21] Nihat Tak. Meta fuzzy functions based feed-forward neural networks with a single hidden layer for forecasting. Journal of Statistical Computation and Simulation, 91(13):2800–2816, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsagris:2020:HTT

[TAVP20] Michail Tsagris, Abdulaziz Alenazi, Kleio-Maria Verrou, and Nikolaos Pandis. Hypothesis testing for two population means: parametric or non-parametric test? Journal of Statistical Computation and Simulation, 90(2):252–270, 2020. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taavoni:2021:MSM

[TAWL21] M. Taavoni, M. Arashi, Wan-Lun Wang, and Tsung-I Lin. Multivariate t semiparametric mixed-effects model for longitudinal data with multiple characteristics. Journal of Statistical Computation and Simulation, 91(2):260–281, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

terBraak:2022:PVR

[tB22] Cajo J. F. ter Braak. Predictor versus response permutation for significance testing in weighted regression and redundancy analysis. Journal of Statistical Computation and Simulation, 92(10):2041–2059, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsompanaki:2021:LIA

[TC21] E. Tsompanaki and C. Caroni. Local influence analysis of the 2PL IRT model for binary responses. Journal of Statistical Computation and Simulation, 91(7):1455–1477, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taconeli:2022:DWD

[TdL22] Cesar Augusto Taconeli and Idemauro Antonio Rodrigues de Lara. Discrete Weibull distribution: different estimation methods under ranked set sampling and simple random sampling. Journal of Statistical Computation and Simulation, 92(8):1740–1762, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tenreiro:2020:BSK

[Ten20] Carlos Tenreiro. Bandwidth selection for kernel density estimation: a Hermite series-based direct plug-in approach. Journal of Statistical Computation and Simulation, 90(18):3433–3453, December 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tian:2021:IWE

[TG21] Yajie Tian and Wenhao Gui. Inference of weighted exponential distribution under progressively Type-II censored competing risks model with electrodes data. Journal of Statistical Computation and Simulation, 91(16):3426–3452, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tran:2021:MCV

[TH21] P. H. Tran and C. Heuchenne. Monitoring the coefficient of variation using variable sampling interval CUSUM control charts. Journal of Statistical Computation and Simulation, 91(3):501– 521, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See correction to garbled paragraph [Ano21].

Tsou:2022:RST

[TH22] Tsung-Shan Tsou and Wei-Cheng Hsiao. Robust score test for treatment effects for paired continuous and ordinal data. Journal of Statistical Computation and Simulation, 92(16):3452– 3467, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2022:NCC

[THL^{+}22] Tzong-Ru Tsai, Chia-Min Hung, Y. L. Lio, Jyun-You Chiang, and H. K. T. Ng. Nonparametric control chart using auxiliary information for smallest extreme value distribution. Journal of Statistical Computation and Simulation, 92(6):1175–1193, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tanil:2022:NPS

[TK22] Halil Tanil and Agah Kozan. A nonparametric partially sequential test for early detection of newly emerging phenomena. Journal of Statistical Computation and Simulation, 92(7): 1426–1437, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomas:2024:OBG

[TK24] Julia T. Thomas and Mahesh Kumar. Optimal Bayesian generalized multiple-dependent state sampling plan for attributes. Journal of Statistical Computation and Simulation, 94(5):1073– 1096, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tan:2021:TND

[TL21] Wushuang Tan and Liu Liu. Truncated normal distributionbased EWMA control chart for monitoring the process mean in the presence of outliers. Journal of Statistical Computation and Simulation, 91(11):2276–2288, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tapia:2020:LRM

[TLGW20] Alejandra Tapia, Victor Leiva, Manuel Galea, and Rachel Werneck. On a logistic regression model with random intercept: diagnostic analytics, simulation and biological application. Journal of Statistical Computation and Simulation, 90(13):2354– 2383, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thanwane:2021:EME

[TMMCS21] Maonatlala Thanwane, Jean-Claude Malela-Majika, Philippe Castagliola, and Sandile Charles Shongwe. The effect of measurement errors on the performance of the homogeneously weighted moving average X^{−} monitoring scheme with estimated parameters. Journal of Statistical Computation and Simulation, 91(7):1306–1330, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Toker:2021:TLE

[TÖSY21] Selma Toker, Nimet Özbay, Gülesen Üstündag Siray, and Ismail Yenilmez. Tobit Liu estimation of censored regression model: an application to Mroz data and a Monte Carlo simulation study. Journal of Statistical Computation and Simulation, 91(6):1061–1091, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tomer:2022:RBE

[TR22] Sanjeev K. Tomer and Himanshu Rai. Robust Bayesian estimation of cumulative incidence function for competing risk data with missing causes. Journal of Statistical Computation and Simulation, 92(9):1781–1804, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tesfamichael:2024:CIP

[TR24] Asmerom Tesfamichael and Kent Riggs. Confidence intervals for a proportion using a fixed-inverse double sampling scheme when the data are subject to false-positive misclassification. Journal of Statistical Computation and Simulation, 94(3):499– 516, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsiotas:2020:AAC

[Tsi20] Georgios Tsiotas. An ABC approach for CAViaR models with asymmetric kernels. Journal of Statistical Computation and Simulation, 90(8):1373–1398, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Telford:2020:PAV

[TTWG20] Alison Telford, Charles C. Taylor, Henry M. Wood, and Arief Gusnanto. Properties and approximate p-value calculation of the Cramér test. Journal of Statistical Computation and Simulation, 90(11):1965–1981, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tavassoli:2022:CKA

[TWN22] Abbas Tavassoli, Yadollah Waghei, and Alireza Nazemi. Comparison of kriging and artificial neural network models for the prediction of spatial data. Journal of Statistical Computation and Simulation, 92(2):352–369, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tackney:2023:NPN

[TWS23] Mia Sato Tackney, Dave Woods, and Ilya Shpitser. Nonmyopic and pseudo-nonmyopic approaches to optimal sequential design in the presence of covariates. Journal of Statistical Computation and Simulation, 93(4):581–603, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tomal:2023:RRE

[TWZ23] Jabed H. Tomal, William J. Welch, and Ruben H. Zamar. Robust ranking by ensembling of diverse models and assessment metrics. Journal of Statistical Computation and Simulation, 93 (1):77–102, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tong:2022:DSA

[TZ22] Jianyang Tong and Xuejing Zhao. Deep survival algorithm based on nuclear norm. Journal of Statistical Computation and Simulation, 92(9):1964–1976, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noor-ul-Amin:2020:MTR

[uA20] Muhammad Noor ul Amin. Memory type ratio and product estimators for population mean for time-based surveys. Journal of Statistical Computation and Simulation, 90(17):3080–3092, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noor-ul-Amin:2021:PAE

[uAR21] Muhammad Noor ul Amin and Afshan Riaz. Performance of adaptive exponentially weighted moving average control chart in the presence of measurement error. Journal of Statistical Computation and Simulation, 91(11):2328–2343, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noor-ul-Amin:2023:DNA

[uAS23] Muhammad Noor ul Amin and Muhammad Atif Sarwar. Design of a new adaptive MEWMA chart to monitor the mean. Journal of Statistical Computation and Simulation, 93(16):2888–2905, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noor-ul-Amin:2020:EPE

[uAT20] Muhammad Noor ul Amin and Muhammad Tayyab. Enhancing the performance of exponential weighted moving average control chart using paired double ranked set sampling. Journal of Statistical Computation and Simulation, 90(6):1118–1130, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Unsal:2022:NBE

[Üns22] Mehmet Güray Ünsal. New Bayesian estimators for the incidence of inefficiency in the cases without a rule of thumb.

Journal of Statistical Computation and Simulation, 92(3):437– 450, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

vanderNest:2022:MFC

[vdNPCvB22] Gavin van der Nest, Valeria Lima Passos, Math J. J. M. Candel, and Gerard J. P. van Breukelen. Model fit criteria curve behaviour in class enumeration — a diagnostic tool for model (mis)specification in longitudinal mixture modelling. Journal of Statistical Computation and Simulation, 92(8):1640–1672, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Valle:2021:PGS

[VLRZ21] Luciana Dalla Valle, Fabrizio Leisen, Luca Rossini, and Weixuan Zhu. A Pólya-Gamma sampler for a generalized logistic regression. Journal of Statistical Computation and Simulation, 91(14):2899–2916, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Varshney:2023:OAM

[VPMA23] Rahul Varshney, Arun Pal, Mradula, and Irfan Ali. Optimum allocation in the multivariate cluster sampling design under Gamma cost function. Journal of Statistical Computation and Simulation, 93(2):312–323, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vanegas:2020:DTD

[VR20] Luis Hernando Vanegas and Luz Marina Rondon. A data transformation to deal with constant under/over-dispersion in Poisson and binomial regression models. Journal of Statistical Computation and Simulation, 90(10):1811–1833, 2020. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vaiciulyte:2020:RPE

[VS20] Jurate Vaiciulyte and Leonidas Sakalauskas. Recursive parameter estimation algorithm of the Dirichlet hidden Markov model. Journal of Statistical Computation and Simulation, 90(2):306– 323, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2022:AHS

[WB22] Chuyi Wang and Patrick Breheny. Adaptive hybrid screening for efficient lasso optimization. Journal of Statistical Computation and Simulation, 92(11):2233–2256, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2022:NBC

[WFY22] Sirao Wang, Kai-Tai Fang, and Huajun Ye. A new biascorrected estimator method in extreme value distributions with small sample size. Journal of Statistical Computation and Simulation, 92(18):3862–3884, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wimbush:2022:SCV

[WGF22] Alexander Wimbush, Nicholas Gray, and Scott Ferson. Singhing with confidence: visualising the performance of confidence procedures. Journal of Statistical Computation and Simulation, 92(13):2686–2702, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2024:HIV

[WKL^{+}24] Mingqiu Wang, Xiaoning Kang, Jiajuan Liang, Kun Wang, and Yuanshan Wu. Heteroscedasticity identification and variable selection via multiple quantile regression. Journal of Statistical Computation and Simulation, 94(2):297–314, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wasim:2023:MRR

[WKS23] Danish Wasim, Sajjad Ahmad Khan, and Muhammad Suhail. Modified robust ridge M-estimators for linear regression models: an application to tobacco data. Journal of Statistical Computation and Simulation, 93(15):2703–2724, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:2021:NTT

[WL21] Tony S. T. Wong and W. K. Li. A new test for tail index with application to Danish fire loss data. Journal of Statistical Computation and Simulation, 91(18):3880–3893, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Wan:2022:NBA

[WLLH22] Zhonglin Wan, Hongyan Li, Yi Luo, and Yirong Huang. A novel Bayesian approach to estimate long memory parameter. Journal of Statistical Computation and Simulation, 92(5):1078– 1091, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:GOS

[WLS^{+}21] Xiaoyi Wang, Baisen Liu, Ning-Zhong Shi, Guo-Liang Tian, and Shurong Zheng. Global one-sample tests for highdimensional covariance matrices. Journal of Statistical Computation and Simulation, 91(10):2051–2073, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:AOS

[WLZ21] Yu Wang, Nhu D. Le, and James V. Zidek. Approximately optimal subset selection for statistical design and modelling. Journal of Statistical Computation and Simulation, 91(12):2490– 2502, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wongoutong:2022:IMM

[Won22] Chantha Wongoutong. Imputation methods for missing response values in the three parts of a central composite design with two factors. Journal of Statistical Computation and Simulation, 92(11):2273–2289, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wei:2023:MRE

[WQZS23] Kecheng Wei, Guoyou Qin, Jiajia Zhang, and Xuemei Sui. Multiply robust estimation of the average treatment effect with missing outcomes. Journal of Statistical Computation and Simulation, 93(10):1479–1495, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

West:2022:SSC

[WR22] P. W. West and D. A. Ratkowsky. Simulation studies comparing fixed effect and mixed models in data sets with multiple measurements in individual sampling units. Journal of Statistical Computation and Simulation, 92(1):81–100, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Wywial:2020:ICP

[WS20] Janusz L. Wywiał and Grzegorz A. Sitek. On influence of clustering population on accuracy of population total estimation. Journal of Statistical Computation and Simulation, 90(2):234– 251, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2022:BCB

[WS22] Cang Wu and Shubin Si. Bivariate copula-based CUSUM charts for monitoring conditional nonlinear processes with first-order autocorrelation. Journal of Statistical Computation and Simulation, 92(16):3373–3399, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2020:BAP

[WTZ20] Yijun Wang, Yincai Tang, and Jiajia Zhang. Bayesian approach for proportional hazards mixture cure model allowing non-curable competing risk. Journal of Statistical Computation and Simulation, 90(4):638–656, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:IRF

[WW21] Hui Wang and Guizhi Wang. Improving random forest algorithm by lasso method. Journal of Statistical Computation and Simulation, 91(2):353–367, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2022:QUS

[WW22a] Qing Wang and Yujie Wei. Quantifying uncertainty of subsampling-based ensemble methods under a U-statistic framework. Journal of Statistical Computation and Simulation, 92(17):3706–3726, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2022:SCR

[WW22b] Yanhong Wu and Wei Biao Wu. Sequential common rate decrease detection, isolation, and estimation in multiple Poisson processes. Journal of Statistical Computation and Simulation, 92(18):3817–3835, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2023:MZT

[WW23] Yi Wu and Xuejun Wang. Marcinkiewicz–Zygmund type strong law of large numbers for weighted sums of random variables with infinite moment and its applications. Journal of Statistical Computation and Simulation, 93(11):1694–1715, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2022:EAD

[WWW22] Qiying Wu, Huiwen Wang, and Shanshan Wang. An efficient approach for discriminant analysis based on adaptive feature augmentation. Journal of Statistical Computation and Simulation, 92(16):3414–3429, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2023:SNC

[WWW23] Yi Wu, Wei Wang, and Xuejun Wang. Sufficient and necessary conditions for the strong consistency of LS estimators in simple linear EV regression models. Journal of Statistical Computation and Simulation, 93(8):1244–1262, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:EEP

[WWWZ21] Weiwei Wang, Yijun Wang, Xianyi Wu, and Xiaobing Zhao. Efficient estimation of panel count data with dependent observation process. Journal of Statistical Computation and Simulation, 91(3):464–476, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:CPP

[WX21] Zhendong Wang and Xingzhong Xu. Calibration of posterior predictive p-values for model checking. Journal of Statistical Computation and Simulation, 91(6):1212–1242, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:SAC

[WXWS21] Shuying Wang, Da Xu, Chunjie Wang, and Jianguo Sun. Semiparametric analysis of case K interval-censored failure time data in the presence of a cured subgroup and informative censoring. Journal of Statistical Computation and Simulation, 91(10):2018–2033, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2022:NEE

[WYBW22] Yi Wu, Wei Yu, Narayanaswamy Balakrishnan, and Xuejun Wang. Nonparametric estimation of expected shortfall via Bahadur-type representation and Berry—Esséen bounds. Journal of Statistical Computation and Simulation, 92(3):544–566, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2021:CMD

[WZ21] Yanxin Wang and Li Zhu. Coordinate majorization descent algorithm for nonconvex penalized regression. Journal of Statistical Computation and Simulation, 91(13):2684–2698, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xue:2022:TSC

[XCXZ22] Huiling Xue, Guanghui Chen, Tingting Xie, and Chun-Hua Zhang. Two-step calibration estimation under multiple auxiliary variables in survey sampling. Journal of Statistical Computation and Simulation, 92(14):3066–3084, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xia:2022:ESM

[XG22] Yunfei Xia and Michael Grabchak. Estimation and simulation for multivariate tempered stable distributions. Journal of Statistical Computation and Simulation, 92(3):451–475, February 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2022:AGM

[XLF22] L. L. Xu, J. Q. Li, and L. Y. Fu. Analysis of GEE with a mixture working correlation matrix for diverging number of covariates. Journal of Statistical Computation and Simulation, 92(4):748– 763, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiong:2020:HTI

[XP20] Li Xiong and Zuoxiang Peng. Heavy tail index estimation based on block order statistics. Journal of Statistical Computation and Simulation, 90(12):2198–2208, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2021:NHA

[XQ21] Min Xu and Zhongfeng Qin. A novel hybrid ARIMA and regression tree model for the interval-valued time series. Journal of Statistical Computation and Simulation, 91(5):1000–1015, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2020:PCB

[XSHC20] Kai Xu, Zhiling Shen, Xudong Huang, and Qing Cheng. Projection correlation between scalar and vector variables and its use in feature screening with multi-response data. Journal of Statistical Computation and Simulation, 90(11):1923–1942, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2023:RPT

[XYCZ23] Kai Xu, Hong Yin, Mengxiao Chen, and Jialong Zhang. A robust permutation test for Kendall’s tau. Journal of Statistical Computation and Simulation, 93(15):2780–2800, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2020:FDS

[XYZ20] Cong Xu, Min Yang, and Jin Zhang. Fast deflation sparse principal component analysis via subspace projections. Journal of Statistical Computation and Simulation, 90(8):1399–1412, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ye:2023:BIS

[YDL23] Rendao Ye, Weixiao Du, and Yiting Lu. Bootstrap inference for skew-normal unbalanced heteroscedastic one-way classification random effects model. Journal of Statistical Computation and Simulation, 93(15):2672–2702, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2020:EEV

[YDS^{+}20] Jun-Qiang Yang, Yuan Dong, Radhey Singh, Cheng Dong, and Xuewen Lu. Efficient estimation of a varying-coefficient partially linear proportional hazards model with current status data. Journal of Statistical Computation and Simulation, 90 (1):90–111, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yan:2021:GFI

[YGWH21] Liang Yan, Juan Geng, Lijun Wang, and Daojiang He. Generalized fiducial inference for the Lomax distribution. Journal of Statistical Computation and Simulation, 91(12):2402–2413, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yue:2020:NAS

[YH20] Mu Yue and Lei Huang. A new approach of subgroup identification for high-dimensional longitudinal data. Journal of Statistical Computation and Simulation, 90(11):2098–2116, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yildirim:2020:REA

[YK20] Vural Yildirim and Yeliz Mert Kantar. Robust estimation approach for spatial error model. Journal of Statistical Computation and Simulation, 90(9):1618–1638, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yalaz:2021:KEP

[YK21] Seçil Yalaz and Özge Kuran. Kernel estimator and predictor of partially linear mixed-effect errors-in-variables model. Journal of Statistical Computation and Simulation, 91(5):934–951, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yamaguchi:2023:RTP

[YKM23] Hikaru Yamaguchi, Masato Kitani, and Hidetoshi Murakami. Robust testing procedures for scale differences in paired data. Journal of Statistical Computation and Simulation, 93(12): 1899–1923, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yao:2020:PIC

[YLHZ20] Cailian Yao, Zhonghua Li, Chuan He, and Jiujun Zhang. A Phase II control chart based on the weighted likelihood ratio test for monitoring polynomial profiles. Journal of Statistical Computation and Simulation, 90(4):676–698, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Yeong:2022:VSI

[YLK^{+}22] W. C. Yeong, Sok Li Lim, Michael B. C. Khoo, Peh Sang Ng, and Z. L. Chong. A variable sampling interval run sum chart for the coefficient of variation. Journal of Statistical Computation and Simulation, 92(15):3150–3166, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yue:2022:CSB

[YLS22] Mu Yue, Jialiang Li, and Baoluo Sun. Conditional sparse boosting for high-dimensional instrumental variable estimation. Journal of Statistical Computation and Simulation, 92(15): 3087–3108, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2024:MSR

[YLYD24] Kai Yang, Ang Li, Xinyang Yu, and Xiaogang Dong. On MCMC sampling in random coefficients self-exciting integervalued threshold autoregressive processes. Journal of Statistical Computation and Simulation, 94(1):164–182, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yun:2020:SAH

[YSJL20] Jonghyun Yun, Minsuk Shin, Ick Hoon Jin, and Faming Liang. Stochastic approximation Hamiltonian Monte Carlo. Journal of Statistical Computation and Simulation, 90(17):3135–3156, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yan:2023:FOB

[YWW23a] Han Yan, Dehui Wang, and Zheqi Wang. First-order binomial autoregressive processes with Markov-switching coefficients. Journal of Statistical Computation and Simulation, 93(9):1378–1402, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2023:ELB

[YWW23b] Kaizhi Yu, Huiqiao Wang, and Christian H. Weiß. An empirical-likelihood-based structural-change test for INAR processes. Journal of Statistical Computation and Simulation, 93 (3):442–458, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2021:ANH

[YXC21] Hanjun Yu, Xinyi Xu, and Di Cao. The adaptive normalhypergeometric-inverted-beta priors for sparse signals. Journal of Statistical Computation and Simulation, 91(2):396–419, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yue:2022:PEC

[YXYL22] Chao Yue, Ma Xuejun, Li Yaguang, and Huang Lei. A penalized estimation for the Cox model with ordinal multinomial covariates. Journal of Statistical Computation and Simulation, 92(6): 1194–1223, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2020:MNA

[YY20] Xinfeng Yang and Xiaodong Yan. Mechanism and a new algorithm for nonconvex clustering. Journal of Statistical Computation and Simulation, 90(4):719–746, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2023:FBQ

[YY23] Hanjun Yu and Lichao Yu. Flexible Bayesian quantile regression for nonlinear mixed effects models based on the generalized asymmetric Laplace distribution. Journal of Statistical Computation and Simulation, 93(15):2725–2750, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yuan:2021:AWE

[YZ21] Yuze Yuan and Yong Zhou. Adaptive-weighted estimation of semi-varying coefficient models with heteroscedastic errors. Journal of Statistical Computation and Simulation, 91(15): 3029–3047, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yan:2022:WTM

[YZK22] Xiaofang Yan, Qi Zheng, and Maiying Kong. Weighted χ^{2} tests for multiple group comparisons in observational studies. Journal of Statistical Computation and Simulation, 92(13):2667– 2685, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2023:ROD

[YZZW23] Miaomiao Yu, Wei Zhao, Yong Zhou, and Chunjie Wu. Robust online detection on highly censored data using a semiparametric EWMA chart. Journal of Statistical Computation and Simulation, 93(9):1403–1419, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zavieh:2022:NMR

[ZADA22] Negar Hossein Zavieh, Mostafa Abouei Ardakan, and Hamed Davari-Ardakani. A new model for the reliability-redundancy allocation problem under the K-mixed redundancy strategy. Journal of Statistical Computation and Simulation, 92(17): 3542–3560, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:LCN

[ZC21] Jun Zhang and Xia Cui. Logarithmic calibration for nonparametric multiplicative distortion measurement errors models. Journal of Statistical Computation and Simulation, 91(13): 2623–2644, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zambom:2022:SNC

[ZCD22] Adriano Zanin Zambom, Julian Alfonso Collazos, and Ronaldo Dias. Selection of the number of clusters in functional data analysis. Journal of Statistical Computation and Simulation, 92(14):2980–2998, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2023:MPB

[ZFLL23] Yongchao Zhao, Ziheng Feng, Min Li, and Xinmin Li. Modified Penalized Blind Kriging for efficiently selecting a global trend model. Journal of Statistical Computation and Simulation, 93(17):3052–3066, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:MFG

[ZFZ22] Xiaochen Zhang, Kuangnan Fang, and Qingzhao Zhang. Multivariate functional generalized additive models. Journal of Statistical Computation and Simulation, 92(4):875–893, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zou:2023:CQR

[ZFZ23] Yuye Zou, Guoliang Fan, and Riquan Zhang. Composite quantile regression for heteroscedastic partially linear varyingcoefficient models with missing censoring indicators. Journal of Statistical Computation and Simulation, 93(3):341–365, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhai:2021:UPO

[ZG21] Ruoshui Zhai and Roee Gutman. Using the potential outcome framework to estimate optimal sample size for cluster randomized trials: a simulation-based algorithm. Journal of Statistical Computation and Simulation, 91(18):3744–3770, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhou:2024:TSN

[ZGO24] Jackson Zhou, Clara Grazian, and John T. Ormerod. Tractable skew-normal approximations via matching. Journal of Statistical Computation and Simulation, 94(5):1016–1034, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:ESP

[ZH22] Jing Zhang and Zhensheng Huang. Efficient simultaneous partial envelope model in multivariate linear regression. Journal of Statistical Computation and Simulation, 92(7):1373–1400, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:SDT

[Zha21a] Shibin Zhang. A spectral domain test of isotropic properties for irregularly spaced spatial data. Journal of Statistical Computation and Simulation, 91(1):151–166, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:IRL

[Zha21b] Tonglin Zhang. Iteratively reweighted least squares with random effects for maximum likelihood in generalized linear mixed effects models. Journal of Statistical Computation and Simulation, 91(16):3404–3425, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:MEM

[Zha22] Jun Zhang. Measurement errors models with exponential parametric multiplicative distortions. Journal of Statistical Computation and Simulation, 92(12):2467–2500, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:2021:MMS

[Zhe21] Songfeng Zheng. A majorization-minimization scheme for L_{2} support vector regression. Journal of Statistical Computation and Simulation, 91(15):3087–3107, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:PAM

[ZHF21] Feipeng Zhang, Xiaoyan Huang, and Caiyun Fan. Prediction accuracy measures for time-to-event models with left-truncated and right-censored data. Journal of Statistical Computation and Simulation, 91(13):2764–2779, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:MES

[ZKG21] Qi Zhang, Sadia Khalil, and Sat Gupta. Mean estimation in the simultaneous presence of measurement errors and non-response using optional RRT models under stratified sampling. Journal of Statistical Computation and Simulation, 91(17):3492–3504, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2024:TDN

[ZLFW24] Yuhao Zhang, Yanhong Liu, Long Feng, and Zhaojun Wang. Testing the differential network between two Gaussian graphical models with false discovery rate control. Journal of Statistical Computation and Simulation, 94(2):424–445, 2024. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2024:EVS

[ZLL24] Weihua Zhao, Rui Li, and Heng Lian. Estimation and variable selection of quantile partially linear additive models for correlated data. Journal of Statistical Computation and Simulation, 94(2):315–345, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:SRR

[ZLLB21] Fode Zhang, Rui Li, Heng Lian, and Dipankar Bandyopadhyay. Sparse reduced-rank regression for multivariate varyingcoefficient models. Journal of Statistical Computation and Simulation, 91(4):752–767, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhong:2022:FPC

[ZLLZ22] Rou Zhong, Shishi Liu, Haocheng Li, and Jingxiao Zhang. Functional principal component analysis estimator for non-Gaussian data. Journal of Statistical Computation and Simulation, 92(13):2788–2801, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2023:BSS

[ZLY23] Jie Zhang, Yang Li, and Qin Yu. Best subset selection with shrinkage: sparse additive hazards regression with the grouping effect. Journal of Statistical Computation and Simulation, 93(18):3382–3402, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zou:2024:AFS

[ZLZ24] Liying Zou, Yi Liu, and Zhonghu Zhang. Adjusted feature screening for ultra-high dimensional missing response. Journal of Statistical Computation and Simulation, 94(3):460–483, 2024. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:2023:MEM

[ZQAH23] Yu Zheng, Siying Qian, Sadia Arshad, and Jianfei Huang. A modified Euler–Maruyama method for Riemann–Liouville stochastic fractional integro-differential equations. Journal of Statistical Computation and Simulation, 93(2):249–265, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zamprogno:2020:PCA

[ZRB^{+}20] Bartolomeu Zamprogno, Valdério A. Reisen, Pascal Bondon, Higor H. Aranda Cotta, and Neyval C. Reis, Jr. Principal component analysis with autocorrelated data. Journal of Statistical Computation and Simulation, 90(12):2117–2135, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zapata:2023:IIW

[ZSS23] Zakry Zapata, Stephen A. Sedory, and Sarjinder Singh. An innovative improvement in Warner’s randomized response device for evasive answer bias. Journal of Statistical Computation and Simulation, 93(2):298–311, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:FAL

[ZSWN21] Zhuan Zhang, Zhenyuan Shen, Hong Wang, and Shu Kay Ng. A fast adaptive Lasso for the Cox regression via safe screening rules. Journal of Statistical Computation and Simulation, 91(14):3005–3027, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zareifard:2021:GSL

[ZTP21] Hamid Zareifard, Vahid Rezaei Tabar, and Dariusz Plewczynski. A Gibbs sampler for learning DAG: a unification for discrete and Gaussian domains. Journal of Statistical Computation and Simulation, 91(14):2833–2853, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2021:LBE

[ZW21] Fengyue Zhang and Lichun Wang. Linear Bayesian estimators for linear models with constraints. Journal of Statistical Computation and Simulation, 91(8):1635–1650, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2023:RHT

[ZW23] Yan Zhang and Yuehua Wu. Robust hypothesis testing in functional linear models. Journal of Statistical Computation and Simulation, 93(15):2563–2581, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2021:AME

[ZWW21] Wei Zhao, Zhijun Wang, and Chunjie Wu. Adaptive multivariate EWMA charts for monitoring sparse mean shifts based on parameter optimization design. Journal of Statistical Computation and Simulation, 91(13):2670–2683, 2021. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Zhao:2022:SBA

[ZXDD22] Yuanying Zhao, Dengke Xu, Xingde Duan, and Jiang Du. A semiparametric Bayesian approach to binomial distribution logistic mixed-effects models for longitudinal data. Journal of Statistical Computation and Simulation, 92(7):1438–1456, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2020:MDM

[ZXG20] Jun Zhang, Wangli Xu, and Yujie Gai. Multiplicative distortion measurement errors linear models with general moment identifiability condition. Journal of Statistical Computation and Simulation, 90(2):291–305, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:SCI

[ZXZ^{+}22] Qi Zhang, Jing Xu, Jianxin Zhao, Hua Liang, and Xinmin Li. Simultaneous confidence intervals for ratios of means of zeroinflated log-normal populations. Journal of Statistical Computation and Simulation, 92(6):1113–1132, 2022. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2020:SOE

[ZY20] Xing Zhang and Zhibin Yan. Second-order extended particle filter with exponential family observation model. Journal of Statistical Computation and Simulation, 90(12):2156–2179, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zuo:2021:TCR

[ZY21] Baishuai Zuo and Chuancun Yin. Tail conditional risk measures for location-scale mixture of elliptical distributions. Journal of Statistical Computation and Simulation, 91(17):3653–3677, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:MLR

[ZY22] Jun Zhang and Yiping Yang. Modal linear regression models with additive distortion measurement errors. Journal of Statistical Computation and Simulation, 92(4):847–874, 2022. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2020:LCP

[ZYFW20] Jun Zhang, Yiping Yang, Sanying Feng, and Zhenghong Wei. Logarithmic calibration for partial linear models with multiplicative distortion measurement errors. Journal of Statistical Computation and Simulation, 90(10):1875–1896, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2022:WCE

[ZYW22] Lu Zhang, Wei Yu, and Xuejun Wang. Weak consistency for the estimators in a semiparametric regression model based on negatively associated random errors. Journal of Statistical Computation and Simulation, 92(1):19–42, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2023:SCB

[ZYW23] Shen Zhang, Keying Ye, and Min Wang. A simple consistent Bayes factor for testing the Kendall rank correlation coefficient. Journal of Statistical Computation and Simulation, 93(6):888– 903, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2020:PSI

[ZZLL20] Weihua Zhao, Fode Zhang, Rui Li, and Heng Lian. Principal single-index varying-coefficient models for dimension reduction in quantile regression. Journal of Statistical Computation and Simulation, 90(5):800–818, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhu:2022:ADM

[ZZY22] Xuehu Zhu, Jun Zhang, and Yiping Yang. Additive distortion measurement errors regression models with exponential calibration. Journal of Statistical Computation and Simulation, 92(14):3020–3044, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.